### v1

Abstract.
Conformal prediction (CP) is an Uncertainty Representation technique that delivers finite-sample calibrated prediction regions for any underlying Machine Learning model, yet its status as an Uncertainty Quantification (UQ) tool has remained conceptually opaque. We adopt a category-theoretic approach to CP—framing it as a morphism, embedded in a commuting diagram, of two newly-defined categories—that brings us three joys. First, we show that—under minimal assumptions—CP is intrinsically a UQ mechanism, that is, its UQ capabilities are a structural feature of the method. Second, we demonstrate that CP bridges (and perhaps subsumes) the Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning. Finally, we show that a conformal prediction region (CPR) is the image of a covariant functor. This observation is relevant to AI privacy: It implies that privacy noise added locally does not break coverage.
Key words and phrases:
Category Theory; Conformal Prediction; Set-Valued Functions
2020 Mathematics Subject Classification:
Primary: 18D99; Secondary: 62G07; 28B20
1. Introduction
Category Theory (CT) can be thought of as a general theory of mathematical structures and their relations. Recently, there has been a surging interest in studying Machine Learning (ML) from a CT perspective, to leverage the unifying power of the latter (thus unifying apparently unrelated ML subfields), and to obtain more profound theoretical results.
For example, Stehlík, (2012) and Song, (2024) survey how to frame statistical learning models into a CT framework. Shiebler et al., (2021) do the same for gradient-based methods, probabilistic methods, and invariant and equivariant learning. Furthermore, Cornish, (2025) expands on the latter by studying the problem of symmetrizing a neural network along a group homomorphism. That is, given two groups $H$ and $G$, they study a procedure that converts $H$-equivariant neural networks to $G$-equivariant ones. More in general, CT approaches to probability and statistics (which play a pivotal role in modern-day ML) are active fields of study, see e.g. Fritz et al., (2025); Bohinen and Perrone, (2025); Ensarguet and Perrone, (2024); Ackerman et al., (2024); Matache et al., (2022), and are widely discussed in the nLab and in the n-Category Café.
Conformal Prediction (CP, introduced formally in Section 2.1) is a popular ML prediction tool. Suppose we are given an $n$-sized exchangeable training data $y^{n}=(y_{1},\ldots,y_{n})$ where each observation $y_{i}$ belongs to the state space $Y$, a confidence level $\alpha\in[0,1]$, and a non-conformity score $\psi$ (i.e. a function that measures “how different” a new observation is with respect to the ones in the training set). CP, then, returns a prediction set, called the Conformal Prediction Region (CPR), that contains the “true value” $y_{n+1}$ with high probability $1-\alpha$ (Vovk et al.,, 2005; Shafer and Vovk,, 2008; Vovk,, 2013; Angelopoulos et al.,, 2024). [^2] For extensions, and more modern references, see e.g. (Caprio et al., 2025b, ; Javanmardi et al.,, 2024; Gibbs et al.,, 2023; Barber et al.,, 2023; Papadopoulos,, 2008). In Caprio et al., 2025a (, Section 5.1), the authors show how CP can be seen as a correspondence (i.e. a set-valued function), thus highlighting a connection between ML and functional analysis (FA, Aliprantis and Border, (2006)). Other such connections have been explored in the literature, see e.g. Korolev, (2022).
This paper builds on the CP-as-a-correspondence intuition to develop a category-theoretic approach to CP, which allows us to expand on how we understand the structural underpinnings of distribution-free uncertainty quantification. [^3] Throughout the paper, we do not use the form “we/our” as a form of pluralis majestatis, but rather to comply to the convention of scientific papers of writing in plural form. We cast CP as a morphism (embedded in a commuting diagram) of two newly-defined categories, and we are also able to frame a CPR as a functor image.
Contributions. This category-theoretic approach grants us three joys, which are briefly summarized next.
  • CP is not only an uncertainty representation, but also an uncertainty quantification tool: The scholars who improperly referred to CP as an Uncertainty Quantification (UQ) methodology, were right all along.
  • CP bridges Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning, thus formalizing an intuition first put forth by Martin, (2022).
  • Seeing the Conformal Prediction Region as the image of a covariant functor directly benefits AI privacy (Fang et al.,, 2025): (i) Any locally added privacy noise leaves coverage guarantees intact, and (ii) In federated learning (Liu et al.,, 2024), each agent need only share a summary object (the credal set $\mathcal{M}$ that we introduce in the following paragraphs) instead of raw data.
As we can see, our Categorical Conformal Prediction approach exhibits many boons, and little to no banes (minimal assumptions on $Y$, choice of $\psi$, and $\alpha$).
To make the paper largely self-contained, we give the necessary background on CP in Section 2. There, we also provide some concepts from the Imprecise Probability (IP) literature. IPs (Walley,, 1991; Augustin et al.,, 2014; Troffaes and de Cooman,, 2014) are mathematical models that allow the scholar to reason in the presence of uncertainty (as for classical probabilities), and ambiguity (that is, ignorance, full or partial, about the “true” data generating process). [^4] For more modern references, see e.g. (Caprio and Gong,, 2023; Caprio and Mukherjee,, 2023; Caprio and Seidenfeld,, 2023; Caprio et al., 2024b, ; Caprio,, 2024, 2025). Their applications to ML form a lively subject field, see e.g. Caprio et al., 2024a ; Sale et al., (2023, 2024); Caprio et al., 2024c ; Lu et al., (2024); Dutta et al., (2024); Chau et al., (2025). Functional Analysis and Category Theory notions are brushed up in Appendices A and B, respectively. The structure of the paper can be summarized as follows.
In Sections 3 and 4, we introduce two categories, $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$, respectively. For both, the objects are topological spaces; The morphisms are, respectively, upper hemicontinuous, and weakly measurable correspondences. We study these categories because continuity and measurability are fundamental—but generally non-equivalent—concepts in functional analysis. Under minimal assumptions on the state space $Y$, on the choice of the credibility level $\alpha$, and on the class $\mathscr{F}$ of non-conformity scores $\psi$, we show that CP is a morphism of both. Viewing CP as a morphism rather than a stand-alone algorithm lets ML engineers plug it into functional-style libraries (e.g. JAX or PyTorch Functorch) and obtain automatic batching, JIT compilation, and differentiability of the score map, while the coverage guarantee remains mathematically intact.
Not only is the CP correspondence a morphism of these categories, but it is also part of a diagram that commutes. As we shall see (in particular in Example 1 and in the discussion thereafter), this discovery implies the intrinsic UQ capabilities of CP.
In Section 5, we show that, under minimal assumption, a diagram that codifies the way Bayesian CP (Fong and Holmes,, 2021), classical CP, and CP-informed imprecise probability theory generate an $\alpha$-level prediction region, commutes. This is a strong indicator that CP forms a bridge between, and perhaps subsumes, the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning (Martin,, 2022).
Section 6 expands on category-theoretic Imprecise Probability theory (Liell-Cock and Staton,, 2024). We formalize the notion of Imprecise Highest Density Region (Coolen,, 1992)—that can be thought of as a “robust version” of a confidence interval which also accounts for distribution uncertainty—as a (covariant) functor $\mathsf{IHDR}_{\alpha}$, $\alpha\in[0,1]$, between the category of closed and convex sets of finitely additive probabilities on $Y$, and that of subsets of $Y$, both with inclusion morphisms. In turn, the Conformal Prediction Region (CPR) can be seen as the image $\mathsf{IHDR}_{\alpha}(\mathcal{M})$ of such a functor at the set of probabilities $\mathcal{M}$ associated with the CP procedure (Cella and Martin,, 2022). This has e.g. the following promising implication. If we have many sources producing credal sets $\mathcal{M}_{j}$, any monotone, measurable or privacy-preserving operation applied locally to each credal set $\mathcal{M}_{j}$ propagates automatically through the functor. The resulting prediction regions, then, still satisfy the global coverage guarantee while only the (possibly privatized) summary objects $\mathcal{M}_{j}$, never the raw data, need be shared.
Finally, Section 7 studies the properties of $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$. In particular, we show that (after slightly restricting the former) they are both monoidal, and find faithful functors between them. In addition, we find which of their subcategories admit a monad, give sufficient conditions for CP to be a morphism in such subcategories, and find a faithful functor between them. Section 8 concludes our work, and discusses some open research questions.
2. Background
In this section, we introduce (Full) Conformal Prediction (Section 2.1), and some basic notions of Imprecise Probability Theory (Section 2.2). The reader familiar with any of these concepts can skip the associated part.
2.1. Full Conformal Prediction
This section is an expanded reproduction of Caprio et al., 2025a (, Section 2.1). It is based on Cella and Martin, (2022, Section 4.1), who summarize work by Shafer and Vovk, (2008) and Vovk et al., (2005) on Transductive (of Full) Conformal Prediction.
Suppose that there is an exchangeable process $\mathcal{Y}_{1},\mathcal{Y}_{2},\ldots$ with distribution $\mathfrak{P}$, where each $\mathcal{Y}_{i}$ is a random element taking values in the measurable space $(Y,\Sigma_{Y})$. Recall that a sequence is exchangeable if, for any $k\in\mathbb{N}$ and any permutation $\text{perm}(\cdot)$, the two random vectors $(\mathcal{Y}_{1},\ldots,\mathcal{Y}_{k})$ and $(\mathcal{Y}_{\text{perm}(1)},\ldots,\mathcal{Y}_{\text{perm}(k)})$ have the same joint distribution. This implies that the marginal distributions of the $\mathcal{Y}_{i}$’s are the same. [^5] Exchangeability does not imply independence, so the standard i.i.d. setup is in a sense more restrictive. Note also that we are not assuming any parametric form for the distribution $\mathfrak{P}$.
We want to solve the following statistical problem. Suppose we observe the first $n$ terms of the process, that is, $\mathcal{Y}^{n}=(\mathcal{Y}_{1},\ldots,\mathcal{Y}_{n})$. With this data, and the assumption of exchangeability, the goal is to predict $\mathcal{Y}_{n+1}$ using a method that is valid or reliable in a certain sense.
Let $\mathcal{Y}^{n+1}=(\mathcal{Y}^{n},\mathcal{Y}_{n+1})$ be an $(n+1)$-dimensional vector consisting of the observable $\mathcal{Y}^{n}$ and the yet-to-be-observed value $\mathcal{Y}_{n+1}$. Consider the transform
$\mathcal{Y}^{n+1}\rightarrow T^{n+1}=(T_{1},\ldots,T_{n+1})$
defined by the rule
$T_{i}\coloneqq\psi_{i}\left(\mathcal{Y}^{n+1}\right)\equiv\psi\left(y^{n+1}_{-i},y_{i}\right)$
for all $i\in\{1,\ldots,n+1\}$, where $y^{n+1}_{-i}=y^{n+1}\setminus\{y_{i}\}$ and $\psi:Y^{n}\times Y\rightarrow\mathbb{R}$ is a fixed function that is invariant to permutations in its first vector argument. Function $\psi$, which is called a non-conformity measure, is constructed in such as way that $\psi_{i}(y^{n+1})$ is small if and only if $y_{i}$ agrees with—i.e. is “close to”—a prediction based on the data $y^{n+1}_{-i}$. Or, the other way around, large values $\psi_{i}(y^{n+1})$ suggest that the observation $y_{i}$ is “strange” and does not conform to the rest of the data $y^{n+1}_{-i}$. The key idea is to define $\psi_{i}(y^{n+1})$ in a way that allows one to compare $y_{i}$ to a suitable summary of $y^{n+1}_{-i}$, e.g.  $\psi_{i}(y^{n+1})=|\text{mean}(y^{n+1}_{-i})-y_{i}|$, for all $i\in\{1,\ldots,n+1\}$. Notice that transformation $\mathcal{Y}^{n+1}\rightarrow T^{n+1}$ preserves exchangeability.
As the value $\mathcal{Y}_{n+1}$ has not yet been observed, and is actually the prediction target, the above calculations cannot be carried out exactly. Nevertheless, the exchangeability-preserving properties of the transformations described above provide a procedure to rank candidate values $\tilde{y}$ of $\mathcal{Y}_{n+1}$ based on the observed $\mathcal{Y}^{n}=y^{n}$, as shown in Algorithm 1, where $\mathbbm{1}[\cdot]$ denotes the indicator function.
Algorithm 1 Full Conformal prediction (CP)
Initialize: data $y^{n}$, non-conformity measure $\psi$, grid of $\tilde{y}$ values
for each $\tilde{y}$ value in the grid do
     set $y_{n+1}=\tilde{y}$ and write $y^{n+1}=y^{n}\cup\{y_{n+1}\}$;
     define $T_{i}=\psi_{i}(y^{n+1})$, for all $i\in\{1,\ldots,n+1\}$;
     evaluate $\pi(\tilde{y},y^{n})=(n+1)^{-1}\sum_{i=1}^{n+1}\mathbbm{1}[T_{i}\geq T_{n+1}]$;
end for
return $\pi(\tilde{y},y^{n})$ for each $\tilde{y}$ on the grid.
The output of Algorithm 1 is a data-dependent function $\tilde{y}\mapsto\pi(\tilde{y},y^{n})\in[0,1]$ that can be interpreted as a measure of plausibility of the assertion that $\mathcal{Y}_{n+1}=\tilde{y}$, given data $y^{n}$. Vovk et al., (2005) refer to the function $\pi$ as conformal transducer. Conformal transducer $\pi$ plays a key role in the construction of conformal prediction regions (CPRs). For any $\alpha\in[0,1]$, the $\alpha$-level CPR is defined as (Vovk,, 2013, Equation (2))
$\mathscr{R}_{\alpha}^{\psi}(y^{n})\coloneqq\{y_{n+1}\in Y:\pi(y_{n+1},y^{n})>\alpha\},$ (1)
and it satisfies
$P\left[\mathcal{Y}_{n+1}\in\mathscr{R}_{\alpha}^{\psi}(y^{n})\right]\geq 1-\alpha,$ (2)
uniformly in $n$ and in $P$ (Vovk et al.,, 2005). That is, (2) is satisfied for all $n\in\mathbb{N}$ and all exchangeable distributions $P$. Note that (2) holds regardless of the choice of the non-conformity measure $\psi$. However, the choice of this function is crucial in terms of the efficiency of conformal prediction, that is, the size of the prediction regions. A simple visual representation of Full CP is given in Figure 1.
Figure 1. A visual representation of Full CP methodology (Caprio et al., 2025a, , Figure 1). The Conformal Prediction Region is obtained as in (1).
Caprio et al., 2025a (, Section 5) point out how Full CP can be written as a correspondence
$\kappa:[0,1]\times Y^{n}\times\mathscr{F}\rightrightarrows Y,\quad(\alpha,y^{n},\psi)\mapsto\kappa(\alpha,y^{n},\psi),$ (3)
where
[…]
and the image of $\kappa$ is defined as
$\kappa(\alpha,y^{n},\psi)\coloneqq\Bigg{\{}y_{n+1}\in Y:\underbrace{\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}\left[\psi\left(y^{n+1}_{-i},y_{i}\right)\geq\psi\left(y^{n+1}_{-(n+1)},y_{n+1}\right)\right]}_{\eqqcolon\pi(y_{n+1},y^{n})}>\alpha\Bigg{\}}.$ (5)
It is easy to see, then, that $\kappa(\alpha,y^{n},\psi)\equiv\mathscr{R}_{\alpha}^{\psi}(y^{n})$.
Full Conformal Prediction (FCP) is not the only way of carrying out conformal prediction. There exists another type called Inductive (or Split) Conformal Prediction (ICP). It assumes exchangeability and it is used to build the same region $\mathscr{R}_{\alpha}^{\psi}(y^{n})$ in (1) having the same guarantee (2), while being computationally less expensive than FCP. Since ICP produces the same region as FCP, which is what we need in light of the goal of our paper, we do not present ICP in detail. We refer the interested reader to (Papadopoulos,, 2008; Angelopoulos et al.,, 2024).
2.2. Imprecise Probabilities
We now introduce a few imprecise probabilistic concepts that will be used when studying CP in the Category Theory framework.
Let $(Y,\Sigma_{Y})$ be a measurable space, and call $\Delta_{Y}$ the space of finitely additive probabilities on $Y$, endowed with the weak topology. A credal set $\mathcal{M}$ is a weak-closed and convex element of $\Delta_{Y}$ (Levi,, 1980). Throughout the paper, we will denote by $\mathscr{C}\subset 2^{\Delta_{Y}}$ the space of nonempty weak-closed and convex subsets of $\Delta_{Y}$. For any non-empty credal set $\mathcal{M}\in\mathscr{C}$, define the upper probability as the upper envelope of $\mathcal{M}$,
$\overline{P}_{\mathcal{M}}(A)\equiv\overline{P}(A)\;=\;\sup_{P\in\mathcal{M}}P(A),\quad A\in\Sigma_{Y}.$
The conjugate lower probability $\underline{P}$ of $\overline{P}$ satisfies $\underline{P}(A)=1-\overline{P}(A^{c})$, for all $A\in\Sigma_{Y}$, thus knowing one immediately gives us the other. In a statistical and ML framework, these concepts allow us to crystallize model ambiguity.
Even in the presence of the latter, we are able to define (a version of) a confidence interval. Fix a credibility level $\alpha\in[0,1]$. Then, the $\alpha$-level Imprecise Highest Density Region (IHDR) of credal set $\mathcal{M}$ is the set $B_{\mathcal{M}}^{\alpha}\subset Y$ such that $\underline{P}(B_{\mathcal{M}}^{\alpha})=1-\alpha$, and its “size is minimal” (Coolen,, 1992). More formally (Caprio et al., 2025a, , Equation (11)),
$B_{\mathcal{M}}^{\alpha}=\bigcap\left\{{A\in\Sigma_{Y}:\underline{P}(A)\geq 1-\alpha}\right\}.$
Notice that $B_{\mathcal{M}}^{\alpha}$ can be written as the image of the following correspondence
$\mathsf{IHDR}:[0,1]\times\mathscr{C}\rightrightarrows Y,\quad(\alpha,\mathcal{M})\mapsto\mathsf{IHDR}(\alpha,\mathcal{M})\coloneqq\bigcap\left\{{A\in\Sigma_{Y}:\underline{P}(A)\geq 1-\alpha}\right\}.$ (6)
Conformal Prediction is associated with a special case of upper probability $\overline{P}$, called a possibility function (Cuzzolin,, 2020). The latter is an upper probability for which there exists a function $\pi:Y\rightarrow[0,1]$ such that
  • $\sup_{y\in\mathbb{Y}}\pi(y)=1$;
  • $\overline{P}(A)=\sup_{y\in A}\pi(y)$, $A\in\Sigma_{Y}$.
Cella and Martin, (2022) show that, under the so-called consonance assumption, i.e. if $\sup_{y_{n+1}\in Y}\pi(y_{n+1},y^{n})=1$, then conformal transducer $\pi$ induces a possibility function $\overline{\Pi}(A)\coloneqq\sup_{y_{n+1}\in A}\pi(y_{n+1},y^{n})$, for all $A\in\Sigma_{Y}$, and in turn a credal set $\mathcal{M}(\overline{\Pi})\coloneqq\{P:P(A)\leq\overline{\Pi}(A),\forall A\in\Sigma_{Y}\}$. Such a credal set can be seen as the image of the correspondence $\mathsf{CRED}:Y^{n}\times\mathscr{F}\rightrightarrows\Delta_{Y}$, where $\mathscr{F}$ is the collection of functions in (4), and $(y^{n},\psi)\mapsto\mathsf{CRED}(y^{n},\psi)\coloneqq$
$\Bigg{\{}P\in\Delta_{Y}:P(A)\leq\underbrace{\sup_{y_{n+1}\in A}\left(\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}\left[\psi\left(y^{n+1}_{-i},y_{i}\right)\geq\psi\left(y^{n+1}_{-(n+1)},y_{n+1}\right)\right]\right)}_{\equiv\sup_{y_{n+1}\in A}\pi(y_{n+1},y^{n})\eqqcolon\overline{\Pi}(A)}\text{, }\forall A\in\Sigma_{Y}\Bigg{\}}.$ (7)
The consonance assumption is needed to ensure that $\mathsf{CRED}(y^{n},\psi)\in\mathscr{C}$. In the remainder of the paper, we sometimes write $\pi(y,y^{n},\psi)\equiv\pi_{\psi}(y,y^{n})$ for notational convenience (we drop the $n+1$ subscript for $y_{n+1}$), and to explicitly acknowledge that the conformal transducer $\pi$ depends on the choice of the non-conformity measure $\psi$.
We also consider a slight modification of $\mathsf{CRED}$, that will considerably simplify our proving endeavors in Sections 3, 4, and 5; We denote it by $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\text{, }\forall A\in\Sigma_{Y}\text{ closed}\}$. As we can see, $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\supseteq\mathcal{M}(\overline{\Pi})$, because it imposes less constraints on the $P$’s: They only have to be setwise dominated on closed sets. The upper probability $\overline{\Pi}$ and its conjugate lower probability $\underline{\Pi}$ (the only characteristics of the credal sets that we use) stay the same for both $\mathsf{CRED}$ and $\mathsf{CRED}_{\text{cl}}$: This justifies our simplifying choice. To avoid notation complications, in the remainder of this work we do not distinguish between $\mathsf{CRED}(y^{n},\psi)$ and $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)$.
A visual representation of the data $\to$ credal set $\to$ IHDR route is given in Figure 2.
Figure 2. A visual representation of the two-step procedure that first extracts credal set $\mathcal{M}(\overline{\Pi})$ from the training set $y^{n}$ as in (7), and then computes the IHDR as in (6) (Caprio et al., 2025a, , Figure 1).
In addition, Caprio et al., 2025a (, Propositions 5 and 6) show that the $\alpha$-level IHDR associated with $\mathcal{M}(\overline{\Pi})$, computed using the conjugate lower probability $\underline{\Pi}$ of $\overline{\Pi}$, corresponds to the Conformal Prediction Region $\mathscr{R}_{\alpha}^{\psi}(y^{n})$; It enjoys the same probabilistic guarantee (2).
3. Conformal Prediction as an Upper Hemicontinuous Morphism
This section introduces category UHCont, and shows that, under minimal assumptions, the Conformal Prediction correspondence $\kappa$ is a morphism of UHCont, embedded in a commuting diagram. The following is inspired by Perrone, (2024, Example 1.1.14).
Definition 1 (Structure UHCont).
Define a structure $\mathbf{UHCont}$ as follows,
  • Objects: Topological spaces $(X,\tau_{X})$.
  • Morphisms: Upper hemicontinuous correspondences $\Phi:X\rightrightarrows Y$ (see Definition 26 in Appendix A).
Theorem 2 (UHCont is a Category).
The structure $\mathbf{UHCont}$ defines a category.
Proof.
We verify the axioms of a category.
(1) Identity morphisms. For each object $X$, define $\mathrm{id}_{X}(x)\coloneqq\{x\}$. Let $V\subseteq X$ be open and suppose $x_{0}\in X$ satisfies $\mathrm{id}_{X}(x_{0})=\{x_{0}\}\subseteq V$. Then, since $x_{0}\in V$ and $V$ is open, there exists an open neighborhood $U=V$ of $x_{0}$ such that for all $x\in U$, $\mathrm{id}_{X}(x)=\{x\}\subseteq V$. Hence, $\mathrm{id}_{X}$ is upper hemicontinuous.
(2) Associativity of composition. Let $\Phi:X\rightrightarrows Y$, $\Psi:Y\rightrightarrows Z$, and $\Theta:Z\rightrightarrows W$. Then for all $x\in X$,
$((\Theta\circ\Psi)\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\bigcup_{z\in\Psi(y)}\Theta(z)=\bigcup_{z\in(\Psi\circ\Phi)(x)}\Theta(z)=(\Theta\circ(\Psi\circ\Phi))(x)$
by the associativity of union. Thus, composition is associative.
(3) Closure under composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be upper hemicontinuous correspondences. By Aliprantis and Border, (2006, Theorem 17.23), the composition $\Psi\circ\Phi$ defined by
$(\Psi\circ\Phi)(x)\coloneqq\bigcup_{y\in\Phi(x)}\Psi(y)$
is also upper hemicontinuous. Thus, the composition of morphisms is again a morphism.
(4) Identity laws. For any morphism $\Phi:X\rightrightarrows Y$ and any $x\in X$,
$(\Phi\circ\mathrm{id}_{X})(x)=\bigcup_{z\in\mathrm{id}_{X}(x)}\Phi(z)=\Phi(x),\quad(1_{Y}\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\{y\}=\Phi(x).$
Hence, $\mathrm{id}_{X}$ acts as a left and right identity.
Therefore, $\mathbf{UHCont}$ satisfies the axioms of a category. ∎
We now show that, under minimal assumptions, the (Full) Conformal Prediction correspondence $\kappa$ in (3) is a morphism of $\mathbf{UHCont}$, and that it is part of a commutative diagram involving the IHDR and the credal set correspondences. Fix any $n\in\mathbb{N}$, and consider the following conditions,
  1. (i)
    $Y$ is compact Hausdorff and $\Sigma_{Y}=\mathcal{B}(Y)$ is its Borel $\sigma$-field.
  2. (ii)
    Each $\psi\in\mathscr{F}$ is jointly continuous on $Y^{n}\times Y$; We can then write $\mathscr{F}\subset C(Y^{n}\times Y)$, the space of jointly continuous real-valued functions on $Y^{n}\times Y$. $\mathscr{F}$ is endowed with the uniform ($\sup$-norm) topology.
  3. (iii)
    No-tie parameter: The credibility level $\alpha$ satisfies $\alpha\notin S_{n+1}\coloneqq\{0,\frac{1}{n+1},\frac{2}{n+1},\ldots,\frac{n}{n+1},1\}$.
  4. (iv)
    $\Delta_{Y}$ is endowed with the weak topology, and $\mathscr{C}\subset 2^{\Delta_{Y}}$ is endowed with the Vietoris topology (see Definition 29 in Appendix A).
The following is a famous classical result in set-valued functional analysis (Aliprantis and Border,, 2006, Theorem 17.11).
Lemma 3 (Closed Graph Theorem).
A correspondence with compact Hausdorff range space has closed graph if and only if it is upper hemicontinuous and closed-valued.
Another important lemma that sets the table for our main result is the following.
Lemma 4 (Upper Semicontinuity of the Upper Envelope $\overline{\Pi}$).
If (i) and (ii) hold, then the map $(y^{n},\psi)\mapsto\overline{\Pi}_{(y^{n},\psi)}(A)=\sup_{y\in A}\pi(y,y^{n},\psi)$ is upper semicontinuous, for all closed sets $A\in\Sigma_{Y}$.
Proof.
First, we show that, if a generic set $K$ is compact and a generic map $f:K\to\mathbb{R}$ is upper semicontinuous (u.s.c.), then $f$ attains its maximum on $K$. Choose $x_{m}$ with $f(x_{m})\geq\sup_{K}f-\tfrac{1}{m}$. A compactness subsequence $x_{m_{k}}\to x_{\star}\in K$ exists. Upper semicontinuity gives $f(x_{\star})\geq\limsup f(x_{m_{k}})=\sup_{K}f.$
Now, we introduce auxiliary maps, and show their continuity. For $i\in\{1,\dots,n+1\}$, put
$g_{i}(y_{n+1},y^{n},\psi)\coloneqq\psi\bigl{(}y^{n+1}_{-i},y_{i}\bigr{)}-\psi\bigl{(}y^{n},y_{n+1}\bigr{)}.$
Because (a) the evaluation $\psi\mapsto\psi(y^{n},y)$ is continuous in the sup-norm, and (b) coordinate deletion $y^{n+1}\mapsto(y^{n+1}_{-i},y_{i})$ is continuous, each $g_{i}$ is continuous on $Y^{n+1}\times\mathscr{F}$.
Let us now show the upper semicontinuity of $\pi_{\psi}$. Fix $(y^{n+1},\psi)$. The set $\{g_{i}\geq 0\}$ is closed; For $k=0,\dots,n+1$ define
$R_{k}\coloneqq\Bigl{\{}(y^{n+1},\psi)\,:\,\pi_{\psi}(y_{n+1},y^{n})\geq\tfrac{k}{n+1}\Bigr{\}}.$
At least $k$ of the inequalities $g_{i}\geq 0$ must hold, so $R_{k}=\bigcup_{\lvert I\rvert=k}\bigcap_{i\in I}\{g_{i}\geq 0\}$, a finite union of finite intersections of closed sets; Hence $R_{k}$ is closed and $\pi_{\psi}$ is u.s.c.
We are now ready to prove the main statement. Fix any closed set $A\in\Sigma_{Y}$. By (i), it is also compact. Let $\bigl{(}y^{n}_{m},\psi_{m}\bigr{)}_{m\geq 1}$ converge to $(y^{n},\psi)$ and set $L\coloneqq\limsup_{m\to\infty}\overline{\Pi}_{(y^{n}_{m},\psi_{m})}(A)$. [^6] If $Y$ is not first countable, we can switch from a sequence to a net argument. We must show $L\leq\overline{\Pi}_{(y^{n},\psi)}(A)$.
Step 1 (Discretization). Because $\pi_{\psi}$ takes values in the finite set $\{0,\tfrac{1}{n+1},\dots,1\}$, there is an integer $k\in\{0,\dots,n+1\}$ and a subsequence (still indexed by $m$) with $\overline{\Pi}_{(y^{n}_{m},\psi_{m})}(A)=k/(n+1)$, for all $m$.
Step 2 (Select Maximizers). By the attained maximum result that we presented earlier, pick $y_{n+1}^{(m)}\in A$ such that $\pi_{\psi_{m}}\bigl{(}y_{n+1}^{(m)},y^{n}_{m}\bigr{)}=k/(n+1)$. Compactness of $A$ yields a further subsequence and a limit $y_{n+1}\in A$ with $y_{n+1}^{(m)}\to y_{n+1}$.
Step 3 (Pass to the Limit). For every $i$, the continuity of $g_{i}$ gives $g_{i}\bigl{(}y_{n+1}^{(m)},y^{n}_{m},\psi_{m}\bigr{)}\to g_{i}\bigl{(}y_{n+1},y^{n},\psi\bigr{)}.$ Since at least $k$ of the $g_{i}$’s are non-negative for each $m$, suppose, for contradiction, that strictly fewer than $k$ are non-negative at the limit. Then there exists $\delta>0$ such that, for all sufficiently large $m$, fewer than $k$ indices satisfy $g_{i}\bigl{(}y_{n+1}^{(m)},y^{n}_{m},\psi_{m}\bigr{)}\geq-\delta$, contradicting the fact that every $m$ has $k$ indices with the same expression $\geq 0$. Hence at least $k$ of the limiting $g_{i}$’s are $\geq 0$, and
$\pi_{\psi}(y_{n+1},y^{n})\;\geq\;\frac{k}{n+1}.$
Step 4 (Take the Supremum). Because $y_{n+1}\in A$,
$\overline{\Pi}_{(y^{n},\psi)}(A)\;\geq\;\pi_{\psi}(y_{n+1},y^{n})\;\geq\;\frac{k}{n+1}\;=\;L.$
Thus $\limsup_{m\to\infty}\overline{\Pi}_{(y^{n}_{m},\psi_{m})}(A)\leq\overline{\Pi}_{(y^{n},\psi)}(A)$, establishing upper semicontinuity. ∎
We are now ready for the main result of this section.
Theorem 5 (The Conformal Prediction Diagram Commutes, $\mathbf{UHCont}$).
Assume that (i)-(iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram
${Y^{n}\times\mathscr{F}}$${\mathscr{C}}$${\Sigma_{Y}}$$\scriptstyle{\mathsf{CRED}}$$\scriptstyle{\kappa_{\alpha}}$$\scriptstyle{\mathsf{IHDR}_{\alpha}}$
commutes in $\mathbf{UHCont}$.
In the Full CP Diagram that factorizes $\kappa_{\alpha}=\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ inside $\mathbf{UHCont}$, for all $\alpha$ satisfying (iii), we indicated $\mathscr{C}$ as the image set of $\mathsf{CRED}$, and $\Sigma_{Y}$ as the image set of $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$. This is an immediate consequence of writing our three correspondences in their functional form $\kappa_{\alpha}^{\text{funct}}:Y^{n}\times\mathscr{F}\to\Sigma_{Y}$, $\mathsf{IHDR}_{\alpha}^{\text{funct}}:\mathscr{C}\to\Sigma_{Y}$, and $\mathsf{CRED}^{\text{funct}}:Y^{n}\times\mathscr{F}\to\mathscr{C}$, where the latter holds under consonance. Notice also that $\Sigma_{Y}=\mathcal{B}(Y)$ by (i).
Proof of Theorem 5.
Fix any $n\in\mathbb{N}$. We first show that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ have a closed graph.
(i) $\kappa$. By (ii) and its definition, the conformal transducer $\pi(\cdot,y^{n},\psi):Y\to[0,1]$ is upper semicontinuous. Because $\alpha$ avoids the grid $S_{n+1}$, let
$\beta(\alpha)\equiv\beta\;\coloneqq\;\min\Bigl{\{}\tfrac{k}{n+1}\in S_{n+1}:\tfrac{k}{n+1}>\alpha\Bigr{\}},\quad\text{so }0<\alpha<\beta\leq 1.$ (8)
Then, define
$G\;\coloneqq\;\bigl{\{}(\alpha,y^{n},\psi,y):\pi(y,y^{n},\psi)\geq\beta\bigr{\}}\subset[0,1]\times Y^{n}\times\mathscr{F}\times Y.$
This is a super-level set of $\pi(\cdot,y^{n},\psi)$, and since $\pi$ is upper semicontinuous, $G$ is closed. In turn, $\operatorname{Gr}(\kappa)=G$ is closed.
(ii) $\mathsf{CRED}$. For every closed $A\in\Sigma_{Y}$, set
$G_{A}\coloneqq\Bigl{\{}(y^{n},\psi,P):P(A)\leq\overline{\Pi}_{(y^{n},\psi)}(A)\Bigr{\}}.$
Weak-continuity of $P\mapsto P(A)$ and upper semicontinuity of the right-hand side (see Lemma 4) make $G_{A}$ closed. Because $\operatorname{Gr}(\mathsf{CRED})=\bigcap_{A\in\Sigma_{Y}}G_{A}$, the graph is closed.
(iii) $\mathsf{IHDR}$. For each $A\in\Sigma_{Y}$ define $f_{A}(\alpha,\mathcal{M})\coloneqq\underline{P}_{\!\mathcal{M}}(A)-(1-\alpha)$. Weak continuity of $P\mapsto P(A)$ and compactness of $\mathcal{M}$ imply that $\mathcal{M}\mapsto\underline{P}_{\!\mathcal{M}}(A)$ is upper semicontinuous in the Vietoris topology (iv), while $\alpha\mapsto 1-\alpha$ is continuous. Hence the set
$H_{A}\coloneqq\bigl{\{}(\alpha,\mathcal{M}):f_{A}(\alpha,\mathcal{M})\geq 0\bigr{\}}$
is closed; Moreover $(\alpha,\mathcal{M})\mapsto H_{A}(\alpha,\mathcal{M})$ is closed-valued by (i). In turn, $\operatorname{Gr}(\mathsf{IHDR})=\bigl{\{}(\alpha,\mathcal{M},y):y\in Y\setminus\bigcap_{A}H_{A}(\alpha,\mathcal{M})\bigr{\}}^{c}=\bigcap_{A}(H_{A}\times Y)$ is closed.
Notice also that $Y$ is compact Hausdorff by (i), and that $\Delta_{Y}$ is compact by (iv) and Banach-Alaoglu, and Hausdorff because it is a subset of a dual Banach space equipped with its weak topology.
It follows that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are upper hemicontinuous by Lemma 3. In turn, they are morphisms of $\mathbf{UHCont}$. Fix now some $\alpha\in[0,1]\setminus S_{n+1}$. The restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ remain upper hemicontinuous. Notice also that $t(\mathsf{CRED})=s(\mathsf{IHDR}_{\alpha})=\mathscr{C}$; That is, the target space of $\mathsf{CRED}$ and the source space of $\mathsf{IHDR}_{\alpha}$ coincide, and are equal to the space of credal sets $\mathscr{C}$. This is clear by looking at the functional forms $\mathsf{CRED}^{\text{funct}}$ and $\mathsf{IHDR}^{\text{funct}}_{\alpha}$ of $\mathsf{CRED}$ and $\mathsf{IHDR}_{\alpha}$, respectively.
$Y$ is a topological space from (i); $[0,1]\times\mathscr{C}$ is a topological space in the product topology by (iv); $Y^{n}\times\mathscr{F}$ is a topological space in the product topology by (ii); $[0,1]\times Y^{n}\times\mathscr{F}$ is a topological space in the product topology by (i) and (ii). Hence, these spaces are all objects of $\mathbf{UHCont}$.
The fact that the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}$, then, is an immediate consequence of Caprio et al., 2025a (, Proposition 5). It states that, for all $\alpha\in[0,1]$, $y^{n}\in Y^{n}$, and $\psi\in\mathscr{F}$, the Conformal Prediction Region $\kappa_{\alpha}(y^{n},\psi)$ corresponds to the $\alpha$-level IHDR of the credal set $\mathsf{CRED}(y^{n},\psi)$. ∎
Let us now give an example in which conditions (i) and (ii) are satisfied. We do not verify (iii), as it is the user that chooses $\alpha$, so the no-tie parameter will always be met, and (iv), because it is a technical assumption that should not worry the practitioner.
Example 1 (An ML-Inspired Setting).
Let $Y=[0,1]^{d}$ with its Euclidean topology (compact and Hausdorff), and let $\varphi_{\theta}\colon Y\to\mathbb{R}^{m}$ be a feed-forward neural network whose activation functions are continuous (e.g. ReLU). Hence, $\varphi_{\theta}$ is continuous in both its parameters $\theta$ and input $y$.
Non-Conformity Measure. For a training set $y^{n}=(y_{1},\dots,y_{n})\in Y^{n}$ and a candidate point $y\in Y$, define
$\psi_{\theta}(y^{n},y)\;=\;-\biggl{\|}\varphi_{\theta}(y)\;-\;\frac{1}{n}\sum_{i=1}^{n}\varphi_{\theta}(y_{i})\biggr{\|}_{2}^{2}.$ (9)
  • Permutation invariance: The sample mean is symmetric in $y_{1},\dots,y_{n}$.
  • Joint continuity: Composition, finite sums, Euclidean norm, and squaring preserve continuity; Thus $(y^{n},y)\mapsto\psi_{\theta}(y^{n},y)$ is jointly continuous on $Y^{n}\!\times\!Y$.
Hence, $\psi_{\theta}\in\mathscr{F}$. Notice that consonance, i.e. $\sup_{y\in Y}\pi(y,y^{n})=1$, does not automatically hold in this example. To enforce it, it is enough to consider the normalized conformal transducer $\tilde{\pi}(y,y^{n})\coloneqq\frac{\pi(y,y^{n})}{\sup_{y\in Y}\pi(y,y^{n})}$ (Cella and Martin,, 2021, Section 7), which is still upper semicontinuous.
Machine Learning Relevance. This example is interesting because of the following implications for Machine Learning.
  1. (1)
    Few-shot / prototype learning: (9) is the squared Euclidean distance to the class prototype $\text{mean}\,\varphi_{\theta}(y_{i})$.
  2. (2)
    Out-Of-Distribution (OOD) detection: The same distance acts as a learned Mahalanobis-style OOD score.
  3. (3)
    Heteroscedastic conformal regression: If $\varphi_{\theta}$ is the penultimate layer of a regression network, (9) yields locally adaptive, yet still valid, prediction intervals.
Hence, this choice of $\psi_{\theta}$ simultaneously fulfills joint continuity and practical usefulness in modern ML workflows. $\triangle$
More in general, for the ML practitioner, it will be enough to
  • check that $Y$ is compact (always satisfied in classification problems and easy to enforce in regression ones), so that (i) is met,
  • pick $\alpha$ outside of $S_{n+1}$, so that (iii) is met,
  • pick a non-conformity score $\psi$ that is invariant to permutations in its first argument and jointly continuous in its arguments, so that (ii) is met,
  • check that the consonance assumption is met, or enforce it e.g. via a transformation, as we did in Example 1,
to be sure that the conditions of Theorem 5 are met. Condition (iv) is technical, and should not worry them.
Theorem 5 clarifies the nature of Conformal Prediction: Although often billed as an uncertainty quantification tool, CP furnishes a representation of predictive uncertainty, but not a quantification in the cardinal sense. The distinction is subtle yet essential: The diameter $\text{diam}(\mathscr{R})$ of a Conformal Prediction Region (CPR) $\mathscr{R}\subseteq Y$ is only defined up to strictly order-preserving (monotone) transformations. [^7] Having assumed a metric $d_{Y}$ on $Y$, $\text{diam}(A)=\sup_{y_{1},y_{2}\in A}d_{Y}(y_{1},y_{2})$, $A\subseteq Y$. A similar argument holds if we consider the volume in place of the diameter of the CPR. Therefore,
[…]
Conformal regions support ordinal comparisons (“larger” versus “smaller”) but not the cardinal arithmetic needed for expected-length-type calibration that variance-based approaches provide. Interpreting region width as a direct proxy for probability mass is therefore as misleading as treating a ranking as if it were a probability distribution.
The commuting diagram, however, reveals a hidden layer: Computing a CP region is equivalent to (i) forming the credal set $\mathcal{M}(\overline{\Pi})$ and (ii) extracting its IHDR. Because $\mathcal{M}(\overline{\Pi})$ lives in the space of probability measures, existing metrics—e.g. those in Abellán et al., (2006); Hüllermeier and Waegeman, (2021); Javanmardi et al., (2023); Hofman et al., (2024)—endow it with a cardinal scale that is able to quantify (different types of) uncertainty, in particular, reducible (epistemic) and irreducible (aleatoric) (Javanmardi et al.,, 2024, 2025; Cabezas et al.,, 2025; Stutts et al.,, 2024; Hanselle et al.,, 2025). This will be “marginal uncertainty”, [^8] That is, averaged over the possible realization $y^{n}$. thus reflecting the marginal guarantee of the Conformal Prediction Region (see equation (2)). We conjecture that the analyst-controlled choices—the significance level $\alpha$, which fixes the tolerated error, and the non-conformity score $\psi\in{\mathscr{F}}$, which embodies modeling assumptions—are primarily associated with epistemic uncertainty. On the other hand, aleatoric uncertainty stems from the intrinsic variability of the exchangeable process that we study, encoded in $\mathfrak{P}$.
Thus, Conformal Prediction’s capacity for predictive uncertainty quantification is not an add-on, but a built-in consequence of the categorical factorization resulting from Theorem 5’s commuting diagram. The scholars who have been calling Conformal Prediction a UQ method were right all along (under the stated assumptions), and Theorem 5 explains why. This expands the work initiated by Cella and Martin, (2022, 2023); Caprio et al., 2025a , who found UQ capabilities of CP methods through an imprecise probabilistic route.
Next, we present a remark that shows that conditions (i)-(iv) cannot be lightened.
Remark 1 (Minimality of Hypotheses (i)-(iv)).
We prove that each hypothesis is indispensable for the simultaneous upper hemicontinuity (u.h.c.) of the three correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$. For every clause, we remove the hypothesis in question while retaining the others, and we show that at least one correspondence fails to be upper hemicontinuous.
  • Drop (i) (compactness of $Y$). Let $Y=\mathbb{R}$ with the Euclidean topology (non-compact), $n=1$, and keep $\psi(y^{1},y)=-y$ (continuous, so (ii) holds). Fix $\alpha=\tfrac{3}{4}\notin\{0,\tfrac{1}{2},1\}$ (satisfies the no-tie rule). Writing $y^{1}=(x)$,
    $\kappa(\alpha,y^{1},\psi)=\{y\in\mathbb{R}:y\geq x\}\;=\;[x,\infty).$
    Take the sequence $x_{k}\nearrow x_{0}=1$. For the open set $V\coloneqq(1,\infty)\subset\kappa(\alpha,x_{0},\psi)=[1,\infty)$, no neighborhood of $(\alpha,x_{0},\psi)$ keeps $\kappa(\alpha,x_{k},\psi)\subseteq V$ (the left end-points $x_{k}$ lie outside $V$). Thus $\kappa$ is not u.h.c. at $(\alpha,x_{0},\psi)$.
  • Drop (ii) (joint continuity of $\psi$). Keep $Y=[0,1]$ (compact) and $n=1$. Define the pathological non-conformity score
    $\psi(y^{1},y)=\begin{cases}1&\text{if }y\in\mathbb{Q}\\ 0&\text{otherwise}\end{cases},$
    which is permutation-invariant but nowhere continuous. Take $\alpha=\tfrac{1}{2}$, $y^{1}=(0)$. Then, $\pi(y,y^{1})=1$ for rational $y$’s and $\pi(y,y^{1})=\tfrac{1}{2}$ for irrational $y$’s, so $\kappa(\alpha,y^{1},\psi)=\mathbb{Q}\cap[0,1]$, a dense set that is not closed. Because nearby bags $\bigl{(}y^{1}+\varepsilon\bigr{)}$ shuffle rationals and irrationals, the graph of $\kappa$ fails to be closed at $(\alpha,y^{1},\psi)$, hence $\kappa$ is not u.h.c.
  • Drop (iii) (no-tie rule). Work again with the continuous $\psi_{\theta}$ of Example 1, $n\geq 1$, and choose $\alpha=\tfrac{1}{n+1}$, which is in $S_{n+1}$. Let $y^{n}$ be such that all non-conformity scores tie (e.g. $y=y_{1}=\cdots=y_{n}$), so $\pi(y,y^{n})=\tfrac{1}{n+1}$, for every $y$. Then, $\kappa(\alpha,y^{n},\psi_{\theta})=\emptyset$, whereas for the sequence $\tfrac{1}{n+1}-\tfrac{1}{k}\eqqcolon\alpha_{k}\downarrow\alpha$ we have $\kappa(\alpha_{k},y^{n},\psi_{\theta})=Y$. Thus $\kappa$ is not u.h.c. at $\alpha$ unless the no-tie exclusion is imposed.
  • Drop (iv) (weak topology on $\Delta_{Y}$). Still with $Y=[0,1]$, equip $\Delta_{Y}$ with the total-variation norm. For any fixed closed Borel set $A$ the map $P\mapsto P(A)$ is not continuous in total variation (take Dirac masses sliding along $A$). Consequently, the constraint $P(A)\leq\overline{\Pi}(A)$ fails to define a closed set in $Y^{n}\times\mathscr{F}\times\Delta_{Y}$, so the graph of $\mathsf{CRED}$ is no longer closed and upper hemicontinuity breaks down.
    Drop (iv) (Vietoris topology on $\mathscr{C}$). Keep the weak topology on $\Delta_{Y}$, but put on $\mathscr{C}\subset 2^{\Delta_{Y}}$ the Hausdorff-metric topology instead of Vietoris. Take $Y=\{0,1\}$ so that $\Delta_{Y}=[0,1]$ via $p\mapsto P(\{1\})=p$, and consider the event $A=\{1\}$. Define the sequence of credal sets
    $\mathcal{M}_{k}=[0,1-\tfrac{1}{k}],\qquad\mathcal{M}_{\infty}=\{1\}.$
    The Hausdorff distance $d_{H}(\mathcal{M}_{k},\mathcal{M}_{\infty})=1/k\to 0$, so $\mathcal{M}_{k}\to\mathcal{M}_{\infty}$ in the Hausdorff topology, but
[…]
    Thus $\mathcal{M}\mapsto\underline{P}_{\!\mathcal{M}}(A)$ is not upper semicontinuous, so $\mathsf{IHDR}$ fails to be u.h.c.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). Therefore, (i)-(iv), together with consonance, are collectively minimal.
An in-depth study of some of the properties of $\mathbf{UHCont}$ is carried out in Section 7. In the next section, instead, we show that a result similar to Theorem 5 can be derived also for a category whose morphisms are weakly measurable correspondences.
4. Conformal Prediction as a Weakly measurable Morphism
This section introduces category $\mathbf{WMeas}_{\text{uc}}$, and shows that, under minimal assumptions, the (Full) Conformal Prediction correspondence $\kappa$ is a morphism of $\mathbf{WMeas}_{\text{uc}}$ as well, embedded in a commuting diagram. The following too is inspired by Perrone, (2024, Example 1.1.14).
Definition 6 (Structure $\mathbf{WMeas}_{\text{uc}}$).
Define a structure $\mathbf{WMeas}_{\text{uc}}$ as follows,
  • Objects: Measurable Polish spaces $(X,\Sigma_{X})$, with $\Sigma_{X}=\mathcal{B}(X)$, the Borel $\sigma$-algebra.
  • Morphisms: Weakly measurable (see Definition 27 in Appendix A), uniformly compact-valued (see Definition 28 in Appendix A) correspondences $\Phi:X\rightrightarrows Y$.
Theorem 7 ($\mathbf{WMeas}_{\text{uc}}$ is a category).
The structure $\mathbf{WMeas}_{\text{uc}}$ defines a category.
Proof.
We verify the four axioms.
(1) and (4) Identity morphisms and Identity laws. For any object $X$, $\operatorname{id}_{X}(x)=\{x\}$ is non-empty compact. Put $K_{\operatorname{id}_{X}}=X$ (compact because $X$ is Polish and we may take the closed unit ball in its metric); The uniformly compact-valued condition holds. For every open $O\subseteq X$ we have $\operatorname{id}_{X}^{-1}(O)=O\in\Sigma_{X}$, so $\operatorname{id}_{X}$ is a morphism. The usual set-theoretic equalities $\Phi\circ\operatorname{id}_{X}=\Phi$ and $\operatorname{id}_{Y}\circ\Phi=\Phi$ prove the identity laws.
(2) Associativity. For morphisms $\Phi\colon X\rightrightarrows Y$, $\Psi\colon Y\rightrightarrows Z$ and $\Theta\colon Z\rightrightarrows W$, the associativity of union gives $\bigl{(}(\Theta\!\circ\!\Psi)\circ\Phi\bigr{)}(x)=(\Theta\!\circ\!(\Psi\!\circ\!\Phi))(x)$, for every $x\in X$.
(3) Closure under composition. Let $\Phi\colon X\rightrightarrows Y$ and $\Psi\colon Y\rightrightarrows Z$ be morphisms. Write $K_{\Phi}\subseteq Y$ and $K_{\Psi}\subseteq Z$ for the uniform compact bounds, see Definition 28.(ii).
Compactness of fibers. For each $x\in X$,
$(\Psi\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\Psi(y)\subseteq\bigcup_{y\in\Phi(x)}K_{\Psi}\subseteq K_{\Psi}.$
Since $K_{\Psi}$ is compact and the union is taken inside it, $(\Psi\circ\Phi)(x)$ is a closed subset of a compact space, hence compact. Thus compactness holds with the same compact bound $K_{\Psi}$, so the uniformly compact-valued condition also holds.
Weak measurability. Fix an open $O\subseteq Z$ and define $H\coloneqq\Psi\circ\Phi$. Because $\Phi$ has non-empty closed values, the Kuratowski-Ryll-Nardzewski theorem (Aliprantis and Border,, 2006, Theorem 18.13) yields Borel selectors $(f_{n})_{n\in\mathbb{N}}\subseteq\mathcal{B}(X,Y)$ with $f_{n}(x)\in\Phi(x)$ and $\overline{\{f_{n}(x):n\in\mathbb{N}\}}=\Phi(x)$. Then
$H(x)\cap O\neq\emptyset\iff\exists n\in\mathbb{N}:\Psi(f_{n}(x))\cap O\neq\emptyset,$
so
$H^{-1}(O)=\bigcup_{n\in\mathbb{N}}f_{n}^{-1}\!\bigl{(}\Psi^{-1}(O)\bigr{)}.$
Each $f_{n}$ is Borel and $\Psi$ is weakly measurable, whence $\Psi^{-1}(O)\in\Sigma_{Y}$ and $f_{n}^{-1}\!\bigl{(}\Psi^{-1}(O)\bigr{)}\in\Sigma_{X}$. A countable union of Borel sets is Borel, so $H^{-1}(O)\in\Sigma_{X}$. Therefore $H$ is weakly measurable, completing closure.
Therefore, $\mathbf{WMeas}_{\text{uc}}$ satisfies the axioms of a category. ∎
Similarly to what we did in Section 3, we now ask ourselves what are the minimal conditions that make correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ uniformly compact-valued and weakly measurable. As we shall see in Theorem 8, they are the same as in Section 3, with (i) and (ii) slightly strengthened to
  1. (i’)
    $Y$ is compact metric and $\Sigma_{Y}$ is its Borel $\sigma$-field.
  2. (ii’)
    Each $\psi\in\mathscr{F}$ is jointly continuous on $Y^{n}\times Y$; $\mathscr{F}$ is Polish in the uniform ($\sup$-norm) topology.
In Remark 2, we show that (i’) is needed to make $\mathsf{IHDR}$ weakly measurable, and (ii’) to make $Y^{n}\times\mathscr{F}$ Polish. We are now ready for the main result of this section.
Theorem 8 (The Conformal Prediction Diagram Commutes, $\mathbf{WMeas}_{\text{uc}}$).
Assume that (i’), (ii’), (iii), (iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram
${Y^{n}\times\mathscr{F}}$${\mathscr{C}}$${\Sigma_{Y}}$$\scriptstyle{\mathsf{CRED}}$$\scriptstyle{\kappa_{\alpha}}$$\scriptstyle{\mathsf{IHDR}_{\alpha}}$
commutes in $\mathbf{WMeas}_{\text{uc}}$.
Notice that, in the Full CP Diagram, $\Sigma_{Y}=\mathcal{B}(Y)$ by Definition 6.
Proof of Theorem 8.
Fix any $n\in\mathbb{N}$. We first show that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are uniformly compact-valued and weakly measurable.
Uniformly compact-valued. All three images lie inside the fixed compact spaces $Y$ (by (i’)) or $\Delta_{Y}$ (by (iv)); We are now left to show that they themselves are compact-valued.
Notice that, under consonance, for all $y^{n}\in Y^{n}$ and all $\psi\in\mathscr{F}$, $\mathsf{CRED}(y^{n},\psi)$ is a credal set, so a closed subset of the compact space $\Delta_{Y}$, and so compact in turn.
Also, for all $\alpha\in[0,1]\setminus S_{n+1}$ and all credal sets $\mathcal{M}$, $\mathsf{IHDR}(\alpha,\mathcal{M})$ is an intersection of closed sets, and so closed itself. In turn, it is a closed subset of the compact space $Y$, so it is itself compact.
An argument similar to point (i) in the Proof of Theorem 5 shows that, for all $y^{n}\in Y^{n}$, all $\psi\in\mathscr{F}$, and all $\alpha\in[0,1]\setminus S_{n+1}$, the set $\kappa(\alpha,y^{n},\psi)=\{y:\pi(y,y^{n},\psi)>\alpha\}=\{y:\pi(y,y^{n},\psi)\geq\beta\}$ is closed, where $\beta\equiv\beta(\alpha)$ was defined in (8). Indeed, it is a super-level set of $\pi(\cdot,y^{n},\psi)$, and the latter is upper semicontinuous. This is because $\{y:\psi(y^{n+1}_{-i},y_{i})\geq\psi(y^{n+1}_{-(n+1)},y_{n+1})\}$ is a closed set, so $\mathbbm{1}[\{y:\psi(y^{n+1}_{-i},y_{i})\geq\psi(y^{n+1}_{-(n+1)},y_{n+1})\}]$ is upper semicontinuous, and the latter is preserved by finite averages. Thus, $\kappa(\alpha,y^{n},\psi)$ is a closed subset of the compact space $Y$, so it is itself compact.
Weak measurability of $\kappa$. Let $\Lambda\coloneqq[0,1]\setminus S_{n+1}$. For every non-empty open set $G\subseteq Y$, put
$f_{G}(\alpha,y^{n},\psi)\;\coloneqq\;\sup_{y\in G}\pi(y,y^{n},\psi),\quad(\alpha,y^{n},\psi)\in\Lambda\times Y^{n}\times\mathscr{F}.$
Since $\pi(\cdot,y^{n},\psi)$ is upper semicontinuous and $G$ is open in the compact space $Y$, $f_{G}$ is itself upper semicontinuous and therefore Borel.
Because $\pi$ only takes values in $S_{n+1}$,
$\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\quad\Longleftrightarrow\quad f_{G}(\alpha,y^{n},\psi)\;\geq\;\beta(\alpha).$
The map $\beta:\Lambda\to S_{n+1}$ is piecewise constant (hence Borel). Set
$\Theta_{G}\coloneqq\bigl{\{}(\alpha,y^{n},\psi)\in\Lambda\times Y^{n}\times\mathscr{F}:f_{G}(\alpha,y^{n},\psi)\geq\beta(\alpha)\bigr{\}}.$
Because $f_{G}$ is Borel and $\beta$ is Borel, $\Theta_{G}$ is a Borel subset of the product space. Therefore
$\bigl{\{}(\alpha,y^{n},\psi):\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\bigr{\}}\;=\;\Theta_{G}\in\mathcal{B}\bigl{(}\Lambda\times Y^{n}\times\mathscr{F}\bigr{)}.$
In turn, for every open $G\subseteq Y$, the selector set $\{\kappa(\cdot)\cap G\neq\emptyset\}$ is measurable, hence the correspondence
$\kappa:\;\Lambda\times Y^{n}\times\mathscr{F}\rightrightarrows Y$
is weakly measurable (Aliprantis and Border,, 2006, Definition 18.1).
Weak measurability of $\mathsf{CRED}$. Let $G\subseteq\Delta_{Y}$ be open. Because $\Delta_{Y}$ is metric and $\mathsf{CRED}$ has compact values, $\mathsf{CRED}(y^{n},\psi)\cap G\neq\emptyset$ if and only if there exists $P\in G$ satisfying all the constraints $P(A_{j})\leq\overline{\Pi}_{(y^{n},\psi)}(A_{j})$, where the $A_{j}$’s are the closed sets in $\Sigma_{Y}$. For each integer $m$, let $G_{m}=\{P\in G:d(P,G^{c})\geq 1/m\}$; $\{G_{m}\}$ is a countable basis for $G$. Hence
$\bigl{\{}(y^{n},\psi):\mathsf{CRED}(y^{n},\psi)\cap G\neq\emptyset\bigr{\}}\;=\;\bigcup_{m=1}^{\infty}\bigcup_{P\in\mathscr{D}_{m}}\bigl{\{}(y^{n},\psi):P(A_{j})\leq\overline{\Pi}_{(y^{n},\psi)}(A_{j}),\;\forall j\bigr{\}},$
where $\mathscr{D}_{m}$ is a countable dense subset of $G_{m}$ (using separability of $\Delta_{Y}$). Each inner set is Borel by Lemma 4; Countable unions preserve measurability, so the whole pre-image is in $\mathcal{B}\bigl{(}Y^{n}\!\times\!\mathscr{F}\bigr{)}$.
Weak measurability of $\mathsf{IHDR}$. For open $G\subseteq Y$, let =$\mathscr{B}=\{B_{k}\}_{k\in\mathbb{N}}$ be a countable closed basis of $Y$. Define $\phi_{k}(\alpha,\mathcal{M})=1-\alpha-\underline{P}_{\!\mathcal{M}}(B_{k}).$ Each $\phi_{k}$ is u.s.c. (by (iv), since $\mathscr{C}$ is endowed with the Vietoris topology and $\mathcal{M}$ is weak-compact), hence Borel on $[0,1]\times\mathscr{C}$. Because $\mathsf{IHDR}(\alpha,\mathcal{M})\cap G\neq\emptyset\iff\exists k:B_{k}\subseteq G,\;\phi_{k}(\alpha,\mathcal{M})\leq 0,$ the desired pre-image is a countable union of Borel sets, hence Borel.
We have now showed that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are uniformly compact-valued and weakly measurable, and thus morphisms of $\mathbf{WMeas}_{\text{uc}}$. Fix now any $\alpha\in[0,1]\setminus S_{n+1}$. It is easy to see that the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ remain uniformly compact-valued and weakly measurable. In addition, as in the Proof of Theorem 5, we have that $t(\mathsf{CRED})=s(\mathsf{IHDR}_{\alpha})=\mathscr{C}$.
Now, recall that the product of two Polish spaces is Polish. Since $Y$ is compact metric by (i’), it is also measurable Polish; $[0,1]\times\mathscr{C}$ is measurable Polish in the product topology by (iv); $Y^{n}\times\mathscr{F}$ is measurable Polish in the product topology by (i’) and (ii’); $[0,1]\times Y^{n}\times\mathscr{F}$ is measurable Polish in the product topology by (i’) and (ii’). Hence, these spaces are all objects of $\mathbf{WMeas}_{\text{uc}}$.
The fact that the Full Conformal Prediction Diagram commutes in $\mathbf{WMeas}_{\text{uc}}$, then, is an immediate consequence of Caprio et al., 2025a (, Proposition 5). ∎
The same considerations that we put forth in Section 3 on the importance of Theorem 5, still hold here for Theorem 8.
Theorems 5 and 8 tell us that the Full Conformal Prediction Diagram commutes when we focus either on the continuity or on the measurability aspects of the conformal prediction correspondence $\kappa$. This further testifies that, under our assumptions, uncertainty quantification is a built-in feature of Conformal Prediction methods.
The following remark shows that the conditions in Theorem 8 are truly minimal.
Remark 2 (Minimality of Hypotheses (i’), (ii’), (iii), (iv)).
Similarly to Remark 1, we prove that each hypothesis is indispensable for the simultaneous uniformly compact-valuedness and weak measurability of the three correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$.
  • Drop (i’) (compactness of $Y$). Let $Y=\mathbb{R}$, $n=1$, $\psi(y^{1},y)=-y$, $\alpha=\frac{3}{4}$. Then $\kappa(\alpha,y^{1},\psi)=[y^{1},\infty)$ is not compact; Uniform compact-valuedness fails.
  • Drop (ii’) (Polishness of $\mathscr{F}$).
[…]
    (b) Continuity. If continuity is also abandoned (e.g. $\psi(y^{1},y)=\mathbbm{1}[y\in\mathbb{Q}]$), $\kappa$’s graph is not Borel, so weak measurability fails.
  • Drop (iii) (no-tie rule). Pick $\alpha=k/(n+1)$, $k\in\{1,\ldots,n+1\}$. Then, $\kappa(\alpha,y^{n},\psi)$ is open but not closed, so individual values are not compact, and uniform compactness is lost.
  • Drop (iv) (weak/Vietoris topologies).
    (i) Put total-variation on $\Delta_{Y}$. Then $P\mapsto P(A)$ is not continuous; Constraints in $\mathsf{CRED}$ are non-Borel, so weak measurability fails.
    (ii) Replace Vietoris with the Hausdorff metric on $\mathscr{C}$. Credal-set limits can break lower-probability continuity; $\mathsf{IHDR}$ fails to be weakly measurable, see Remark 1.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). Hence (i’), (ii’), (iii), (iv), together with consonance, are jointly minimal for the uniform compact-valuedness and weak measurability of $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$.
5. Unifying Bayesian, Frequentist, and Imprecise Reasoning
In this section, we show how a diagram involving Bayesian, Conformal, and Imprecise Probabilistic prediction methods, commute in category $\mathbf{UHCont}$. We focus on the latter rather than $\mathbf{WMeas}_{\text{uc}}$ because its objects and morphisms have less restricting requirements.
Since the diagram commuting is an extension of the Full CP Diagram in Theorem 5, we retain assumptions (i)-(iv).
Let $(\Theta,\Sigma_{\Theta})$ be a measurable parameter space, and call $\lambda_{\Theta}$ a fixed $\sigma$-finite Borel measure on $(\Theta,\Sigma_{\Theta})$. Call $\Delta_{\Theta}^{\mathrm{dens}}$ the space of finitely additive probability measures on $\Theta$ admitting a density with respect to $\lambda_{\Theta}$ (Bell and Hagood,, 1988). [^9] This requirement essentially restricts us to focus on countably additive probabilities, see e.g. also here. Endow $\Delta_{\Theta}^{\mathrm{dens}}$ with the $L^{1}(\lambda_{\Theta})$ topology. Consider now the following conditions,
  1. (v)
    Measurability. There exists a de Finetti kernel $k_{n}:\Theta\times Y^{n}\rightarrow\mathbb{R}_{+}$ such that, under parameter $\theta$, the joint density of $y^{n}$ is $p(y^{n}\mid\theta)=k_{n}(\theta,y^{n})$ (Hoff,, 2009), and the predictive for $y_{n+1}$ is
    $p_{P}(y_{n+1}\mid y^{n})=\frac{\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))P(\text{d}\theta)}{\int_{\Theta}k_{n}(\theta,y^{n})P(\text{d}\theta)},$
    where $P\in\Delta^{\text{dens}}_{\Theta}$. We assume that $k_{n}(\theta,y^{n})$ is $\Sigma_{\Theta}$-measurable in $\theta$ for each $y^{n}\in Y^{n}$, and Borel in $y^{n}$ for each $\theta$.
  2. (vi)
    Pointwise $y$-continuity a.e. For $\lambda_{\Theta}$-almost all $\theta$, the map $y^{n}\mapsto k_{n}(\theta,y^{n})$ is continuous on $Y^{n}$.
  3. (vii)
    Integrable envelope. There exists $M(\theta)\equiv M\in L^{1}(\lambda_{\Theta})$ with $k_{n+1}(\theta,(y^{n},y_{n+1}))\;\leq\;M(\theta)$ for every $(y^{n},y_{n+1})\in Y^{n+1}$ and $\lambda_{\Theta}$-almost all $\theta$.
  4. (viii)
    Prior integrability. Every admissible prior probability measure $P\in\Delta_{\Theta}^{\mathrm{dens}}$ satisfies $\int_{\Theta}M(\theta)P(\text{d}\theta)<\infty.$ That is, calling $p=\text{d}P/\text{d}\lambda_{\Theta}$ the Radon-Nykodim derivative of $P$ with respect to $\lambda_{\Theta}$, $\int_{\Theta}p(\theta)\,M(\theta)\lambda_{\Theta}(\text{d}\theta)\;<\;\infty$.
  5. (ix)
    Positive marginal likelihood. For every $y^{n}\in Y^{n}$ and every admissible prior probability measure $P\in\Delta_{\Theta}^{\mathrm{dens}}$,
    $D_{P}(y^{n})\coloneqq\int_{\Theta}k_{n}(\theta,y^{n})P(\text{d}\theta)=\int_{\Theta}k_{n}(\theta,y^{n})p(\theta)\lambda_{\Theta}(\text{d}\theta)\;>\;0.$
For $P\in\Delta_{\Theta}^{\mathrm{dens}}$ having density $p$, and data $y^{n}\in Y^{n}$, set
$p_{P}\left(y_{n+1}\mid y^{n}\right)\coloneqq\frac{N_{P}(y_{n+1},y^{n})}{D_{P}(y^{n})},$
where $N_{P}(y_{n+1},y^{n})\coloneqq\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))p(\theta)\lambda_{\Theta}(\text{d}\theta)$. By (ix), it is well-defined.
Because $y\mapsto p_{P}(y\mid y^{n})$ is continuous (we prove this in Lemma 9) and $Y^{n+1}$ is compact by (i), the natural codomain for such a map is the Banach space $C(Y^{n}\times Y)$ with norm $\|f\|_{\infty}\coloneqq\sup_{(y^{n},y)\in Y^{n}\times Y}|f(y^{n},y)|$. Define
$\mathsf{BCP}:\;Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}\rightarrow Y^{n}\times C(Y^{n}\times Y),\quad\mathsf{BCP}\bigl{(}y^{n},P\bigr{)}\coloneqq\bigl{(}y^{n},\psi_{(y^{n},P)}\bigr{)},$
where
$\psi_{(y^{n},P)}(y_{n+1})\coloneqq-\,p_{P}(y_{n+1}\mid y^{n});$
We call it $\mathsf{BCP}$ because it is evocative of an acronym for Bayesian Conformal Prediction (Fong and Holmes,, 2021). Next, we have an important lemma.
Lemma 9.
Under (v)-(ix), the following hold for every fixed $n\in\mathbb{N}$,
  • For every $P\in\Delta_{\Theta}^{\mathrm{dens}}$, the predictive map $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ is jointly continuous on $Y^{n}\times Y=Y^{n+1}$;
  • The map $\mathsf{BCP}$ is continuous when $C(Y^{n}\times Y)\equiv C(Y^{n+1})$ is equipped with the sup-norm.
Proof.
Throughout, let $P\in\Delta_{\Theta}^{\mathrm{dens}}$ be fixed and write $\text{d}P/\text{d}\lambda_{\Theta}=p$.
Step 1. Continuity of the numerator $N_{P}$. Define
$g(\theta,y_{n+1},y^{n})\coloneqq k_{n+1}(\theta,(y^{n},y_{n+1}))p(\theta);$
We have $\lvert g\rvert\leq p(\theta)M(\theta)$ by (vii). Assumption (vi) gives continuity of $(y_{n+1},y^{n})\mapsto g(\theta,y_{n+1},y^{n})$ for $\lambda_{\Theta}$-almost all $\theta$. The envelope $pM^{n+1}\in L^{1}(\lambda_{\Theta})$ by (viii); Hence, dominated convergence yields continuity of $(y_{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$.
Step 2. Continuity and positivity of the denominator $D_{P}$. The integrand in $D_{P}(y^{n})$ is $g$, but “stopped” at observation $n$, so the same argument as Step 1 shows that $y^{n}\mapsto D_{P}(y^{n})$ is continuous. Positivity follows from (ix).
Step 3. Joint (and uniform) continuity of the predictive. Because $Y^{n+1}$ is compact by (i) and $D_{P}$ never vanishes,
$p_{P}(y_{n+1}\mid y^{n})=\frac{N_{P}(y_{n+1},y^{n})}{D_{P}(y^{n})}$
is continuous in $(y_{n+1},y^{n})$ and, by compactness, uniformly continuous. This proves joint continuity of the map $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$.
Step 4. Continuity in the prior. Let $P_{m}\to P$ in $L^{1}(\lambda_{\Theta})$, with densities $p_{m}\coloneqq\text{d}P_{m}/\text{d}\lambda_{\Theta}$. From (vii)-(viii)
$\bigl{\|}N_{P_{m}}-N_{P}\bigr{\|}_{\infty}\leq\int_{\Theta}\lvert p_{m}-p\rvert\,M\,\text{d}\lambda_{\Theta}\xrightarrow[m\to\infty]{}0,$
and similarly $\|D_{P_{m}}-D_{P}\|_{\infty}\to 0.$ Since $D_{P}$ is bounded below (by Step 2) and $D_{P_{m}}$ converges uniformly,
$\|p_{P_{m}}(\cdot\mid y^{n})-p_{P}(\cdot\mid y^{n})\|_{\infty}\xrightarrow[m\to\infty]{}0\quad\text{uniformly in }y^{n}.$
Step 5. Continuity of $\mathsf{BCP}$. Fix a point $(y^{n},P)$ in the domain. If $y^{n}_{m}\to y^{n}\text{ and }P_{m}\to P$ (as in Step 4), then
$\|\psi_{(y^{n}_{m},P_{m})}-\psi_{(y^{n},P)}\|_{\infty}\leq\|p_{P_{m}}(\cdot\mid y^{n}_{m})-p_{P_{m}}(\cdot\mid y^{n})\|_{\infty}+\|p_{P_{m}}(\cdot\mid y^{n})-p_{P}(\cdot\mid y^{n})\|_{\infty}.$
The first term vanishes by uniform continuity from Step 3, the second by Step 4. Hence, $\|\psi_{(y^{n}_{m},P_{m})}-\psi_{(y^{n},P)}\|_{\infty}\to 0$, and the first component $y^{n}_{m}\to y^{n}$ by assumption. Therefore, $\mathsf{BCP}$ is continuous. ∎
Conditions (v)-(ix) are indispensable for Lemma 9 to hold. This is examined in the following remark.
Remark 3 (Minimality of (v)-(ix)).
The five conditions (v)-(ix) are truly minimal for Lemma 9 to hold. If any single hypothesis is dropped, one can construct a Bayesian model for which $\mathsf{BCP}$ fails to be continuous (or even well-defined). To see this, consider the following.
  • Drop (v) (Measurability of $k$). Without measurability in $\theta$, the integrals defining $N_{P}$ and $D_{P}$ need not exist as Lebesgue integrals.
  • Drop (vi) ($y$-continuity a.e.). If $k_{n}(\theta,\cdot)$ has a jump in $y$, then $D_{P}$—and hence $p_{P}$—can jump as a function of $y$, breaking continuity.
[…]
  • Drop (viii) (Prior integrability). If $\int_{\Theta}M(\theta)P(\text{d}\theta)=\infty$ for some $P$, then $N_{P}$ or $D_{P}$ may diverge or lose continuity.
  • Drop (ix) (Positive marginal likelihood). If $D_{P}(y^{n})=0$ for some $y^{n}$, the posterior density $p_{P}$ is undefined there, so $\mathsf{BCP}$ cannot be continuous.
In practice, conditions (v)-(ix) are almost built-in to any reasonable Bayesian model. Let us briefly expand on this. First, notice that almost every statistical model (Gaussian, exponential family, mixture, etc.) has a likelihood density continuous and measurable in $y$, so that (v) is met.
Condition (vi) guarantees that, as we vary our data $y^{n}$, the integrand in $N_{P}$ and $D_{P}$ moves continuously, so dominated convergence applies. In practice, standard density kernels (Normal, Laplace, Student-$t$, etc.) are smooth in $y$.
The uniform-in-$y$ bound of (vii) is the key to swapping limits and integrals. If $Y$ is compact (as it is in our setting, by (i)) and $k_{n}(\theta,\cdot)$ is continuous, we can often take $M(\theta)=\sup_{y\in Y}k_{n}(\theta,y)$, which is finite and $L^{1}$-integrable under mild assumptions on $k_{n}$.
Since $N_{P}$ involves $k_{n+1}$, the moment condition in (viii) ensures finiteness and continuity of the predictive density. Common priors (Gaussian, Gamma, Beta, compact support) have finite moments of all orders, so (viii) is automatic.
Condition (ix) prevents division by zero in the posterior $p_{P}(y_{n+1}\mid y^{n})=N_{P}/D_{P}$. In well-posed Bayesian models, the prior’s support overlaps every parameter value that could generate the observed data, so $D_{P}$ never vanishes.
It is safe to say, then, that hypotheses (v)-(ix) are the minimal technical machinery needed to ensure the posterior predictive is well-defined and varies continuously in both the data $y^{n}$ and the prior $P$. In practical modeling, they hold almost automatically.
Before the main result of this section, we need to introduce a further condition on the posterior predictive Cumulative Density Function (CDF).
  1. (x)
    Well-behaved CDF. For each fixed $(y^{n},P)\in Y^{n}\times\Delta_{\Theta}^{\text{dens}}$, the CDF
    $F(c;\,y^{n},P)\coloneqq\int_{\{\,p_{P}(\cdot\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\text{d}y$
    is continuous and strictly increasing in a neighborhood of its $(1-\alpha)$-quantile
    $c(\alpha,y^{n},P)=\inf\{\,c:F(c;y^{n},P)\geq 1-\alpha\}.$
Call then $\mathsf{QUANT}:[0,1]\times Y^{n}\times\Delta_{\Theta}^{\text{dens}}\rightrightarrows Y$ the set-valued map that extracts the $\alpha$-level set of the posterior predictive distribution. That is,
$(\alpha,y^{n},P)\mapsto\mathsf{QUANT}(\alpha,y^{n},P)\coloneqq H\bigl{(}c(\alpha,y^{n},P),\,y^{n},P\bigr{)},$
where
$H:\mathbb{R}\times Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}\;\rightrightarrows\;Y,\qquad H\bigl{(}c,y^{n},P\bigr{)}\coloneqq\{\,y\in Y:p_{P}(y\mid y^{n})\geq c\}.$
Lemma 10.
Under assumptions (i)-(x), the correspondence $\mathsf{QUANT}$ is upper hemicontinuous.
Proof.
First, we show that under (i)-(x), the correspondence $H$ is upper hemicontinuous. Fix $(c_{0},y^{n}_{0},P_{0})$ and let $V\supset H(c_{0},y^{n}_{0},P_{0})$ be open in $Y$. Since $p_{P}(y\mid y^{n})$ is jointly continuous and $Y$ is compact, $H(c_{0},y^{n}_{0},P_{0})=\{y:p_{P_{0}}(y\mid y^{n}_{0})\geq c_{0}\}$ is closed and contained in $V$. Thus there is $\varepsilon>0$ such that
$\bigl{\{}y:p_{P_{0}}(y\mid y^{n}_{0})\geq c_{0}-\varepsilon\bigr{\}}\;\subset\;V.$
By joint continuity of $(c,y^{n},P,y)\mapsto p_{P}(y\mid y^{n})$, there is a neighborhood $U$ of $(c_{0},y^{n}_{0},P_{0})$ in ×$\mathbb{R}\times Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}$ such that for all $(c,y^{n},P)\in U$ and all $y\in Y$,
$p_{P}(y\mid y^{n})\;\geq\;c\quad\Longrightarrow\quad p_{P_{0}}(y\mid y^{n}_{0})\;\geq\;c_{0}-\varepsilon.$
Hence for every $(c,y^{n},P)\in U$,
$H(c,y^{n},P)=\{y:p_{P}(y\mid y^{n})\geq c\}\;\subset\;\{y:p_{P_{0}}(y\mid y^{n}_{0})\geq c_{0}-\varepsilon\}\;\subset\;V,$
proving upper hemicontinuity.
Now, we show that under (i)-(x), function $c$ is continuous. By (x), for each fixed $(y^{n},P)$ the function $c\mapsto F(c;y^{n},P)$ is continuous and strictly increasing near the target quantile. Standard results on the continuity of inverse functions then imply that
$(\alpha,y^{n},P)\;\longmapsto\;c(\alpha,y^{n},P)$
is continuous at every point, since it solves $F\bigl{(}c(\alpha,y^{n},P);y^{n},P\bigr{)}=1-\alpha$ uniquely in $c$, and both the left side and the right side vary continuously in $(\alpha,y^{n},P)$.
Finally, we put everything together to show the upper hemicontinuity of $\mathsf{QUANT}$. View $\mathsf{QUANT}$ as the composition
[…]
As we showed before, $H$ is upper hemicontinuous in its three arguments, and $c(\alpha,y^{n},P)$ is continuous, hence he singleton-valued correspondence associated with $c$ (that we still denote by $c$ for ease of notation) is upper hemicontinuous. The composition of two upper hemicontinuous correspondences is again upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Therefore, $\mathsf{QUANT}$ is upper hemicontinuous. ∎
Let us add a brief discussion on condition (x).
Remark 4 (On the Well-Behavior of the Predictive CDF).
In many continuous Bayesian models, $p_{P}(y\mid y^{n})$ is a continuous, strictly positive function on compact $Y$ (for us, this holds by (i)). As $c$ varies, the level-sets $\{y:p_{P}(y\mid y^{n})\leq c\}$ move continuously, so $F(c)$ has no jumps. Since $p_{P}$ has no plateaus of positive volume, for any $c<c^{\prime}$, the set $\{c<p_{P}(y\mid y^{n})\leq c^{\prime}\}$ has strictly positive mass, forcing $F(c^{\prime})-F(c)>0$. Hence, $F(c)$ is both continuous and strictly increasing at the target quantile.
The condition may fail in two cases, rather negligible in practice. If $p_{P}(y\mid y^{n})\equiv c$ on a subset of $Y$ of positive measure, then $F(c)$ is constant across that plateau, violating strict increase. Also, if the predictive has atoms (point-mass components), $F(c)$ jumps at those mass points and fails to be continuous.
Lemmas 9 and 10 are crucial in light of the following, which is the main result of this section.
Theorem 11 (Unifying Bayes, Conformal, and Imprecise Prediction in $\mathbf{UHCont}$).
Assume that (i)-(x) hold. Consider the restrictions $\kappa_{\alpha}$, $\mathsf{IHDR}_{\alpha}$, and $\mathsf{QUANT}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, the following diagram
${Y^{n}\times\mathscr{F}}$${\mathscr{C}}$${Y^{n}\times\Delta^{\text{dens}}_{\Theta}}$${\Sigma_{Y}}$$\scriptstyle{\mathsf{CRED}}$$\scriptstyle{\kappa_{\alpha}}$$\scriptstyle{\mathsf{IHDR}_{\alpha}}$$\scriptstyle{\mathsf{BCP}}$$\scriptstyle{\mathsf{QUANT}_{\alpha}}$
commutes in $\mathbf{UHCont}$, that is, $\mathsf{QUANT}_{\alpha}=\kappa_{\alpha}\circ\mathsf{BCP}=\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}\circ\mathsf{BCP}$.
Proof.
First, notice that we do not require consonance directly because it is immediately satisfied as a result of (i) and $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ being jointly continuous on $Y^{n}\times Y=Y^{n+1}$.
Notice also that, by (i) and $\Delta^{\text{dens}}_{\Theta}$ being endowed the $L^{1}(\lambda_{\Theta})$ topology, $Y^{n}\times\Delta^{\text{dens}}_{\Theta}$ is a topological space in the product topology, hence a well-defined object of $\mathbf{UHCont}$.
In addition, by Lemma 9, function $\mathsf{BCP}$ is continuous. The latter implies that the singleton-valued correspondence associated with $\mathsf{BCP}$ (that we still denote by $\mathsf{BCP}$ for ease of notation) is upper hemicontinuous. Moreover, by Lemma 10, correspondence $\mathsf{QUANT}$ (and also, trivially, its restriction $\mathsf{QUANT}_{\alpha}$ to some $\alpha$ satisfying (iii)) is upper hemicontinuous as well. In turn, they are well-defined morphisms of $\mathbf{UHCont}$.
Furthermore, we have that $t(\mathsf{BCP})=s(\kappa_{\alpha})=s(\mathsf{CRED})=Y^{n}\times\mathscr{F}$. That is, the target space of $\mathsf{BCP}$ coincides with the source spaces of both $\kappa_{\alpha}$ and $\mathsf{CRED}$. This is a consequence of
  • Fong and Holmes, (2021), who show that the negative posterior predictive density is a valid non-conformity score (i.e. it is an element of $\mathscr{F}$ in (4));
  • Lemma 9, which shows that map $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ is jointly continuous on $Y^{n}\times Y=Y^{n+1}$ (and so also $\psi_{(y^{n},P)}(y^{n+1})$ is continuous, since we simply multiply $p_{P}(y_{n+1}\mid y^{n})$ by $-1$). This shows that $\psi_{(y^{n},P)}(y^{n+1})\equiv\psi_{P}(y^{n},y^{n+1})$ satisfies (ii).
Now, given our choice of non-conformity score, the Conformal Prediction Region, that is, the image $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ of the Conformal correspondence $\kappa_{\alpha}$, coincides with the $\alpha$-level set $\mathsf{QUANT}_{\alpha}(y^{n},P)$ of the posterior predictive distribution, for all $\alpha$ that satisfies (iii). To see this, notice that, by definition, a candidate $y\in Y$ lies in $\mathscr{R}^{\psi_{(y^{n},P)}}_{\alpha}(y^{n})\equiv\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ if and only if
|$|\{\,i:\psi_{(y^{n},P)}(y_{i})\geq\psi_{(y^{n},P)}(y_{n+1})\}|=|\{\,i:p_{P}(y_{i}\mid y^{n})\leq p_{P}(y_{n+1}\mid y^{n})\}|>\;k\coloneqq\lceil(n+1)\alpha\rceil.$
But among the $n$ scores $\psi_{(y^{n},P)}(y_{1}),\ldots,\psi_{(y^{n},P)}(y_{n})$, the event |$|\{i:\psi_{(y^{n},P)}(y_{i})\geq s\}|>k$ holds precisely when $s\leq\psi_{(y^{n},P)}(y_{(k)})$, where $\{\psi_{(y^{n},P)}(y_{(1)}),\ldots,\psi_{(y^{n},P)}(y_{(n)})\}$ are the ordered values of $\{\psi_{(y^{n},P)}(y_{1}),\ldots,\psi_{(y^{n},P)}(y_{n})\}$. Therefore,
$y\in\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})\quad\iff\quad\psi_{(y^{n},P)}(y_{n+1})\;\leq\;\psi_{(y^{n},P)}(y_{(k)}).$
Set $c(\alpha,y^{n},P)=\psi_{(y^{n},P)}(y_{(k)})$. Then,
$\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\{\,y_{n+1}\in Y:\psi_{(y^{n},P)}(y_{n+1})\leq c(\alpha,y^{n},P)\},$
that is, $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ is exactly the level set of the posterior predictive density at level $c(\alpha,y^{n},P)$.
By Caprio et al., 2025a (, Proposition 5), we also know that $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi}))$, for all $\alpha$ that satisfies (iii). The remaining arrows follow directly from Theorem 5. ∎
Theorem 11 is the central payoff of our categorical treatment, and therefore the greatest joy (so far!) of looking at Conformal Prediction from a categorical angle. It confirms the intuition put forth in Martin, (2022) and Caprio et al., 2025a (, Section 5.2), that CP is a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning. In the category UHCont, the model-based Bayesian and imprecise methods, and the model-free conformal construction, yield the same $\alpha$-level prediction region, thus showing a very profound connections between these three only apparently far apart prediction mechanisms.
Notice also that Theorem 11 allows us to tie CP to the concept of e-posterior (Grünwald,, 2023), thus further testifying the central role of CP in modern Machine Learning.
To see this, denote by $\Delta_{Y}^{\text{dens}}$ the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$. Since we observe the realizations $y^{n}=(y_{1},\ldots,y_{n})$ of the first $n$ terms $\mathcal{Y}_{1},\ldots,\mathcal{Y}_{n}$ of an exchangeable process (as pointed out in Section 2.1), the likelihood density $p(y^{n}\mid\theta)\equiv\ell(y^{n}\mid\theta)$ can be written as (de Finetti,, 1974, 1975)
$\ell(y^{n}\mid\theta)=\int\left[\prod_{i=1}^{n}h(y_{i}\mid\varsigma)\right]G_{\theta}(\text{d}\varsigma),$
where $h(\cdot\mid\varsigma)$ is the density of an element $L_{\varsigma}\in\Delta_{Y}^{\text{dens}}$ (given a latent parameter $\varsigma$), and $G_{\theta}$ is a mixing distribution indexed by the parameter $\theta\in\Theta$.
Consider then a credal prior set $\mathcal{P}_{\text{prior}}$, that is, a nonempty convex and weak-closed (and hence also closed in the $L^{1}(\lambda_{\Theta})$ topology) subset of $\Delta_{\Theta}^{\text{dens}}$. Let $\overline{P}$ be its upper probability, and $\overline{p}\coloneqq\sup_{P\in\mathcal{P}_{\text{prior}}}\frac{\text{d}P}{\text{d}\lambda_{\Theta}}$ be the upper prior density; The lower prior density $\underline{p}$ is defined similarly, with $\sup$ replaced by $\inf$.
Proposition 12 (The Upper Posterior is an e-Posterior).
Suppose that $\overline{P}$ is $2$-alternating, that is, $\overline{P}(A_{1}\cup A_{2})+\overline{P}(A_{1}\cap A_{2})\leq\overline{P}(A_{1})+\overline{P}(A_{2})$, for all $A_{1},A_{2}\in\Sigma_{\Theta}$, and that $\sup_{y^{n}\in Y^{n}}\ell(y^{n}\mid\theta)<\infty$, for all $\theta\in\Theta$. Then, the upper posterior density
$\overline{p}(\theta\mid y^{n})=\frac{\ell(y^{n}\mid\theta)\overline{p}(\theta)}{\underline{\ell}(y^{n})},$ (10)
where ¯$\underline{\ell}(y^{n})\coloneqq\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)$, is an e-posterior (Grünwald,, 2023).
Proof.
First, notice that the upper posterior density in (10) is well defined if $\overline{P}$ is $2$-alternating and $\sup_{y^{n}\in Y^{n}}\ell(y^{n}\mid\theta)<\infty$, for all $\theta\in\Theta$, by Wasserman and Kadane, (1990); Caprio et al., 2024b .
Then, recall that an e-posterior is a function $p^{\text{e}}(\theta\mid y^{n})=1/Z_{\theta}(y^{n})$, where $Z_{\theta}(y^{n})$ is an e-variable, that is, a statistic for which $\mathbb{E}_{Y^{n}\sim L_{\theta}}[Z_{\theta}(Y^{n})]\leq 1$. Call $\lambda_{Y}^{\otimes n}$ the $n$-fold product of $\lambda_{Y}$; We have
$\mathbb{E}_{Y^{n}\sim L_{\theta}}\left[\frac{1}{\overline{p}(\theta\mid Y^{n})}\right]$ $=\mathbb{E}_{Y^{n}\sim L_{\theta}}\left[\frac{\underline{\ell}(Y^{n})}{\ell(Y^{n}\mid\theta)\overline{p}(\theta)}\right]$
$=\int_{Y^{n}}\frac{\underline{\ell}(y^{n})}{\ell(y^{n}\mid\theta)\overline{p}(\theta)}\ell(y^{n}\mid\theta)\lambda_{Y}^{\otimes n}(\text{d}y^{n})$
$=\frac{1}{\overline{p}(\theta)}\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})$
$\leq 1\iff\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})\leq\overline{p}(\theta),\quad\forall\theta\in\Theta.$
Now, notice that
$\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})$ $=\int_{Y^{n}}\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\lambda_{Y}^{\otimes n}(\text{d}y^{n})$
$=\int_{\Theta}\underline{p}(\theta)\int_{Y^{n}}\ell(y^{n}\mid\theta)\lambda_{Y}^{\otimes n}(\text{d}y^{n})\lambda_{\Theta}(\mathrm{d}\theta)$
$=\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta),$
where the last-but-one equality comes from Fubini-Tonelli, and the last one comes from $\ell(y^{n}\mid\theta)$ being a proper density. Hence, the upper posterior density $\overline{p}(\theta\mid y^{n})$ is an e-posterior if and only if
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),\quad\forall\theta\in\Theta.$
But this corresponds to the coherence condition in Walley, (1991, Section 2.5), which is always met when $\mathcal{P}_{\text{prior}}\neq\emptyset$. The proof is then concluded. ∎
A consequence of Proposition 12 is that, if
$\overline{\Pi}(A)\equiv\overline{\Pi}_{(y^{n},\psi)}(A)=\int_{A}\underbrace{\int_{\Theta}\ell(y_{n+1}\mid\theta)\overline{p}(\theta\mid y^{n})\lambda_{\Theta}(\text{d}\theta)}_{\eqqcolon\overline{p}(y_{n+1}\mid y^{n})}\lambda_{Y}(\text{d}y),\quad\forall A\in\Sigma_{Y},$ (11)
then the $\alpha$-level Imprecise Highest Density Region $\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi}))$ in Theorem 11 is a prediction region that is associated with an e-posterior. The equality in (11) only holds if the upper posterior predictive density $\overline{p}(y_{n+1}\mid y^{n})$ collapses to a point mass.
Proposition 13 (Coincidence of Upper Probabilities implies Degeneracy).
If (11) holds, then then there exist elements $y^{\ast}\in Y$ and $\theta^{\ast}\in\Theta$ such that
$\overline{p}(y_{n+1}\mid y^{n})=\delta_{y^{\ast}}(y)\quad(\lambda_{Y}\text{-a.s.}),\qquad\overline{p}\!\left(\theta\mid y^{n}\right)=\delta_{\theta^{\ast}}(\theta),\qquad\ell\!\left(y\mid\theta^{\ast}\right)=\delta_{y^{\ast}}(y).$
Proof.
Let $\overline{\Pi}_{\mathrm{B}}(A)\coloneqq\int_{A}\overline{p}(y_{n+1}\mid y^{n})\lambda_{Y}(\text{d}y)$, for all $A\in\Sigma_{Y}$.
Step 1. Maxitivity versus additivity. For disjoint measurable sets $A,B\in\Sigma_{Y}$,
$\overline{\Pi}(A\cup B)=\max\{\overline{\Pi}(A),\overline{\Pi}(B)\},\qquad\overline{\Pi}_{\mathrm{B}}(A\cup B)=\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B),$
because $\overline{\Pi}$ is maxitive while $\overline{\Pi}_{\mathrm{B}}$ is additive. Equality (11) therefore forces
$\max\{\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\}\;=\;\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B).$ (12)
Step 2. A 00-$1$ law for $\overline{\Pi}_{\mathrm{B}}$. If $10<\overline{\Pi}_{\mathrm{B}}(A)<$ 1 and $10<\overline{\Pi}_{\mathrm{B}}(B)<$ 1 with $A\cap B=\emptyset$, then (12) is impossible because $\max\{x,y\}<x+y$ for $x,y\in(0,1)$. Hence for every measurable $A$, it must be that
$\overline{\Pi}_{\mathrm{B}}(A)\in\{0,1\}.$ (13)
Step 3. Concentration on a single point. Define the set $S=\{y\in Y:\overline{p}(y_{n+1}\mid y^{n})>0\}$. If $\lambda_{Y}(S)=0$ then $\overline{\Pi}_{\mathrm{B}}\equiv 0$, contradicting $\overline{\Pi}_{\mathrm{B}}(Y)=1$. Thus $\lambda_{Y}(S)>0$ and (13) with $A=S$ gives $\overline{\Pi}_{\mathrm{B}}(S)=1$. Now take any two disjoint measurable subsets $A,B\subseteq S$ with positive $\lambda_{Y}$-measure. By (13) at most one of them can have $\overline{\Pi}_{\mathrm{B}}=1$, so we must have $\overline{\Pi}_{\mathrm{B}}(A)=\overline{\Pi}_{\mathrm{B}}(B)=0$. This is possible only if every measurable subset of $S$ with positive $\lambda_{Y}$-measure has $\overline{\Pi}_{\mathrm{B}}(\cdot)=1$, which forces $S$ to be a $\lambda_{Y}$-null neighborhood of a single point $y^{\ast}$ (up to $\lambda_{Y}$-null sets). Hence $\overline{\Pi}_{\mathrm{B}}$ (and therefore $\overline{p}(y_{n+1}\mid y^{n})$) is the point mass $\delta_{y^{\ast}}$, proving the first claim.
Step 4. Collapse of the posterior and the likelihood. Because
[…]
the mixture representation can hold only if $\overline{p}(\theta\mid y^{n})=\delta_{\theta^{\ast}}(\theta)$ and $\ell(y\mid\theta^{\ast})=\delta_{y^{\ast}}(y)$. ∎
Although Proposition 13 shows that the equality in (11) holds only under very restrictive conditions, it nevertheless hints at the possibility of an arrow in the diagram of Theorem 11 connecting $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ and $\mathscr{C}$ via an e-posterior-based method. [^10] To be precise, for such an arrow to be a morphism of $\mathbf{UHCont}$, we need to substitute $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ with $Y^{n}\times\mathscr{C}_{\Theta}^{\text{dens}}$, where $\mathscr{C}_{\Theta}^{\text{dens}}$ denotes the set of convex and weak-closed subsets of $\Delta_{\Theta}^{\text{dens}}$. In that case, $\mathsf{BCP}$ and $\mathsf{QUANT}_{\alpha}$ both take as input the pair $(y^{n},\{P\})$, where $\{P\}\subset\mathcal{P}_{\text{prior}}$. A deeper study of the broader interest of the e-posterior approach to Conformal Prediction within a category theory framework is left for future research.
6. Conformal Prediction Regions as Images of the Imprecise Highest Density Region Functor
In this section, we see how the correspondence $\mathsf{IHDR}_{\alpha}$ can be seen as a functor between two well-known categories. As a consequence, the Conformal Prediction Region can be seen as a functor image.
Theorem 14 (Functoriality of the $\,\alpha$-IHDR).
Let $(Y,\Sigma_{Y})$ be a measurable space, and let $\Delta_{Y}$ be the set of all finitely additive probability measures on $\Sigma_{Y}$ endowed with the weak topology. Let
[…]
and regard both $\mathscr{C}$ and $\Sigma_{Y}$ as (pre-)categories whose only morphisms are inclusions $\subseteq$”. Fix $\alpha\in[0,1]$. For $\mathcal{M}\in\mathscr{C}$, define
$B_{\mathcal{M}}^{\alpha}\;\coloneqq\;\bigcap\Bigl{\{}A\in\Sigma_{Y}:\underline{P}_{\mathcal{M}}(A)\,\geq 1-\alpha\Bigr{\}}.$
Set
$\mathsf{IHDR}_{\alpha}:\mathscr{C}\rightarrow\Sigma_{Y},\quad\mathcal{M}\mapsto B_{\mathcal{M}}^{\alpha},\qquad\mathsf{IHDR}_{\alpha}(\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2})\coloneqq B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha}.$
Then, $\mathsf{IHDR}_{\alpha}$ is a covariant functor.
Proof.
(i) Objects. Every $A$ in the defining intersection for $B_{\mathcal{M}}^{\alpha}$ is measurable; Arbitrary intersections of sets in $\Sigma_{Y}$ remain in $\Sigma_{Y}$, hence $B_{\mathcal{M}}^{\alpha}\in\Sigma_{Y}$. Thus $\mathsf{IHDR}_{\alpha}$ maps objects of $\mathscr{C}$ to objects of $\Sigma_{Y}$.
(ii) Morphisms (monotonicity). Suppose $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$. Let $y\in B_{\mathcal{M}_{1}}^{\alpha}$. If $y\notin B_{\mathcal{M}_{2}}^{\alpha}$, then by definition of $B_{\mathcal{M}_{2}}^{\alpha}$, there exists $A\in\Sigma_{Y}$ such that
$\underline{P}_{\mathcal{M}_{2}}(A)\;\geq 1-\alpha,\qquad y\notin A.$
Because $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$, the same inequality holds with $\mathcal{M}_{2}$ replaced by $\mathcal{M}_{1}$, so $A$ belongs to the family of sets whose intersection defines $B_{\mathcal{M}_{1}}^{\alpha}$. Consequently $y\in A$, contradicting $y\notin A$. Therefore $y\in B_{\mathcal{M}_{2}}^{\alpha}$, and $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}.$ Hence $\mathsf{IHDR}_{\alpha}$ maps the arrow $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ to the arrow $B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha}$.
(iii) Identities. For each $\mathcal{M}\in\mathscr{C}$, the identity morphism is $\mathcal{M}\hookrightarrow\mathcal{M}$. Since inclusion is reflexive, $B_{\mathcal{M}}^{\alpha}\hookrightarrow B_{\mathcal{M}}^{\alpha}$ is the identity in $\Sigma_{Y}$, so $\mathsf{IHDR}_{\alpha}(\mathrm{id}_{\mathcal{M}})=\mathrm{id}_{B_{\mathcal{M}}^{\alpha}}$.
(iv) Composition. Given $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}\subseteq\mathcal{M}_{3}$, the composite morphism in $\mathscr{C}$ is $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{3}$. By monotonicity, $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}\subseteq B_{\mathcal{M}_{3}}^{\alpha},$ so
$\mathsf{IHDR}_{\alpha}(\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{3})=B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{3}}^{\alpha}$ $=(B_{\mathcal{M}_{2}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{3}}^{\alpha})\circ(B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha})$
$=\mathsf{IHDR}_{\alpha}(\mathcal{M}_{2}\hookrightarrow\mathcal{M}_{3})\circ\mathsf{IHDR}_{\alpha}(\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}).$
Since $\mathsf{IHDR}_{\alpha}$ respects objects, morphisms, identities and composition, it is a covariant functor. ∎
Remark 5 (Contravariant Variants).
Replacing $B_{\mathcal{M}}^{\alpha}$ by its complement $Y\setminus B_{\mathcal{M}}^{\alpha}$ yields a contravariant functor. Allowing $\alpha$ itself to vary introduces a functor $\mathsf{IHDR}\colon[0,1]\times\mathscr{C}\rightarrow\Sigma_{Y}$ that is antitone in the first argument and covariant in the second.
To the best of our knowledge, Theorem 14 is the first time that an imprecise probabilistic notion such as the IHDR is framed as a functor between two categories, thus adding to the field of category-theoretic imprecise probability theory (Liell-Cock and Staton,, 2024).
We also note in passing that we will not delve into the examination of categories $(\mathscr{C},\subseteq)$ and $(\Sigma_{Y},\subseteq)$ because they are two special instances of the well-studied category $\mathbf{Sub}(X)$ of subsets of a generic set $X$, endowed with inclusion morphisms. Here we abuse notation: We write $(\mathscr{C},\subseteq)$ to denote the category whose objects are the elements of $\mathscr{C}$, and whose morphisms are the inclusions, and similarly for $(\Sigma_{Y},\subseteq)$.
We have the following corollary, which allows us to see the Conformal Prediction Region outputted by a Full Conformal Prediction procedure as an object image of functor $\mathsf{IHDR}_{\alpha}$.
Corollary 14.1 (Conformal Prediction as a Functor Image).
Suppose that the conformal transducer $\pi(\cdot,y^{n})$ is consonant, i.e.
$\sup_{y_{n+1}\in Y}\pi(y_{n+1},y^{n})=1.$
Then, for all $\alpha\in[0,1]$, all $y^{n}\in Y^{n}$, and all $\psi\in\mathscr{F}$, we have that
$\mathscr{R}_{\alpha}^{\psi}(y^{n})=\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi})).$
That is, the $\alpha$-level Conformal Prediction Region $\mathscr{R}_{\alpha}^{\psi}(y^{n})$ is the image under functor $\mathsf{IHDR}_{\alpha}$ of a well-defined credal set
$\mathcal{M}(\overline{\Pi})\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\coloneqq\sup_{y_{n+1}\in A}\pi(y_{n+1},y^{n})\text{, }\forall A\in\Sigma_{Y}\},$
which is an object of category $(\mathscr{C},\subseteq)$.
Proof.
Immediate from Theorem 14 and Caprio et al., 2025a (, Proposition 5). ∎
Viewing the conformal-prediction map $\mathsf{IHDR}_{\alpha}$ as a covariant functor immediately yields practical AI privacy benefits. Suppose that we have many sources producing credal sets $\mathcal{M}_{j}$. First, any differentially-private or otherwise noise-adding transformation applied to each site’s credal set $\mathcal{M}_{j}$ propagates through the functor without violating the coverage guarantee, so statistical validity is preserved “for free”.
Second, collaboration can proceed in a strictly federated manner: Every agent needs to transmit only its (possibly privatized) summary object $\mathcal{M}_{j}$—never the underlying data—thereby minimizing exposure while still enabling the consortium to construct a globally valid prediction region.
7. Further Properties of our Categories
In this section, we ask ourselves whether we can further characterize the categories that we introduced earlier in Sections 3 and 4. Throughout the section, if restrictions have to be imposed on the objects or morphisms of a category, we will report them as superscript and subscripts of the ensuing subcategory, respectively.
7.1. Monoidal Categories
The first, natural question is whether they are monoidal.
Alas, $\mathbf{UHCont}$ is not, as upper hemicontinuity is not preserved under product in general unless the correspondences are compact-valued. We can define a slight strengthening of $\mathbf{UHCont}$, though, which is monoidal.
[…]
Define a structure $\mathbf{UHCont}_{\text{c}}$ as follows,
  • Objects: Topological spaces $(X,\tau_{X})$.
  • Morphisms: Compact-valued, upper hemicontinuous correspondences $\Phi:X\rightrightarrows Y$.
Theorem 16 ($\mathbf{UHCont}_{\text{c}}$ is a Monoidal Category).
The structure $\mathbf{UHCont}_{\text{c}}$ is a monoidal category under the Cartesian product of topological spaces.
Proof.
We verify the monoidal category structure.
(1) Tensor Product of Objects and Morphisms. For objects $X,Y$, define $X\otimes Y\coloneqq X\times Y$, the Cartesian product with the product topology.
Given morphisms $\Phi_{1}:X\rightrightarrows X^{\prime}$ and $\Phi_{2}:Y\rightrightarrows Y^{\prime}$, define their tensor product as
$(\Phi_{1}\otimes\Phi_{2})(x,y)\coloneqq\Phi_{1}(x)\times\Phi_{2}(y).$
Since compact sets are preserved under finite products, $\Phi_{1}(x)\times\Phi_{2}(y)$ is compact for all $(x,y)\in X\times Y$. Also, the product of upper hemicontinuous correspondences with compact values is upper hemicontinuous with compact values (Aliprantis and Border,, 2006, Theorem 17.28). Thus, $\Phi_{1}\otimes\Phi_{2}$ is a morphism in $\mathbf{UHCont}_{\text{c}}$.
(2) Unit Object. Let the unit object be the singleton space $I\coloneqq\{*\}$, with the trivial topology. For any object $X$, the canonical identifications
$\lambda_{X}:I\times X\rightarrow X,\quad\rho_{X}:X\times I\rightarrow X$
are homeomorphisms. They define morphisms in $\mathbf{UHCont}_{\text{c}}$ via single-valued, continuous correspondences (e.g., $(x,*)\mapsto x$) with compact singleton values.
(3) Associativity. Define the associator as the canonical homeomorphism,
$\alpha_{X,Y,Z}:(X\times Y)\times Z\rightarrow X\times(Y\times Z).$
Its graph defines a single-valued, continuous correspondence with compact-valued singleton fibers. Hence, it is a morphism in the category.
(4) Coherence Conditions. The coherence diagrams (pentagon and triangle) commute strictly, as they correspond to natural identifications of Cartesian products in the category of topological spaces. Thus, the associators and unitors satisfy the coherence conditions. ∎
Given our previous proofs, it is easy to see that, under conditions (i)-(iv), the Full Conformal Prediction correspondence $\kappa$ is a morphism of $\mathbf{UHCont}_{\text{c}}$. This is because $Y$ is compact Hausdorff, and the graph of $\kappa$ is closed by the Proof of Theorem 5. By the Closed Graph Theorem (Lemma 3), this means that $\kappa$ is upper hemicontinuous and closed-valued; But $Y$ is compact Hausdorff, which implies that $\kappa$ is compact-valued.
As perhaps expected from the proof of Theorem 16, the category $\mathbf{WMeas}_{\text{uc}}$ is monoidal.
Theorem 17 ($\mathbf{WMeas}_{\text{uc}}$ is a Monoidal Category).
The structure $\mathbf{WMeas}_{\text{uc}}$ forms a monoidal category under the Cartesian product.
Proof.
Let the unit object be the singleton space $I\coloneqq\{*\}$, with the trivial topology, and
$(X,\Sigma_{X})\,\otimes\,(X^{\prime},\Sigma_{X^{\prime}})\;\coloneqq\;(X\times X^{\prime},\,\Sigma_{X}\!\otimes\!\Sigma_{X^{\prime}}),\qquad(F\otimes G)(x,x^{\prime})\;\coloneqq\;F(x)\times G(x^{\prime}).$
1. Tensor is a bifunctor. For morphisms $F:X\!\Rightarrow\!Y$ and $G:X^{\prime}\!\Rightarrow\!Y^{\prime}$ the map $(F\otimes G)(x,x^{\prime})=F(x)\times G(x^{\prime})$ is:
Uniformly compact. If $K_{F}\subset Y$ and $K_{G}\subset Y^{\prime}$ witness compactness, $K_{F}\times K_{G}$ is compact by Tychonoff, and $(F\otimes G)(x,x^{\prime})\subset K_{F}\times K_{G}$ for all $(x,x^{\prime})$.
Weakly measurable. Let $O=\bigcup_{k\in\mathbb{N}}(B_{k}\times C_{k})$ be an open subset of $Y\times Y^{\prime}$ written as a countable union of basic rectangles (recall that $Y,Y^{\prime}$ are second-countable). Then,
$(F\otimes G)^{-1}(O)\;=\;\bigcup_{k\in\mathbb{N}}\bigl{(}\{x:F(x)\cap B_{k}\!\neq\emptyset\}\times\{x^{\prime}:G(x^{\prime})\cap C_{k}\!\neq\emptyset\}\bigr{)},$
a countable union of measurable rectangles, hence measurable in $\Sigma_{X}\!\otimes\!\Sigma_{X^{\prime}}$.
Functoriality is immediate: $(G_{1}\!\circ\!F_{1})\otimes(G_{2}\!\circ\!F_{2})=(G_{1}\otimes G_{2})\!\circ\!(F_{1}\otimes F_{2})$ because product distributes over union, and $1_{X}\otimes 1_{X^{\prime}}=1_{X\times X^{\prime}}$.
2. Associativity, unit, symmetry. Because $\otimes$ is the cartesian product on objects and the point-wise cartesian product on sets, we have equalities (not merely isomorphisms),
$(X\otimes Y)\otimes Z=X\otimes(Y\otimes Z),\qquad X\otimes\{\ast\}=X=\{\ast\}\otimes X.$
The associator and unitors are therefore identities, so the monoidal structure is strict. The symmetry is the flip map $\sigma_{X,Y}:(x,y)\mapsto(y,x)$; naturality and the hexagon condition hold because the cartesian product is symmetric in the category Set, and every construction is inherited verbatim.
3. Coherence. Since all coherence morphisms are identities, MacLane’s pentagon and triangle commute trivially.
Hence $(\mathbf{WMeas}_{\mathrm{uc}},\otimes,\{\ast\})$ is a strict symmetric monoidal category. ∎
7.2. Functors Between Categories
Another question that we ask ourselves is whether there is a canonical functor between $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$ (or vice versa). The answer, as the attentive reader may expect, is no. The reason is that there exists a non-functorial gap between measurability and continuity, two different concepts that also motivated us to consider categories $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$ to begin with.
Functors do exist if we focus on restrictions of $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$. In particular, call $\mathbf{WMeas}_{\text{uc,uhc}}\subset\mathbf{WMeas}_{\text{uc}}$ the category whose objects are measurable Polish spaces with the Borel $\sigma$-algebra, and whose morphisms are uniformly compact-valued, weakly measurable, and upper hemicontinuous correspondences. Then, the following holds.
Theorem 18 (Faithful Functor between $\mathbf{WMeas}_{\text{uc,uhc}}$ and $\mathbf{UHCont}$).
There exists a faithful functor
$\mathcal{F}:\mathbf{WMeas}_{\text{uc,uhc}}\rightarrow\mathbf{UHCont}$
defined by
  • $\mathcal{F}(X,\tau_{X},\mathcal{B}(\tau_{X}))\coloneqq(X,\tau_{X})$,
  • $\mathcal{F}(\Phi)\coloneqq\Phi$ (viewed as a topological correspondence).
Proof.
We verify that $\mathcal{F}$ is a functor.
(1) Identity. The identity correspondence $\mathrm{id}_{X}(x)=\{x\}$ is trivially compact-valued, upper hemicontinuous, and weakly measurable. Thus, $\mathcal{F}(\mathrm{id}_{X})=\mathrm{id}_{X}$ is a valid morphism in $\mathbf{UHCont}$.
(2) Composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be morphisms in $\mathbf{WMeas}_{\text{uc,uhc}}$. Then,
  • Both are uniformly compact-valued and upper hemicontinuous;
  • Their composition $\Psi\circ\Phi$ is also uniformly compact-valued (Aliprantis and Border,, 2006, Page 558 and (a consequence of) Example 17.24) and upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23);
  • Weak measurability is preserved under composition (by the Proof of Theorem 7).
So $\Psi\circ\Phi$ is a morphism in $\mathbf{WMeas}_{\text{uc,uhc}}$, and $\mathcal{F}(\Psi\circ\Phi)=\mathcal{F}(\Psi)\circ\mathcal{F}(\Phi)$.
(3) Faithfulness. If $\Phi\neq\Psi$ as correspondences in $\mathbf{WMeas}_{\text{uc,uhc}}$, then $\mathcal{F}(\Phi)\neq\mathcal{F}(\Psi)$ in $\mathbf{UHCont}$. So $\mathcal{F}$ is injective on morphisms. Therefore, $\mathcal{F}$ is a well-defined faithful functor. ∎
Functor $\mathcal{F}$ in Theorem 18 “forgets” the measurable structure but preserves the topological one. The condition of weak measurability ensures that the image $\Phi$ remains a valid morphism in both categories.
It is also worth mentioning that, in general, $\mathcal{F}$ does not have a limit in $\mathbf{UHCont}$ (Perrone,, 2024, Definition 3.1.5). This is because the representability conditions fail: Universal lifting of upper hemicontinuous cones is obstructed by the loss of measurability information via $\mathcal{F}$. This is a consequence of the incompatibility between continuity-based and measurability-based structures, exactly the same structural gap that made defining functor $\mathcal{F}$ nontrivial in the first place.
Let now $\mathbf{UHCont}_{\text{c}}^{\text{Pa}}\subset\mathbf{UHCont}$ be the category whose objects are Polish topological spaces, and whose morphisms are upper hemicontinuous, uniformly compact-valued correspondences having analytic graphs. Then, the following holds.
Theorem 19 (Faithful Functor between $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}$ and $\mathbf{WMeas}_{\text{uc}}$).
There exists a faithful functor
$\mathcal{G}:\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}\rightarrow\mathbf{WMeas}_{\text{uc}}$
defined by
  • $\mathcal{G}(X,\tau_{X})\coloneqq(X,\tau_{X},\mathcal{B}(\tau_{X}))$,
  • $\mathcal{G}(\Phi)\coloneqq\Phi$, viewing the correspondence as a measurable one.
Proof.
First, notice that if $X$ is a Polish space with topology $\tau_{X}$, then $(X,\tau_{X},\mathcal{B}(\tau_{X}))$ is a measurable Polish space, a valid object in $\mathbf{WMeas}_{\text{uc}}$.
Also, let $\Phi:X\rightrightarrows Y$ be a morphism in $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}$, i.e. uniformly compact-valued, upper hemicontinuous, and with analytic graph. Then, since $\Phi$ is compact-valued, it follows that $\Phi$ is weakly measurable (Aliprantis and Border,, 2006, Theorem 18.10). Thus $\Phi$ is a valid morphism in $\mathbf{WMeas}_{\text{uc}}$.
We now verify the functorial properties.
(1) Identity. The identity correspondence $x\mapsto\{x\}$ has analytic graph, is trivially upper hemicontinuous and (uniformly) compact-valued, hence maps to the identity in $\mathbf{WMeas}_{\text{uc}}$.
(2) Composition. The composition of two (uniformly) compact-valued, upper hemicontinuous, analytic-graph correspondences again satisfies all three properties (using standard results from multivalued analysis), so $\mathcal{G}(\Psi\circ\Phi)=\mathcal{G}(\Psi)\circ\mathcal{G}(\Phi)$. [^11] The graph of the composite $G_{\Psi\circ\Phi}=\bigl{\{}(x,z):\exists y,\,(x,y)\in G_{\Phi},\;(y,z)\in G_{\Psi}\bigr{\}}$ is the analytic image (Souslin projection) of the analytic set $G_{\Phi}\times Z\cap X\times G_{\Psi}$; Hence it is analytic.
(3) Faithfulness. Since $\mathcal{G}$ is the identity on morphisms, it is injective on $\operatorname{Hom}$-sets. ∎
Notice that the analytic graph condition in Theorem 19 is essential. Upper hemicontinuity alone does not imply weak measurability; Hence, $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{Pa}}$ is the minimal natural subcategory for which functor $\mathcal{G}$ is defined.
We also mention that, just as continuity cannot capture all measurable behavior, measurability cannot reconstruct topological properties like upper hemicontinuity. Therefore, $\mathcal{G}$ typically fails to admit a limit object in $\mathbf{WMeas}_{\text{uc}}$ (Perrone,, 2024, Definition 3.1.5) representing $\text{Cone}(-,\mathcal{G})$, for precisely the same reason $\mathcal{F}$ does: The two categories encode fundamentally distinct structures.
7.3. Monads
We now ask ourselves whether we can find monads on the categories $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$. The general answer is no, but we can find interesting subcategories for which a monad can be found.
Definition 20 (Structure $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$).
Define the structure $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ as follows,
  • Objects: Compact Hausdorff spaces $X$.
  • Morphisms: Compact-valued, upper hemicontinuous $\Phi:X\rightrightarrows Y$.
It is immediate to see that structure $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ is a subcategory of $\mathbf{UHCont}$. It admits a monad.
Theorem 21 (Vietoris Monad on $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$).
There exists a monad $(T,\eta_{X},\nu_{X})$ on the category $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$, where
  • $T(X)\coloneqq\mathcal{K}(X)$, the space of compact subsets of $X$, endowed with the Vietoris topology;
  • For a morphism $\Phi:X\rightrightarrows Y$, define $T(\Phi):\mathcal{K}(X)\rightrightarrows\mathcal{K}(Y)$ by
    $T(\Phi)(K)\;\coloneqq\;\bigl{\{}\,L\in\mathcal{K}(Y)\;:\;L\subset\Phi[K]\bigr{\}},\quad\Phi[K]\coloneqq\bigcup_{x\in K}\Phi(x);$
  • The unit $\eta_{X}:X\rightarrow\mathcal{K}(X)$ is given by $\eta_{X}(x)=\{x\}$;
  • The multiplication $\nu_{X}:\mathcal{K}(\mathcal{K}(X))\rightarrow\mathcal{K}(X)$ is defined by
    $\nu_{X}(\mathcal{A})\coloneqq\bigcup_{A\in\mathcal{A}}A,$
    where $\mathcal{K}(\mathcal{K}(X))$ is the hyperspace of nonempty, compact collections of compact subsets of $X$.
Proof.
We verify the monad axioms.
(1) Well-definedness.
  • $T(X)$ is a compact Hausdorff space (classical result).
  • We now verify that $T(\Phi)$ is a morphism in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. By Lemma 3, $\Phi$ has a closed graph; It also has compact values. It follows from standard results on the Vietoris topology (Michael,, 1951) that
    • The map $K\mapsto T(\Phi)(K)$ is upper hemicontinuous;
    • The values $T(\Phi)(K)$ are nonempty, compact subsets of $\mathcal{K}(Y)$, hence compact in the Vietoris topology.
    Thus, $T(\Phi)$ is again a compact-valued, upper hemicontinuous correspondence. Finally, for morphisms $\Phi:X\!\rightrightarrows\!Y$ and $\Psi:Y\!\rightrightarrows\!Z$ we can see that
    $L\subset(\Psi\circ\Phi)[K]\iff\exists M\subset\Phi[K]\;\bigl{(}L\subset\Psi[M]\bigr{)},$
    hence $T(\Psi\circ\Phi)=T(\Psi)\circ T(\Phi)$; Thus $T$ is indeed a functor.
(2) Naturality of $\eta_{X}$. For each $x\in X$, $\eta_{X}(x)=\{x\}$ is a continuous map (singletons are closed in compact Hausdorff spaces and form a continuous embedding into $\mathcal{K}(X)$), and hence upper hemicontinuous and compact-valued.
(3) Naturality of $\nu_{X}$. Given $\mathcal{A}\in\mathcal{K}(\mathcal{K}(X))$, $\nu_{X}(\mathcal{A})=\bigcup\mathcal{A}$ is continuous in the Vietoris topology, [^12] For basic opens $\langle U_{1},\dots,U_{m}\rangle\subset\mathcal{K}(X)$ we have $\nu_{X}^{-1}\!\langle U_{1},\dots,U_{m}\rangle=\bigl{\langle}\,\langle U_{1}\rangle,\dots,\langle U_{m}\rangle,\,\langle U_{1}\cup\cdots\cup U_{m}\rangle\bigr{\rangle},$ which is open in $\mathcal{K}(\mathcal{K}(X))$, see Definition 29. and hence upper hemicontinuous and compact-valued.
(4) Monad laws.
  • Left unit: $\nu_{X}\circ T(\eta_{X})(K)=\bigcup_{x\in K}\eta_{X}(x)=\bigcup_{x\in K}\{x\}=K$.
  • Right unit: $\nu_{X}\circ\eta_{T(X)}(K)=\nu_{X}(\{K\})=K$.
  • Associativity: $\nu_{X}\circ T(\nu_{X})=\nu_{X}\circ\nu_{T(X)}$ holds due to the associativity of set union.
Therefore, $(T,\eta_{X},\nu_{X})$ is a monad on $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. ∎
The monad in Theorem 21 is the classical Vietoris monad (Garner,, 2020), adapted to the category of compact Hausdorff spaces with compact-valued, upper hemicontinuous correspondences.
We now show that, for any $n\in\mathbb{N}$, the (minimal) sufficient conditions for the Full Conformal Prediction diagram to commute in category $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ are the same that make it commute in category $\mathbf{UHCont}$.
Theorem 22 (The Conformal Prediction Diagram Commutes, $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$).
Assume that (i)-(iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$, $\mathsf{CRED}_{\alpha}$, and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}_{\alpha}=\kappa_{\alpha}$. That is, the Full CP Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$.
Proof.
We show that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are compact-valued. The claim, then, follows from an argument similar to that in the Proof of Theorem 5. Throughout, $Y$ is compact Hausdorff by (i); Every closed subset of $Y$ is therefore compact.
1. $\kappa$. Fix $(\alpha,y^{n},\psi)$ with $\alpha\notin S_{n+1}$, so to satisfy (iii). Define $\beta(\alpha)\coloneqq\min\{g\in S_{n+1}:g>\alpha\}$ as in (8). The conformal transducer $\pi(y,y^{n},\psi)$ is an upper semicontinuous map of $y$: Every summand in its definition is the indicator of a closed set and finite averages preserve u.s.c. Hence
$\kappa(\alpha,y^{n},\psi)=\bigl{\{}y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\bigr{\}},$
a super-level set of a u.s.c. function, therefore closed in $Y$, and thus compact.
2. $\mathsf{CRED}$. For a countable generator of closed sets $\{A_{j}\}\subset\Sigma_{Y}$,
$\mathsf{CRED}(y^{n},\psi)=\bigcap_{j}\bigl{\{}P\in\Delta_{Y}:P(A_{j})\leq\overline{\Pi}_{y^{n},\psi}(A_{j})\bigr{\}}.$
Since $P\mapsto P(A_{j})$ is weak-continuous and $j\mapsto\overline{\Pi}_{y^{n},\psi}(A_{j})$ is u.s.c. by Lemma 4, each half-space is closed in the compact space $\Delta_{Y}$; Intersection keeps closedness, and in turn compactness.
3. $\mathsf{IHDR}$. For fixed $(\alpha,\mathcal{M})$, put
$C_{\alpha}(\mathcal{M})\coloneqq\bigl{\{}A\in\Sigma_{Y}:\underline{P}_{\!\mathcal{M}}(A)\geq 1-\alpha\bigr{\}}.$
Each $A\in C_{\alpha}(\mathcal{M})$ is closed (complements of open sets have lower probability $\leq\alpha$). Hence $\mathsf{IHDR}(\alpha,\mathcal{M})=\bigcap_{A\in C_{\alpha}(\mathcal{M})}A$ is an intersection of closed subsets of the compact $Y$, hence compact. ∎
Theorem 22 gives us minimal sufficient conditions so that (i) Full Conformal Prediction is a morphism of a category, $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$, that has a monad, and (ii) the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. We now look at a subcategory of $\mathbf{WMeas}_{\text{uc}}$ that has a monad.
Definition 23 (Structure $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$).
Define a structure $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ as follows,
  • Objects: Compact Polish spaces $X$ equipped with their Borel $\sigma$-algebra $\mathcal{B}(X)$.
  • Morphisms: Uniformly compact-valued, weakly measurable correspondences $\Phi:X\rightrightarrows Y$.
It is immediate to see that structure $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ is a subcategory of $\mathbf{WMeas}_{\text{uc}}$, as its objects need to be compact as well. We now show that it admits a monad.
Let $X$ be an object of $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$. Define an endofunctor $\mathcal{K}\colon\mathbf{WMeas}_{\text{uc}}^{\text{c}}\to\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, a unit transformation $\eta_{X}\equiv\eta:\mathrm{Id}_{\mathbf{WMeas}_{\text{uc}}^{\text{c}}}\!\Rightarrow\!\mathcal{K}$ and a multiplication $\nu_{X}\equiv\nu:\mathcal{K}\!\circ\!\mathcal{K}\Rightarrow\mathcal{K}$ by
$\mathcal{K}(X)$$\coloneqq\bigl{\{}C\subseteq X:C\neq\emptyset,\;C\text{ compact}\bigr{\}},$
$\mathcal{K}(\Phi)(C)$$\coloneqq\bigl{\{}\,\overline{\Phi[C]}\bigr{\}},\quad C\in\mathcal{K}(X),$
$\eta_{X}(x)$$\coloneqq\bigl{\{}\,\{x\}\bigr{\}},\quad x\in X,$
$\nu_{X}(\mathcal{C})$$\coloneqq\bigl{\{}\bigcup_{C\in\mathcal{C}}C\bigr{\}},\quad\mathcal{C}\in\mathcal{K}\!\bigl{(}\mathcal{K}(X)\bigr{)}.$
Theorem 24 (Vietoris monad on $\mathbf{WMeas}_{\mathrm{uc}}^{\mathrm{c}}$).
The triple $(\mathcal{K},\eta,\nu)$ satisfies the monad axioms inside $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$. Moreover, $\mathcal{K}(X)\not\cong X$ for any $X$ other than a singleton, so the monad is non-trivial.
Proof.
1. $\mathcal{K}$ is an endofunctor.
  1. (1)
    (Objects). For any compact Polish space $X$, $\mathcal{K}(X)$ endowed with the Vietoris topology is again compact Polish; Its Borel $\sigma$-algebra makes $\mathcal{K}(X)$ an object of the category $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$.
  2. (2)
    (Morphisms). Given $\Phi:X\rightrightarrows Y$ in $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, each fiber of $\mathcal{K}(\Phi)$ is a singleton $\{\overline{\Phi[C]}\}$, hence compact; The common compact bound is $K_{\Phi}$, the one already bounding $\Phi$. Weak measurability follows because inverse images of Vietoris subbasic open sets can be expressed via the measurable “hit” and “miss” sets of $\Phi$ (studied below). Functoriality is immediate,
    $\mathcal{K}(\operatorname{id}_{X})=\operatorname{id}_{\mathcal{K}(X)},\qquad\mathcal{K}(\Psi\circ\Phi)=\mathcal{K}(\Psi)\circ\mathcal{K}(\Phi),$
    since closures commute with unions in compact spaces.
2. The unit $\eta$. For every $X$, $\eta_{X}$ has compact fibres $\{x\}$, is uniformly bounded by $\mathcal{K}(X)$, and is weakly measurable because $\eta_{X}^{u}(\langle U\rangle)\coloneqq\{x\in X:\eta_{X}(x)\cap\langle U\rangle\neq\emptyset\}=U\in\mathcal{B}(X)$ for each open $U\subseteq X$. Naturality of $\eta$ is trivial: $\mathcal{K}(\Phi)\circ\eta_{X}=\eta_{Y}\circ\Phi$ for every $\Phi:X\rightrightarrows Y$.
3. The multiplication $\nu$. For $\mathcal{C}\in\mathcal{K}\bigl{(}\mathcal{K}(X)\bigr{)}$, the union $\bigcup_{C\in\mathcal{C}}C$ is compact (union of a compact family of compact subsets in a compact space). Thus $\nu_{X}(\mathcal{C})\in\mathcal{K}(X)$, giving well-defined singleton fibres. Uniform boundedness is by $\mathcal{K}(X)$ itself. Weak measurability is checked on subbasic Vietoris open sets:
  • Hit part: $\nu_{X}^{-1}\!\langle U\rangle=\bigl{\{}\mathcal{C}:\exists C\in\mathcal{C},\;C\cap U\neq\emptyset\bigr{\}},$ which is open in the Vietoris topology of $\mathcal{K}(\mathcal{K}(X))$.
  • Miss part: $\nu_{X}^{-1}\langle U_{1},\dots,U_{m}\rangle=\bigl{\{}\mathcal{C}:\textstyle\bigcup_{C\in\mathcal{C}}C\subseteq U_{1}\cup\dots\cup U_{m},\;C\cap U_{i}\neq\emptyset\text{ }foralli\bigr{\}},$ again Vietoris-open.
Naturality of $\nu$ follows from $\nu_{Y}\circ\mathcal{K}(\mathcal{K}(\Phi))=\mathcal{K}(\Phi)\circ\nu_{X}$, since direct images commute with unions and closure.
4. Monad laws. For every $X$ the diagrams
$\mathcal{K}^{3}(X)\;\xrightarrow{\mathcal{K}\nu_{X}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)\quad=\quad\mathcal{K}^{3}(X)\;\xrightarrow{\nu_{\mathcal{K}(X)}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)$
and
$\mathcal{K}(X)\;\xrightarrow{\mathcal{K}\eta_{X}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)\quad$ $=\quad\mathcal{K}(X)\;\xrightarrow{\mathrm{id}}\;\mathcal{K}(X)$
$=\quad\mathcal{K}(X)\;\xrightarrow{\eta_{\mathcal{K}(X)}}\;\mathcal{K}^{2}(X)\;\xrightarrow{\nu_{X}}\;\mathcal{K}(X)$
commute because unions of singletons reproduce the original compact set, and unions are associative.
5. Non-triviality. If $X$ has at least two points, then $\mathcal{K}(X)$ is strictly larger than $X$ under any continuous, measure-compatible identification; Hence $(\mathcal{K},\eta,\nu)$ is not isomorphic to the identity monad. ∎
The monad that we found on the category $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$ is again the classic Vietoris monad, adapted to the category of compact Polish spaces (with their Borel $\sigma$-algebra), with uniformly compact-valued, weakly measurable correspondences. It is immediate to see that Theorem 8 still holds for $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$, since we simply require the extra assumption that the objects are compact measurable Polish spaces.
We have a nice byproduct of focusing on the (sub)categories that we considered in this section.
Proposition 25 (Forgetful Embedding between $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ and $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$).
There is a faithful functor
$F\;:\;\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}\;\rightarrow\;\mathbf{WMeas}_{\text{uc}}^{\text{c}},$
defined as follows.
  1. (Obj)
    For every compact Hausdorff space $X$, pick any compatible complete separable metric $d_{X}$; The resulting topological space is compact Polish. Set $F(X)\coloneqq(X,d_{X})$ equipped with its Borel $\sigma$–algebra.
  2. (Mor)
    For a morphism $\Phi:X\rightrightarrows Y$ in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ (i.e. compact–valued and upper hemicontinuous), let
    $F(\Phi)\;\coloneqq\;\Phi\qquad(\text{same graph and fibres}).$
With these assignments, $F$ preserves identities and composition, and on every hom-set it is injective. Hence, $F$ is faithful.
Proof.
1. $F(\Phi)$ is a morphism in $\mathbf{WMeas}_{\text{uc}}^{\text{c}}$. Because $Y$ itself is compact, $\Phi(x)\subseteq Y$ for all $x$, so $K_{\Phi}\coloneqq Y$ is a common compact bound; Thus $F(\Phi)$ is uniformly compact-valued. To show weak measurability, fix an open $O\subseteq Y$ and put $H\coloneqq\{x\in X:\Phi(x)\cap O\neq\emptyset\}.$ If $x_{0}\notin H$ then $\Phi(x_{0})\subseteq Y\setminus O$ (a closed set). Upper hemicontinuity yields a neighborhood $U$ of $x_{0}$ with $\Phi(U)\subseteq Y\setminus O$, so $U\subseteq X\setminus H$. Hence $X\setminus H$ is open and $H$ is closed, thus Borel; This is precisely the weak measurability condition for $F(\Phi)$.
2. Functoriality. $F$ leaves the underlying correspondence unchanged, so $F(\mathrm{id}_{X})=\mathrm{id}_{F(X)}$ and $F(\Psi\circ\Phi)=F(\Psi)\circ F(\Phi)$, for all composable $\Phi,\Psi$.
3. Faithfulness. On each hom-set, $F$ acts as the identity map $\Phi\mapsto\Phi,$ hence it is injective.
4. Domain of $F$. If every object of $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ is assumed to be metrizable (equivalently, second-countable), then the object assignment is defined on the whole category. Otherwise, $F$ can still be applied to the metrizable subcategory, for which the conclusion holds verbatim. ∎
Alas, there is no faithful functor in the reverse direction. Weakly measurable correspondences admit far more pathological behavior than upper hemicontinuous ones. Any attempt to encode all that extra information inside upper hemicontinuous morphology must necessarily collapse distinct measurable maps, breaking faithfulness.
8. Conclusion
In this paper, we took a Category Theory route to show that Full Conformal Prediction exhibits intrinsic UQ capabilities. This is a result of framing it as a morphism of two categories, UHCont and $\textbf{WMeas}_{\text{uc}}$, and of showing that it is part of a diagram that commutes. As byproduct, we were able to prove that CP methods bridge Bayesian, frequentist, and imprecise approaches to statistical prediction. We also studied the properties of such categories, and we pointed out how a Conformal Prediction Region can be seen as a functor image.
This work opens up the study of Categorical Conformal Prediction, and hints at many unsolved problems. For example, it is an open question how to leverage the structure of the functional-analysis-flavored categories UHCont and $\textbf{WMeas}_{\text{uc}}$ to learn further aspects of the CP methodology.
In addition, it is left to inspect whether studying CP as a morphism of different categories can allow us to remove the need for assumptions altogether, in particular compactness of the state-space $Y$.
Finally, it is left for the future to show whether our results extend to Split Conformal Prediction, and to any other extension of classical Conformal Prediction that ML researchers introduced in recent years.
Acknowledgments
We wish to express our gratitude to Sam Staton and Paolo Perrone for insightful discussions on the possibility of formulating Conformal Prediction in the Category Theory framework. We are also extremely grateful to Sam Staton for his invitation to the University of Oxford, where the idea in the present paper germinated. Finally, we thank Yusuf Sale and Eyke Hüllermeier for stimulating discussions on the Uncertainty Quantification capabilities of Conformal Prediction, Sangwoo Park for his idea that led to Proposition 12, and Alessandro Zito for his patience in helping us shape the abstract and introduction.
Appendix A Background in Functional Analysis
The concepts we present in this section are studied in detail in Aliprantis and Border, (2006, Chapters 17 and 18).
A correspondence $\Phi:X\rightrightarrows Y$ from a set $X$ to a set $Y$ assigns to each $x\in X$ a subset $\Phi(x)\in 2^{Y}$. The image of a set $A\subseteq X$ under $\Phi$ is the set $\Phi(A)\coloneqq\cup_{x\in A}\Phi(x)$. Given two correspondences $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$, their composition is defined naturally as
$\Psi\circ\Phi:X\rightrightarrows Z,\quad(\Psi\circ\Phi)(x)\coloneqq\bigcup_{y\in\Phi(x)}\Psi(y).$
The graph of $\Phi$ is given by $\text{Gr}\Phi\coloneqq\{(x,y)\in X\times Y:y\in\Phi(x)\}$.
We now turn our attention to continuity and measurability notions for correspondences.
Definition 26 (Upper Hemicontinuity).
Let $X$ and $Y$ be topological spaces. A correspondence $\Phi:X\rightrightarrows Y$ is upper hemicontinuous at $x_{0}\in X$ if for every open set $V\subseteq\Phi(x_{0})$, there exists an open neighborhood $U$ of $x_{0}$ such that for all $x\in U$, $\Phi(x)\subseteq V$.
We say that $\Phi$ is upper hemicontinuous (uhc) if it is upper hemicontinuous at every point $x\in X$.
As the name suggests, an alternative notion of continuity for correspondences, called lower hemicontinuity, can be defined (Aliprantis and Border,, 2006, Definition 17.2). In this work, we focus primarily on uhc for its historical importance, especially in fields such as economics and game theory (Börgers,, 1991).
The following is the definition of how a correspondence can be considered measurable in the weakest possible sense.
Definition 27 (Weakly Measurable Correspondence).
Let $(X,\Sigma_{X})$ be a measurable space and $(Y,\tau_{Y})$ a topological space. A correspondence $\Phi:X\rightrightarrows Y$ is said to be weakly measurable if for every open set $O\subseteq Y$, the set
$\Phi^{u}(O)\coloneqq\{x\in X:\Phi(x)\cap O\neq\emptyset\}\in\Sigma_{X}.$
Notice that weak measurability has nothing to do with weak topologies, and that other (stronger) measurability notions can be found in Aliprantis and Border, (2006, Definition 18.1).
Definition 28 (Uniformly Compact-Valued Correspondences).
Let $(X,\Sigma_{X})$ and $(Y,\Sigma_{Y})$ be measurable Polish spaces. A correspondence $\Phi\colon X\rightrightarrows Y$ is called uniformly compact-valued if
  1. (i)
    (compact fibers) $\Phi(x)\subseteq Y$ is non-empty compact, for every $x\in X$;
  2. (ii)
    (uniform compact bound) there exists a one and the same compact set $K_{\Phi}\subseteq Y$ such that $\Phi(x)\subseteq K_{\Phi}$, for all $x\in X$.
Finally, we also recall the definition of the Vietoris topology.
Definition 29 (Vietoris Topology).
Given a compact Hausdorff space $X$, the Vietoris topology on the hyperspace $\mathcal{K}(X)$ of nonempty compact subsets of $X$ is generated by sets of the form
$\langle U_{1},\dots,U_{n}\rangle\coloneqq\left\{{K\in\mathcal{K}(X):K\subseteq\bigcup_{i=1}^{n}U_{i}\text{ and }K\cap U_{i}\neq\emptyset,\text{ for all }i}\right\},$
where each $U_{i}$ is open in $X$.
Appendix B Background in Category Theory
In this section, we brush up on some Category Theory basic notions that we use throughout the paper when deriving our main results. The definitions that we provide all come from Perrone, (2024).
Definition 30 (Category).
A category $\mathbf{C}$ consists of
  • a collection $\mathbf{C}_{0}$ of objects (written $X,Y,Z,\dots$);
  • a collection $\mathbf{C}_{1}$ of morphisms (written $f,g,h,\dots$);
together with
  1. (1)
    for every morphism $f$ there are two distinguished objects called its source and target, written $s(f)$ and $t(f)$ (we abbreviate $f\colon X\rightarrow Y$ when $s(f)=X$, $t(f)=Y$);
  2. (2)
    for every object $X$ an identity morphism $\mathrm{id}_{X}\colon X\rightarrow X$;
  3. (3)
[…]
satisfying the unitality laws $f\circ\mathrm{id}_{X}=f=\mathrm{id}_{Y}\circ f$ and the associativity law $(h\circ g)\circ f=h\circ(g\circ f)$, whenever the expressions are well defined.
Categories are the building blocks in CT. It is natural, then, to ask how to relate them.
Definition 31 (Functor).
Let $\mathbf{C},\mathbf{D}$ be categories. A functor $F\colon\mathbf{C}\rightarrow\mathbf{D}$ assigns
  • to every object $X$ of $\mathbf{C}$ an object $FX$ of $\mathbf{D}$;
  • to every morphism $f\colon X\rightarrow Y$ in $\mathbf{C}$ a morphism $Ff\colon FX\rightarrow FY$ in $\mathbf{D}$,
such that
$F(\mathrm{id}_{X})=\mathrm{id}_{FX},\qquad F(g\circ f)=Fg\circ Ff$
for all composable $f,g$ in $\mathbf{C}$.
A functor $F\colon\mathbf{C}\rightarrow\mathbf{D}$ is called faithful if, for every objects $X,X^{\prime}$ of $\mathbf{C}$ and every morphism $f,f^{\prime}:X\rightarrow X^{\prime}$ of $\mathbf{C}$, we have that $Ff=Ff^{\prime}$ in $\mathbf{D}$ implies $f=f^{\prime}$ in $\mathbf{C}$.
Following up on the question that we asked ourselves before, we can query whether there is a convenient choice of relating two functors.
Definition 32 (Natural transformation).
Let $\mathbf{C},\mathbf{D}$ be categories and let
$F,G:\mathbf{C}\rightarrow\mathbf{D}$
be functors. A natural transformation $\alpha:F\Rightarrow G$ assigns to every object $X$ of $\mathbf{C}$ a morphism
$\alpha_{X}:FX\rightarrow GX$
such that for every morphism $f:X\rightarrow Y$ in $\mathbf{C}$
$Gf\circ\alpha_{X}\;=\;\alpha_{Y}\circ Ff.$
Natural transformation $\alpha$ is a natural isomorphism if for each object $X$ of $\mathbb{C}$, the component $\alpha_{X}:FX\rightarrow GX$ is an isomorphism. In that case, functors $F$ and $G$ are said to be (naturally) isomorphic.
Then, we turn our attention to a special structure on a category, called a monad, that can be thought of as a consistent way of extending spaces to include generalized elements and generalized functions of a specific type. It is of special interest in computer science.
Definition 33 (Monad on a category).
For a category $\mathbf{C}$, a monad is a triple
$(T,\eta,\mu)$
consisting of
  • a functor $T:\mathbf{C}\rightarrow\mathbf{C}$;
  • a natural transformation (the unit)
    $\eta:\operatorname{id}_{\mathbf{C}}\;\Rightarrow\;T;$
  • a natural transformation (the multiplication)
    $\mu:T\!\circ T\;\Rightarrow\;T;$
subject to the associativity and unit laws
$\mu\circ T\mu\;=\;\mu\circ\mu T,\qquad\mu\circ T\eta\;=\;\mu\circ\eta T\;=\;\operatorname{id}_{T}.$
Finally, we are interested in finding a category whose objects and morphisms can be multiplied together in a way that is associative and unital, similarly to what we have for monoids (but only up to isomorphisms). Such categories, called monoidal categories, have wide applications, ranging from information theory to probability.
Definition 34 (Monoidal category).
A monoidal category is a tuple
$(\mathbf{C},\otimes,I,\alpha,\lambda,\rho)$
where
  • $\mathbf{C}$ is a category;
  • $\otimes:\mathbf{C}\times\mathbf{C}\rightarrow\mathbf{C}$ is a product, the tensor product (Perrone,, 2024, Definition 1.3.49); [^13] We do not introduce the tensor product formally in this background section because its definition follows the natural intuition of how a (cartesian) product behaves.
  • $I$ is an object of $\mathbf{C}$ (the unit);
  • $\alpha_{A,B,C}\!:\!A\otimes(B\otimes C)\!\rightarrow\!(A\otimes B)\otimes C$ (associator), $\lambda_{A}\!:I\otimes A\!\rightarrow\!A$, $\rho_{A}\!:A\otimes I\!\rightarrow\!A$ are natural isomorphisms,
satisfying Mac Lane’s pentagon and triangle equalities, for all objects $A,B,C,D$ of $\mathbf{C}$
$\alpha_{A,B,C\otimes D}\!\circ\!\alpha_{A\otimes B,C,D}\;=\;(\alpha_{A,B,C}\otimes\operatorname{id}_{D})\!\circ\!\alpha_{A,B\otimes C,D}\!\circ\!(\operatorname{id}_{A}\otimes\alpha_{B,C,D}),$
$(\operatorname{id}_{A}\otimes\lambda_{B})\!\circ\!\alpha_{A,I,B}\;=\;\rho_{A}\otimes\operatorname{id}_{B}.$