### v3

Abstract.
Conformal prediction (CP) is an Uncertainty Representation technique that delivers finite-sample calibrated prediction regions for any underlying Machine Learning model. Its status as an Uncertainty Quantification (UQ) tool, though, has remained conceptually opaque: While Conformal Prediction Regions (CPRs) give an ordinal representation of uncertainty (larger regions typically indicate higher uncertainty), they lack the capability to cardinally quantify it (twice as large regions do not imply twice the uncertainty). We adopt a category-theoretic approach to CP—framing it as a morphism, embedded in a commuting diagram, of two newly-defined categories—that brings us three joys. First, we show that—under minimal assumptions—CP is intrinsically a UQ mechanism, that is, its cardinal UQ capabilities are a structural feature of the method. Second, we demonstrate that CP bridges the Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning. Finally, we show that a CPR is the image of a covariant functor. This observation is relevant to AI privacy: It implies that privacy noise added locally does not break the global coverage guarantee.
Key words and phrases:
Category Theory; Conformal Prediction; Set-Valued Functions
2020 Mathematics Subject Classification:
Primary: 18D99; Secondary: 62G07; 28B20
1. Introduction
Category Theory (CT) can be thought of as a general theory of mathematical structures and their relations. Recently, there has been a surging interest in studying Machine Learning (ML) from a CT perspective, to leverage the unifying power of the latter (thus unifying apparently unrelated ML subfields), and to obtain more profound theoretical results.
For example, Stehlík, (2012) and Song, (2024) survey how to frame statistical learning models into a CT framework. Shiebler et al., (2021) do the same for gradient-based and probabilistic methods, and invariant and equivariant learning. Furthermore, Cornish, (2025) expands on the latter by studying the problem of symmetrizing a neural network along a group homomorphism. That is, given two groups $H$ and $G$, they study a procedure that converts $H$-equivariant neural networks to $G$-equivariant ones. More in general, CT approaches to probability and statistics (which play a pivotal role in modern-day ML) are active fields of study, see e.g. Fritz et al., (2025); Bohinen and Perrone, (2025); Ensarguet and Perrone, (2024); Ackerman et al., (2024); Matache et al., (2022), and are widely discussed in the nLab and in the n-Category Café.
Conformal Prediction (CP, introduced formally in Section 2.1) is a popular ML prediction tool. Suppose we are given an $n$-sized exchangeable training data $y^{n}=(y_{1},\ldots,y_{n})$ where each observation $y_{i}$ belongs to the state space $Y$, a confidence level $\alpha\in[0,1]$, and a non-conformity score $\psi$ (i.e. a function that measures “how different” a new observation is with respect to the ones in the training set). CP, then, returns a prediction set, called the Conformal Prediction Region (CPR), that contains the “true value” $y_{n+1}$ with high probability $1-\alpha$ (Vovk et al.,, 2005; Shafer and Vovk,, 2008; Vovk,, 2013; Angelopoulos et al.,, 2024). [^2] For extensions, and more modern references, see e.g. (Caprio et al., 2025b, ; Javanmardi et al.,, 2024; Gibbs et al.,, 2023; Barber et al.,, 2023; Papadopoulos,, 2008). In Caprio et al., 2025a (, Section 5.1), the authors show how CP can be seen as a correspondence (i.e. a set-valued function, sometimes referred to as relation (Perrone,, 2024)), thus highlighting a connection between ML and functional analysis (FA, Aliprantis and Border, (2006)). Other such connections have been explored in the literature, see e.g. Korolev, (2022).
This paper builds on the CP-as-a-correspondence intuition to develop a category-theoretic approach to CP, which allows us to expand on how we understand the structural underpinnings of distribution-free uncertainty quantification. [^3] Throughout the paper, we do not use the form “we/our” as a form of pluralis majestatis, but rather to comply to the convention of scientific papers of writing in plural form. We cast CP as a morphism (embedded in a commuting diagram) of two newly-defined categories, and we are also able to frame a CPR as a functor image.
Contributions. This category-theoretic approach grants us three joys, which are briefly summarized next.
  • Quantified uncertainty. CP is not only an uncertainty (ordinal) representation, but also an uncertainty (cardinal) quantification tool: The scholars who improperly referred to CP as an Uncertainty Quantification (UQ) methodology, were right all along.
    • As a byproduct, the user will now be able to see how much the choice of non-conformity measure $\psi$ affects (different types of) uncertainty in the analysis.
  • Conceptual Unification. CP bridges Bayesian, frequentist, and imprecise probabilistic approaches to predictive statistical reasoning, thus formalizing an intuition first put forth by Martin, (2022). This is a conceptual contribution that shows the central (and possibly unifying) role that CP plays in prediction theory.
  • Privacy by Design. Seeing the Conformal Prediction Region as the image of a covariant functor directly benefits AI privacy (Fang et al.,, 2025): (i) Locally added privacy noise leaves coverage guarantees intact, and (ii) In federated learning (Liu et al.,, 2024), each agent need only share a summary object (the credal set $\mathcal{M}$, introduced in the following paragraphs) instead of raw data.
As we can see, our Categorical Conformal Prediction approach exhibits many boons, and little to no banes (minimal assumptions on $Y$, choice of $\psi$, and $\alpha$).
To make the paper largely self-contained, we give the necessary background on CP in Section 2. There, we also provide some concepts from the Imprecise Probability (IP) literature. IPs (Walley,, 1991; Augustin et al.,, 2014; Troffaes and de Cooman,, 2014) are mathematical models that allow the scholar to reason in the presence of uncertainty (as for classical probabilities), and ambiguity (that is, ignorance, full or partial, about the “true” data generating process). [^4] For more modern references, see e.g. (Caprio and Gong,, 2023; Caprio and Mukherjee,, 2023; Caprio and Seidenfeld,, 2023; Caprio et al., 2024b, ; Caprio,, 2024, 2025). Their applications to ML form a lively subject field, see e.g. Caprio et al., 2024a ; Sale et al., (2023, 2024); Caprio et al., 2024c ; Lu et al., (2024); Dutta et al., (2024); Chau et al., (2025). Functional Analysis and Category Theory notions are brushed up in Appendices C and D, respectively. The structure of the paper can be summarized as follows.
In Section 3, we introduce category $\mathbf{UHCont}$. Its objects are topological spaces, and its morphisms are upper hemicontinuous correspondences. Under minimal assumptions on the state space $Y$, on the choice of the credibility level $\alpha$, and on the class $\mathscr{F}$ of non-conformity scores $\psi$, we show that CP is a morphism of $\mathbf{UHCont}$. Viewing CP as a morphism rather than a stand-alone algorithm lets ML engineers plug it into functional-style libraries (e.g. JAX or PyTorch Functorch) and obtain automatic batching, JIT compilation, and differentiability of the score map, while the coverage guarantee remains mathematically intact.
CP is also a morphism of category $\mathbf{WMeas}_{\text{uc}}$ (which we briefly introduce in Section 3, and explore in depth in Appendix A to keep the paper easily readable) whose objects are compact Polish spaces and whose morphisms are weakly measurable, uniformly compact-valued correspondences. We study both $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$ because continuity and measurability are fundamental—but generally non-equivalent—concepts in functional analysis.
Not only is the CP correspondence a morphism of these categories, but it is also part of a commuting diagram in both. As we shall see (in particular in Example 1 and in the discussion thereafter), this discovery implies the intrinsic (cardinal) UQ capabilities of CP.
In Section 4, we show that, under minimal assumption, a diagram that codifies the way Bayesian CP (Fong and Holmes,, 2021), classical CP, and CP-informed imprecise probability theory generate an $\alpha$-level prediction region, commutes asymptotically. This is a strong indicator that CP forms a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning (Martin,, 2022).
Section 5 expands on category-theoretic Imprecise Probability theory (Liell-Cock and Staton,, 2024). We formalize the notion of Imprecise Highest Density Region (Coolen,, 1992)—that can be thought of as a “robust version” of a confidence interval which also accounts for distribution uncertainty—as a (covariant) functor $\mathsf{IHDR}_{\alpha}$, $\alpha\in[0,1]$, between the category of closed and convex sets of finitely additive probabilities on $Y$, and that of subsets of $Y$, both with inclusion morphisms. In turn, the Conformal Prediction Region (CPR) can be seen as the image $\mathsf{IHDR}_{\alpha}(\mathcal{M})$ of such a functor at the set of probabilities $\mathcal{M}$ associated with the CP procedure (Cella and Martin,, 2022). This has e.g. the following promising implication. If we have many sources producing credal sets $\mathcal{M}_{j}$, then privacy-preserving operation applied locally to each credal set $\mathcal{M}_{j}$ propagates automatically through the functor. The resulting prediction regions, then, still satisfy the global coverage guarantee while only the (possibly privatized) summary objects $\mathcal{M}_{j}$, never the raw data, need be shared.
Finally, Section 6 studies the properties of $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$. In particular, we show that (after slightly restricting the former) they are both monoidal, and find faithful functors between them. In addition, we find which of their subcategories admit a monad, give sufficient conditions for CP to be a morphism in such subcategories, and find a faithful functor between them. Section 7 concludes our work, and discusses some open research questions.
2. Background
In this section, we introduce (Full) Conformal Prediction (Section 2.1), and some basic notions of Imprecise Probability Theory (Section 2.2). The reader familiar with any of these concepts can skip the associated part.
2.1. Full Conformal Prediction
This section is an expanded reproduction of Caprio et al., 2025a (, Section 2.1). It is based on Cella and Martin, (2022, Section 4.1), who summarize work by Shafer and Vovk, (2008) and Vovk et al., (2005) on Transductive (or Full) Conformal Prediction.
Suppose that there is an exchangeable process $\mathcal{Y}_{1},\mathcal{Y}_{2},\ldots$ with distribution $\mathfrak{P}$, where each $\mathcal{Y}_{i}$ is a random element taking values in the measurable space $(Y,\Sigma_{Y})$. Recall that a sequence is exchangeable if, for any $k\in\mathbb{N}$ and any permutation $\text{perm}(\cdot)$, the two random vectors $(\mathcal{Y}_{1},\ldots,\mathcal{Y}_{k})$ and $(\mathcal{Y}_{\text{perm}(1)},\ldots,\mathcal{Y}_{\text{perm}(k)})$ have the same joint distribution. This implies that the marginal distributions of the $\mathcal{Y}_{i}$’s are the same. [^5] Exchangeability does not imply independence, so the standard i.i.d. setup is in a sense more restrictive. Note also that we are not assuming any parametric form for the distribution $\mathfrak{P}$.
We want to solve the following statistical problem. Suppose we observe the first $n$ terms of the process, that is, $\mathcal{Y}^{n}=(\mathcal{Y}_{1},\ldots,\mathcal{Y}_{n})$. With this data, and the assumption of exchangeability, the goal is to predict $\mathcal{Y}_{n+1}$ using a method that is valid or reliable in a certain sense.
Let $\mathcal{Y}^{n+1}=(\mathcal{Y}^{n},\mathcal{Y}_{n+1})$ be an $(n+1)$-dimensional vector consisting of the observable $\mathcal{Y}^{n}$ and the yet-to-be-observed value $\mathcal{Y}_{n+1}$. Consider the transform
$\mathcal{Y}^{n+1}\rightarrow T^{n+1}=(T_{1},\ldots,T_{n+1})$
defined by the rule
$T_{i}\coloneqq\psi_{i}\left(\mathcal{Y}^{n+1}\right)\equiv\psi\left(y^{n+1}_{-i},y_{i}\right)$
for all $i\in\{1,\ldots,n+1\}$, where $y^{n+1}_{-i}=y^{n+1}\setminus\{y_{i}\}$ and $\psi:Y^{n}\times Y\rightarrow\mathbb{R}$ is a fixed function that is invariant to permutations in its first vector argument. Function $\psi$, which is called a non-conformity measure, is constructed in such as way that $\psi_{i}(y^{n+1})$ is small if and only if $y_{i}$ agrees with—i.e. is “close to”—a prediction based on the data $y^{n+1}_{-i}$. Or, the other way around, large values $\psi_{i}(y^{n+1})$ suggest that the observation $y_{i}$ is “strange” and does not conform to the rest of the data $y^{n+1}_{-i}$. The key idea is to define $\psi_{i}(y^{n+1})$ in a way that allows us to compare $y_{i}$ to a suitable summary of $y^{n+1}_{-i}$, e.g. $\psi_{i}(y^{n+1})=|\text{mean}(y^{n+1}_{-i})-y_{i}|$, for all $i\in\{1,\ldots,n+1\}$. Notice that transformation $\mathcal{Y}^{n+1}\rightarrow T^{n+1}$ preserves exchangeability.
As the value $\mathcal{Y}_{n+1}$ has not yet been observed, and is actually the prediction target, the above calculations cannot be carried out exactly. Nevertheless, the exchangeability-preserving properties of the transformations described above provide a procedure to rank candidate values $\tilde{y}$ of $\mathcal{Y}_{n+1}$ based on the observed $\mathcal{Y}^{n}=y^{n}$, as shown in Algorithm 1, where $\mathbbm{1}[\cdot]$ denotes the indicator function.
Algorithm 1 Full Conformal prediction (CP)
Initialize: data $y^{n}$, non-conformity measure $\psi$, grid of $\tilde{y}$ values
for each $\tilde{y}$ value in the grid do
  set $y_{n+1}=\tilde{y}$ and write $y^{n+1}=y^{n}\cup\{y_{n+1}\}$;
  define $T_{i}=\psi_{i}(y^{n+1})$, for all $i\in\{1,\ldots,n+1\}$;
  evaluate $\pi(\tilde{y},y^{n})=(n+1)^{-1}\sum_{i=1}^{n+1}\mathbbm{1}[T_{i}\geq T_{n+1}]$;
end for
return $\pi(\tilde{y},y^{n})$ for each $\tilde{y}$ on the grid.
The output of Algorithm 1 is a data-dependent function $\tilde{y}\mapsto\pi(\tilde{y},y^{n})\in[0,1]$ that can be interpreted as a measure of plausibility of the assertion that $\mathcal{Y}_{n+1}=\tilde{y}$, given data $y^{n}$. Vovk et al., (2005) refer to the function $\pi$ as conformal transducer. Conformal transducer $\pi$ plays a key role in the construction of conformal prediction regions (CPRs). For any $\alpha\in[0,1]$, the $\alpha$-level CPR is defined as (Vovk,, 2013, Equation (2))
$\mathscr{R}_{\alpha}^{\psi}(y^{n})\coloneqq\{y_{n+1}\in Y:\pi(y_{n+1},y^{n})>\alpha\},$ (1)
and it satisfies
$P\left[\mathcal{Y}_{n+1}\in\mathscr{R}_{\alpha}^{\psi}(y^{n})\right]\geq 1-\alpha,$ (2)
uniformly in $n$ and in $P$ (Vovk et al.,, 2005). That is, (2) is satisfied for all $n\in\mathbb{N}$ and all exchangeable distributions $P$. Note that (2) holds regardless of the choice of the non-conformity measure $\psi$. However, the choice of this function is crucial in terms of the efficiency of conformal prediction, that is, the size of the prediction regions. A simple visual representation of Full CP is given in Figure 1.
Figure 1. A visual representation of Full CP methodology (Caprio et al., 2025a, , Figure 1). The Conformal Prediction Region is obtained as in (1).
Caprio et al., 2025a (, Section 5) point out how Full CP can be written as a correspondence
$\kappa:[0,1]\times Y^{n}\times\mathscr{F}\rightrightarrows Y,\quad(\alpha,y^{n},\psi)\mapsto\kappa(\alpha,y^{n},\psi),$ (3)
where
[…]
and the image of $\kappa$ is defined as
$\kappa(\alpha,y^{n},\psi)\coloneqq\Bigg{\{}y_{n+1}\in Y:\underbrace{\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}\left[\psi\left(y^{n+1}_{-i},y_{i}\right)\geq\psi\left(y^{n+1}_{-(n+1)},y_{n+1}\right)\right]}_{\eqqcolon\pi(y_{n+1},y^{n})}>\alpha\Bigg{\}}.$ (5)
It is easy to see, then, that $\kappa(\alpha,y^{n},\psi)\equiv\mathscr{R}_{\alpha}^{\psi}(y^{n})$.
Full Conformal Prediction (FCP) is not the only way of carrying out conformal prediction. There exists another type called Inductive (or Split) Conformal Prediction (ICP). It assumes exchangeability and it is used to build the same region $\mathscr{R}_{\alpha}^{\psi}(y^{n})$ in (1) having the same guarantee (2), while being computationally less expensive than FCP. Since ICP produces the same region as FCP, which is what we need in light of the goal of our paper, we do not present ICP in detail. We refer the interested reader to (Papadopoulos,, 2008; Angelopoulos et al.,, 2024).
2.2. Imprecise Probabilities
We now introduce a few imprecise probabilistic concepts that will be used when studying CP in the Category Theory framework.
Let $(Y,\Sigma_{Y})$ be a measurable space, and call $\Delta_{Y}$ the space of finitely additive probabilities on $Y$, endowed with the weak topology. A credal set $\mathcal{M}$ is a weak-closed and convex element of $\Delta_{Y}$ (Levi,, 1980). Throughout the paper, we will denote by $\mathscr{C}\subset 2^{\Delta_{Y}}$ the space of nonempty weak-closed and convex subsets of $\Delta_{Y}$. For any nonempty credal set $\mathcal{M}\in\mathscr{C}$, define the upper probability as the upper envelope of $\mathcal{M}$,
$\overline{P}_{\mathcal{M}}(A)\equiv\overline{P}(A)=\sup_{P\in\mathcal{M}}P(A),\quad A\in\Sigma_{Y}.$
The conjugate lower probability $\underline{P}_{\mathcal{M}}\equiv\underline{P}$ of $\overline{P}$ satisfies $\underline{P}(A)=1-\overline{P}(A^{c})$, for all $A\in\Sigma_{Y}$, thus knowing one immediately gives us the other. In a statistical and ML framework, these concepts allow us to crystallize model ambiguity.
Even in the presence of such ambiguity, we are able to define (a version of) a confidence interval. Fix a credibility level $\alpha\in[0,1]$. Then, the $\alpha$-level Imprecise Highest Density Region (IHDR) of credal set $\mathcal{M}$ is the set $B_{\mathcal{M}}^{\alpha}\subset Y$ such that $\underline{P}(B_{\mathcal{M}}^{\alpha})=1-\alpha$, and its “size is minimal” (Coolen,, 1992). More formally (Caprio et al., 2025a, , Equation (11)),
$B_{\mathcal{M}}^{\alpha}=\bigcap\left\{{A\in\Sigma_{Y}:\underline{P}(A)\geq 1-\alpha}\right\}.$
Notice that $B_{\mathcal{M}}^{\alpha}$ can be written as the image of the following correspondence
$\mathsf{IHDR}:[0,1]\times\mathscr{C}\rightrightarrows Y,\quad(\alpha,\mathcal{M})\mapsto\mathsf{IHDR}(\alpha,\mathcal{M})\coloneqq\bigcap\left\{{A\in\Sigma_{Y}:\underline{P}(A)\geq 1-\alpha}\right\}.$ (6)
Conformal Prediction is associated with a special case of upper probability $\overline{P}$, called a possibility function (Cuzzolin,, 2020). The latter is an upper probability for which there exists a function $\pi:Y\rightarrow[0,1]$ such that
  • $\sup_{y\in Y}\pi(y)=1$;
  • $\overline{P}(A)=\sup_{y\in A}\pi(y)$, $A\in\Sigma_{Y}$.
Cella and Martin, (2022) show that, under the so-called consonance assumption, i.e. if $\sup_{y_{n+1}\in Y}\pi(y_{n+1},y^{n})=1$, then conformal transducer $\pi$ induces a possibility function $\overline{\Pi}(A)\coloneqq\sup_{y_{n+1}\in A}\pi(y_{n+1},y^{n})$, for all $A\in\Sigma_{Y}$, and in turn a credal set $\mathcal{M}(\overline{\Pi})\coloneqq\{P:P(A)\leq\overline{\Pi}(A),\forall A\in\Sigma_{Y}\}$. Such a credal set can be seen as the image of the correspondence $\mathsf{CRED}:Y^{n}\times\mathscr{F}\rightrightarrows\Delta_{Y}$, where $\mathscr{F}$ is the collection of functions in (4), and $(y^{n},\psi)\mapsto\mathsf{CRED}(y^{n},\psi)\coloneqq$
$\Bigg{\{}P\in\Delta_{Y}:P(A)\leq\underbrace{\sup_{y_{n+1}\in A}\left(\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}\left[\psi\left(y^{n+1}_{-i},y_{i}\right)\geq\psi\left(y^{n+1}_{-(n+1)},y_{n+1}\right)\right]\right)}_{\equiv\sup_{y_{n+1}\in A}\pi(y_{n+1},y^{n})\eqqcolon\overline{\Pi}(A)}\text{, }\forall A\in\Sigma_{Y}\Bigg{\}}.$ (7)
The consonance assumption is needed to ensure that $\mathsf{CRED}(y^{n},\psi)\in\mathscr{C}$; The elements of $\mathsf{CRED}(y^{n},\psi)$ are characterized in Martin, (2025, Theorem 1). In the remainder of the paper, we sometimes write $\pi(y,y^{n},\psi)\equiv\pi_{\psi}(y,y^{n})$ for notational convenience (we drop the $n+1$ subscript for $y_{n+1}$), and to explicitly acknowledge that the conformal transducer $\pi$ depends on the choice of the non-conformity measure $\psi$.
We also consider a slight modification of $\mathsf{CRED}$, that will significantly simplify our proving endeavors in Sections 3 and 4, and in Appendix A. We denote it by $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\text{, }\forall A\in\Sigma_{Y}\text{ closed}\}$. As we can see, $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)\supseteq\mathsf{CRED}(y^{n},\psi)=\mathcal{M}(\overline{\Pi})$, because it imposes less constraints on the $P$’s: They only have to be setwise dominated on closed sets. The upper probability $\overline{\Pi}$ and its conjugate lower probability $\underline{\Pi}$ (the only characteristics of the credal sets that we use) stay the same for both $\mathsf{CRED}$ and $\mathsf{CRED}_{\text{cl}}$: This justifies our simplifying choice. To avoid notation complications, in the remainder of this work we do not distinguish between $\mathsf{CRED}(y^{n},\psi)$ and $\mathsf{CRED}_{\text{cl}}(y^{n},\psi)$.
A visual representation of the data $\to$ credal set $\to$ IHDR route is given in Figure 2.
Figure 2. A visual representation of the two-step procedure that first extracts credal set $\mathcal{M}(\overline{\Pi})$ from the training set $y^{n}$ as in (7), and then computes the IHDR as in (6) (Caprio et al., 2025a, , Figure 1).
In addition, Caprio et al., 2025a (, Propositions 5 and 6) show that the $\alpha$-level IHDR associated with $\mathcal{M}(\overline{\Pi})$, computed using the conjugate lower probability $\underline{\Pi}$ of $\overline{\Pi}$, corresponds to the Conformal Prediction Region $\mathscr{R}_{\alpha}^{\psi}(y^{n})$; It enjoys the same probabilistic guarantee (2).
3. Conformal Prediction as an Upper Hemicontinuous Morphism
This section introduces category UHCont, and shows that, under minimal assumptions, the Conformal Prediction correspondence $\kappa$ is a morphism of UHCont, embedded in a commuting diagram. The following is inspired by Perrone, (2024, Example 1.1.14).
Definition 1 (Structure UHCont).
Define a structure $\mathbf{UHCont}$ as follows,
  • Objects: Topological spaces $(X,\tau_{X})$.
  • Morphisms: Upper hemicontinuous correspondences $\Phi:X\rightrightarrows Y$ (see Definition 25 in Appendix C).
Theorem 2 (UHCont is a Category).
The structure $\mathbf{UHCont}$ defines a category.
Proof.
We verify the axioms of a category. Recall that the composition $\Psi\circ\Phi$ of two correspondences $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ is given by $(\Psi\circ\Phi)(x)\coloneqq\bigcup_{y\in\Phi(x)}\Psi(y)$.
(1) Identity morphisms. For each object $X$, define $\mathrm{id}_{X}(x)\coloneqq\{x\}$. Let $V\subseteq X$ be open and suppose $x_{0}\in X$ satisfies $\mathrm{id}_{X}(x_{0})=\{x_{0}\}\subseteq V$. Then, since $x_{0}\in V$ and $V$ is open, there exists an open neighborhood $U=V$ of $x_{0}$ such that for all $x\in U$, $\mathrm{id}_{X}(x)=\{x\}\subseteq V$. Hence, $\mathrm{id}_{X}$ is upper hemicontinuous.
(2) Associativity of composition. Let $\Phi:X\rightrightarrows Y$, $\Psi:Y\rightrightarrows Z$, and $\Theta:Z\rightrightarrows W$. Then for all $x\in X$,
$((\Theta\circ\Psi)\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\bigcup_{z\in\Psi(y)}\Theta(z)=\bigcup_{z\in(\Psi\circ\Phi)(x)}\Theta(z)=(\Theta\circ(\Psi\circ\Phi))(x)$
by the associativity of union. Thus, composition is associative.
(3) Closure under composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be upper hemicontinuous correspondences. By Aliprantis and Border, (2006, Theorem 17.23), the composition $\Psi\circ\Phi$ is also upper hemicontinuous. Thus, the composition of morphisms is again a morphism.
(4) Identity laws. For any morphism $\Phi:X\rightrightarrows Y$ and any $x\in X$,
$(\Phi\circ\mathrm{id}_{X})(x)=\bigcup_{z\in\mathrm{id}_{X}(x)}\Phi(z)=\Phi(x),\quad(1_{Y}\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\{y\}=\Phi(x).$
Hence, $\mathrm{id}_{X}$ acts as a left and right identity.
Therefore, $\mathbf{UHCont}$ satisfies the axioms of a category. ∎
We now show that, under minimal assumptions, the (Full) Conformal Prediction correspondence $\kappa$ in (3) is a morphism of $\mathbf{UHCont}$, and that it is part of a commutative diagram involving the IHDR and the credal set correspondences. Fix any $n\in\mathbb{N}$, and consider the following conditions,
  1. (i)
    $Y$ is compact Hausdorff and $\Sigma_{Y}=\mathcal{B}(Y)$ is its Borel $\sigma$-field.
  2. (ii)
    Each $\psi\in\mathscr{F}$ is jointly continuous on $Y^{n}\times Y$; We can then write $\mathscr{F}\subset C(Y^{n}\times Y)$, the space of jointly continuous real-valued functions on $Y^{n}\times Y$. $\mathscr{F}$ is endowed with the uniform ($\sup$-norm) topology.
  3. (iii)
    No-tie parameter: The credibility level $\alpha$ satisfies $\alpha\notin S_{n+1}\coloneqq\{0,\frac{1}{n+1},\frac{2}{n+1},\ldots,\frac{n}{n+1},1\}$.
  4. (iv)
    $\Delta_{Y}$ is endowed with the weak topology, and $\mathscr{C}\subset 2^{\Delta_{Y}}$ is endowed with the Vietoris topology (see Definition 28 in Appendix C).
The following is a famous classical result in set-valued functional analysis (Aliprantis and Border,, 2006, Theorem 17.11).
Lemma 3 (Closed Graph Theorem).
A correspondence with compact Hausdorff range space has closed graph if and only if it is upper hemicontinuous and closed-valued.
Another important lemma that sets the table for our main result is the following. To improve readability, we defer its proof to Appendix B.
Lemma 4 (Upper Semicontinuity of the Upper Envelope $\overline{\Pi}$).
If (i) and (ii) hold, then the map $(y^{n},\psi)\mapsto\overline{\Pi}_{(y^{n},\psi)}(A)=\sup_{y\in A}\pi(y,y^{n},\psi)$ is upper semicontinuous, for all closed sets $A\in\Sigma_{Y}$.
We are now ready for the main result of this section.
Theorem 5 (The Conformal Prediction Diagram Commutes, $\mathbf{UHCont}$).
Assume that (i)-(iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram
$Y^{n}\times\mathscr{F}$$\mathscr{C}$$\Sigma_{Y}$$\scriptstyle{\mathsf{CRED}}$$\scriptstyle{\kappa_{\alpha}}$$\scriptstyle{\mathsf{IHDR}_{\alpha}}$
commutes in $\mathbf{UHCont}$.
In the Full CP Diagram that factorizes $\kappa_{\alpha}=\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ inside $\mathbf{UHCont}$, for all $\alpha$ satisfying (iii), we indicated $\mathscr{C}$ as the image set of $\mathsf{CRED}$, and $\Sigma_{Y}$ as the image set of $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$. This is an immediate consequence of writing our three correspondences in their functional form $\kappa_{\alpha}^{\text{funct}}:Y^{n}\times\mathscr{F}\to\Sigma_{Y}$, $\mathsf{IHDR}_{\alpha}^{\text{funct}}:\mathscr{C}\to\Sigma_{Y}$, and $\mathsf{CRED}^{\text{funct}}:Y^{n}\times\mathscr{F}\to\mathscr{C}$, where the latter holds under consonance. Notice also that $\Sigma_{Y}=\mathcal{B}(Y)$ by (i).
Let us also emphasize that we only use $\mathsf{IHDR}_{\alpha}$ through its restriction to the image set $\operatorname{Im}(\mathsf{CRED})\subset\mathscr{C}$ (which is trivially an object of $\mathbf{UHCont}$). We denote it by $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}\equiv\mathsf{IHDR}_{\alpha}\mid_{\operatorname{Im}(\mathsf{CRED})}:\operatorname{Im}(\mathsf{CRED})\rightrightarrows Y$. We do not claim that the stand alone correspondence $\mathsf{IHDR}_{\alpha}:\mathscr{C}\rightrightarrows Y$ is upper hemicontinuous on all of $\mathscr{C}$; This is not needed for the diagram to commute inside $\mathbf{UHCont}$, and may fail without additional assumptions.
Proof of Theorem 5.
Fix any $n\in\mathbb{N}$. We first show that $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ have a closed graph.
(i) $\kappa_{\alpha}$. By (ii) and its definition, the conformal transducer $\pi(\cdot,y^{n},\psi):Y\to[0,1]$ is upper semicontinuous. Pick any $\alpha$ that avoids the grid $S_{n+1}$, i.e. $\alpha\in\Lambda\coloneqq[0,1]\setminus S_{n+1}$. Let
$\beta(\alpha)\equiv\beta\coloneqq\min\Bigl{\{}\tfrac{k}{n+1}\in S_{n+1}:\tfrac{k}{n+1}>\alpha\Bigr{\}},\quad\text{so }0<\alpha<\beta\leq 1.$ (8)
Then, define
$G\coloneqq\bigl{\{}(\alpha,y^{n},\psi,y):\pi(y,y^{n},\psi)\geq\beta\bigr{\}}\subset[0,1]\times Y^{n}\times\mathscr{F}\times Y.$
This is a superlevel set of $\pi(\cdot,y^{n},\psi)$, and since $\pi$ is upper semicontinuous, $G$ is closed. In turn, $\operatorname{Gr}(\kappa_{\alpha})=G$ is closed.
(ii) $\mathsf{CRED}$. For every closed $A\in\Sigma_{Y}$, set
$G_{A}\coloneqq\Bigl{\{}(y^{n},\psi,P):P(A)\leq\overline{\Pi}_{(y^{n},\psi)}(A)\Bigr{\}}.$
Weak-continuity of $P\mapsto P(A)$ and upper semicontinuity of the right-hand side (see Lemma 4) make $G_{A}$ closed. Because $\operatorname{Gr}(\mathsf{CRED})=\bigcap_{A\in\Sigma_{Y}\text{, }A\text{ closed}}G_{A}$, the graph is closed.
(iii) $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$. Under the consonance assumption, for each $(y^{n},\psi)$, the credal set $\mathsf{CRED}(y^{n},\psi)$ is generated by the possibility $\overline{\Pi}_{(y^{n},\psi)}(A)=\sup_{y\in A}\pi_{\psi}(y,y^{n})$, $A\in\Sigma_{Y}$. Recall that, for any $\alpha\notin S_{n+1}$, we defined $\beta(\alpha)\coloneqq\min\{\tfrac{k}{n+1}\in S_{n+1}:\tfrac{k}{n+1}>\alpha\}$. Then, by (Caprio et al., 2025a, , Proposition 5), we have that
$\mathsf{IHDR}_{\alpha}\big{(}\mathsf{CRED}(y^{n},\psi)\big{)}=\Bigl{\{}y\in Y:\pi_{\psi}(y,y^{n})\geq\beta(\alpha)\Bigr{\}}.$
By Lemma 4, $y\mapsto\pi_{\psi}(y,y^{n})$ is upper semicontinuous on $Y$, so the $\alpha$-cut $\{\pi_{\psi}(\cdot,y^{n})\geq\beta(\alpha)\}$ is closed in $Y$. Moreover, since $(y,y^{n},\psi)\mapsto\pi_{\psi}(y,y^{n})$ is upper semicontinuous, the graph
$\operatorname{Gr}\bigl{(}\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}\bigr{)}=\Bigl{\{}\bigl{(}y^{n},\psi,y\bigr{)}\in Y^{n}\times\mathscr{F}\times Y:\pi_{\psi}(y,y^{n})\geq\beta(\alpha)\Bigr{\}}$
is closed. Therefore, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ has closed graph.
Notice also that $Y$ is compact Hausdorff by (i), and that $\Delta_{Y}$ is compact by (iv) and Banach-Alaoglu, [^6] To be precise, let $B_{b}(Y,\Sigma_{Y})$ be the Banach space of bounded $\Sigma_{Y}$-measurable functions with the sup norm. Its (isometric) dual is $ba(\Sigma_{Y})$, the space of bounded finitely additive signed measures (charges) on $\Sigma_{Y}$, with the variation norm. Equip $ba(\Sigma_{Y})$ with the weak topology. Then, $\Delta_{Y}$ is a weak-closed subset of the weak-compact unit ball of $ba(\Sigma_{Y})$, hence $\Delta_{Y}$ is compact Hausdorff in this topology. and Hausdorff because it is a subset of a dual Banach space equipped with its weak topology.
It follows that $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$ are upper hemicontinuous by Lemma 3. In turn, they are morphisms of $\mathbf{UHCont}$.
$Y$ is a topological space from (i); $\operatorname{Im}(\mathsf{CRED})$ and $\mathscr{C}$ are topological spaces by (iv); $Y^{n}\times\mathscr{F}$ is a topological space in the product topology by (ii). Hence, these spaces are all objects of $\mathbf{UHCont}$.
The fact that the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}$, then, is an immediate consequence of Caprio et al., 2025a (, Proposition 5). It states that, for all $\alpha\in[0,1]$, $y^{n}\in Y^{n}$, and $\psi\in\mathscr{F}$, the Conformal Prediction Region $\kappa_{\alpha}(y^{n},\psi)$ corresponds to the $\alpha$-level IHDR of the credal set $\mathsf{CRED}(y^{n},\psi)$, in formulas $\mathsf{IHDR}_{\alpha}(\mathsf{CRED}(y^{n},\psi))$. Of course, here we restrict to $\alpha\in[0,1]\setminus S_{n+1}$. ∎
Let us now give an example in which conditions (i) and (ii) are satisfied. We do not verify (iii), as it is the user that chooses $\alpha$, so the no-tie parameter will always be met, and (iv), because it is a technical assumption that should not worry the practitioner.
Example 1 (An ML-Inspired Setting).
Let $Y=[0,1]^{d}$ with its Euclidean topology (compact and Hausdorff), and let $\varphi_{\theta}\colon Y\to\mathbb{R}^{m}$ be a feed-forward neural network whose activation functions are continuous (e.g. ReLU). Hence, $\varphi_{\theta}$ is continuous in both its parameters $\theta$ and input $y$.
Non-Conformity Measure. For a training set $y^{n}=(y_{1},\dots,y_{n})\in Y^{n}$ and a candidate point $y\in Y$, define
$\psi_{\theta}(y^{n},y)=-\biggl{\|}\varphi_{\theta}(y)-\frac{1}{n}\sum_{i=1}^{n}\varphi_{\theta}(y_{i})\biggr{\|}_{2}^{2}.$ (9)
  • Permutation invariance: The sample mean is symmetric in $y_{1},\dots,y_{n}$.
  • Joint continuity: Composition, finite sums, Euclidean norm, and squaring preserve continuity; Thus $(y^{n},y)\mapsto\psi_{\theta}(y^{n},y)$ is jointly continuous on $Y^{n}\times Y$.
Hence, $\psi_{\theta}\in\mathscr{F}$. Notice that consonance, i.e. $\sup_{y\in Y}\pi(y,y^{n})=1$, does not automatically hold in this example. To enforce it, it is enough to consider the normalized conformal transducer $\tilde{\pi}(y,y^{n})\coloneqq\frac{\pi(y,y^{n})}{\sup_{y\in Y}\pi(y,y^{n})}$ (Cella and Martin,, 2021, Section 7), which is still upper semicontinuous because the denominator is a strictly positive constant for fixed $y^{n}$, and it returns the same CP Region (after the corresponding rescaling of $\alpha$).
Machine Learning Relevance. This example is interesting because of the following implications for Machine Learning.
  1. (1)
    Few-shot / prototype learning: (9) is the squared Euclidean distance to the class prototype $\text{mean}\varphi_{\theta}(y_{i})$.
  2. (2)
    Out-Of-Distribution (OOD) detection: The same distance acts as a learned Mahalanobis-style OOD score.
  3. (3)
    Heteroscedastic conformal regression: If $\varphi_{\theta}$ is the penultimate layer of a regression network, (9) yields locally adaptive, yet still valid, prediction sets.
Hence, this choice of $\psi_{\theta}$ simultaneously fulfills joint continuity and practical usefulness in modern ML workflows. $\diamond$
More in general, for the ML practitioner, it will be enough to
  • check that $Y$ is compact (always satisfied in classification problems and easy to enforce in regression ones), so that (i) is met,
  • pick a non-conformity score $\psi$ that is invariant to permutations in its first argument and jointly continuous in its arguments, so that (ii) is met,
  • pick $\alpha$ outside of $S_{n+1}$, so that (iii) is met,
  • check that the consonance assumption is met, or enforce it e.g. via a transformation, as we did in Example 1,
to be sure that the conditions of Theorem 5 are met. Condition (iv) is technical, and should not worry them.
Theorem 5 clarifies the nature of Conformal Prediction: Although often billed as an uncertainty quantification tool, CP furnishes a representation of predictive uncertainty, but not a quantification in the cardinal sense. The distinction is subtle yet essential: CP Regions depend only on the ranks of the non-conformity scores: For any strictly increasing transform $f$, $\kappa_{\alpha}(\psi)=\kappa_{\alpha}(f\circ\psi)$. Thus, while $\mathrm{diam}(\mathscr{R})$ is numerically well-defined once a metric on $Y$ is fixed, [^7] A similar argument holds if we consider the volume in place of the diameter of the CPR. there is no canonical cardinal scale of “amount of uncertainty” attached to $\mathscr{R}$. In particular,
$\mathrm{diam}(\mathscr{R}_{1})=3\mathrm{diam}(\mathscr{R}_{2})\not\Rightarrow\text{``$\mathscr{R}_{1}$ is three times more uncertain than $\mathscr{R}_{2}$''.}$
CP natively supports ordinal comparisons (larger/smaller regions), not cardinal arithmetic.
The commuting factorization $\kappa_{\alpha}=\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}$, however, reveals a hidden layer. Indeed, computing a CP region is equivalent to (i) forming the credal set $\mathcal{M}(\overline{\Pi})$ and (ii) extracting its IHDR. Because $\mathcal{M}(\overline{\Pi})$ is a credal set (i.e. it “lives” in the space of probability measures), existing metrics—e.g. those in Abellán et al., (2006); Hüllermeier and Waegeman, (2021); Javanmardi et al., (2023); Hofman et al., (2024); Chau et al., (2025)—endow it with a cardinal scale that is able to quantify (different types of) uncertainty, in particular, reducible (epistemic) and irreducible (aleatoric) (Javanmardi et al.,, 2024, 2025; Cabezas et al.,, 2025; Stutts et al.,, 2024; Hanselle et al.,, 2025). This will be “marginal uncertainty”, [^8] That is, averaged over the possible realization $y^{n}$. thus reflecting the marginal guarantee of the Conformal Prediction Region (see equation (2)). We conjecture that the analyst-controlled choices—the significance level $\alpha$, which fixes the tolerated error, and the non-conformity score $\psi\in{\mathscr{F}}$, which embodies modeling assumptions—are primarily associated with epistemic uncertainty. On the other hand, aleatoric uncertainty stems from the intrinsic variability of the exchangeable process that we study, encoded in $\mathfrak{P}$.
Thus, Conformal Prediction’s capacity for predictive uncertainty quantification is not an add-on, but a built-in consequence of the categorical factorization resulting from Theorem 5’s commuting diagram. The scholars who have been calling Conformal Prediction a UQ method were right all along (under the stated assumptions), and Theorem 5 explains why. This expands the work initiated by Cella and Martin, (2022, 2023); Caprio et al., 2025a , who found UQ capabilities of CP methods through an imprecise probabilistic route.
Next, we present a remark that shows that conditions (i)-(iv) cannot be lightened.
Remark 1 (Minimality of Hypotheses (i)(iv)).
We prove that each hypothesis is indispensable for the simultaneous upper hemicontinuity (u.h.c.) of the three correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$. For each clause, we remove exactly one hypothesis and exhibit a single parameter value at which at least one of the three maps fails to be u.h.c. [^9] We use u.h.c. for both “upper hemicontinuity” and “upper hemicontinuous”. Throughout, all remaining hypotheses (plus consonance) continue to hold.
  1. (i)
    Dropping (i) (compactness of $Y$). Let $Y=\mathbb{R}$ with its usual (non-compact) topology, take $n=1$, and define a continuous, permutation-invariant score that depends only on the first argument,
    $\psi(y^{1},y)\coloneqq g(y^{1}),\qquad g(x)=\begin{cases}1-e^{x},&x\leq 0,\\ x,&0<x<1,\\ 1,&x\geq 1.\end{cases}$
    Fix the closed set $A\coloneqq(-\infty,0]$ and write $x$ for $y^{1}$. For $n=1$ the conformal transducer satisfies $\pi(y,x)=\tfrac{1}{2}\bigl{(}1+\mathbbm{1}[g(y)\geq g(x)]\bigr{)}.$ Hence
    $\overline{\Pi}_{x}(A)=\sup_{y\leq 0}\pi(y,x)=\begin{cases}1,&g(x)<\sup_{y\leq 0}g(y)=1,\\ \tfrac{1}{2},&g(x)\geq 1.\end{cases}$
    Since $\sup_{y\leq 0}g(y)=1$ is not attained on $A$ while $g(1)=1$, we obtain $\overline{\Pi}_{x}(A)=\mathbbm{1}[x<1]+\tfrac{1}{2}\mathbbm{1}[x\geq 1],$ which has an upward limsup at $x=1$, $\limsup_{x\uparrow 1}\overline{\Pi}_{x}(A)=1>\tfrac{1}{2}=\overline{\Pi}_{1}(A).$ Thus the envelope is not upper semicontinuous at $x=1$. Consequently, $\mathsf{CRED}$ fails to be u.h.c. without compactness of $Y$.
  2. (ii)
    Dropping (ii) (joint continuity of $\psi$). Let $Y=[0,1]$ (compact Hausdorff) and $n=1$. Define the permutation-invariant but discontinuous score
    $\psi(y^{1},y)\coloneqq g(y^{1}),\qquad g(x)=\mathbbm{1}[x\in\mathbb{Q}].$
    Fix $\alpha=\tfrac{3}{4}$ so that $\beta(\alpha)=1$, and take $y^{1}=x$. Then for $n=1$, $\pi(y,x)=\tfrac{1}{2}\bigl{(}1+\mathbbm{1}[g(y)\geq g(x)]\bigr{)}.$ If $x\in\mathbb{Q}$, $g(x)=1$ and
    $\kappa(\alpha,x,\psi)=\{y\in[0,1]:g(y)\geq 1\}=[0,1]\cap\mathbb{Q}.$
    If $x\notin\mathbb{Q}$, then $g(x)=0$ and $\pi(y,x)=1$ for all $y$, hence $\kappa(\alpha,x,\psi)=[0,1].$ Let $t\in(0,1)\setminus\mathbb{Q}$ be fixed and set $V\coloneqq[0,1]\setminus\{t\}$, which is open in the subspace topology and satisfies $\kappa(\alpha,x,\psi)\subset V$ when $x\in\mathbb{Q}$. For any neighborhood $U$ of a rational $x$, there exists $x^{\prime}\in U\setminus\mathbb{Q}$ with $\kappa(\alpha,x^{\prime},\psi)=[0,1]\not\subset V$. Thus $\kappa$ is not upper hemicontinuous at rational $x$. Joint continuity of $\psi$ is therefore necessary.
  3. (iii)
    Dropping (iii) (no-tie rule). Let $\psi\in\mathscr{F}$ be the constant score $\psi(y^{n},y)\equiv 0$. Then for every $y,y^{n}$, all nonconformity scores are equal, hence $\pi(y,y^{n})=1$. Take $\alpha=1\in S_{n+1}$ (violating (iii)). Because $\kappa$ uses the strict inequality $\pi>\alpha$,
    $\kappa(1,y^{n},\psi)=\emptyset.$
    For any sequence $\alpha_{k}\uparrow 1$ with $\alpha_{k}<1$, we have $\kappa(\alpha_{k},y^{n},\psi)=Y$ for all $k$. Therefore $\kappa$ is not upper hemicontinuous at $\alpha=1$; The no-tie parameter restriction is essential.
  4. (iv)
    Dropping (iv) (topology on $\mathscr{C}$). Endow $\Delta_{Y}$ with the weak topology, but equip $\mathscr{C}$ with the lower Vietoris (hit) topology, generated by the subbasic sets $\langle U\rangle_{-}\coloneqq\{\mathcal{M}\in\mathscr{C}:\mathcal{M}\cap U\neq\emptyset\}$, where $U\subset\Delta_{Y}$ is weak open. Let $Y=[0,1]$ (so (i) holds) and fix $c\in(0,1]$. Set $A\coloneqq\{0\}$, $P_{0}\coloneqq\delta_{0}$, $P_{c}\coloneqq\delta_{c}$, and define
    $\mathcal{M}_{k}\coloneqq\operatorname{co}(\{P_{0},P_{c}\})\in\mathscr{C}\quad(k\geq 1),\qquad\mathcal{M}_{\infty}\coloneqq\{P_{0}\}\in\mathscr{C},$
    where $\text{co}(\cdot)$ denotes the convex hull operator. Then, $\mathcal{M}_{k}\to\mathcal{M}_{\infty}$ in the lower Vietoris topology: For every weak open neighborhood $U$ of $P_{0}$ we have $P_{0}\in\mathcal{M}_{k}\cap U$ for all $k$. For any $E\in\Sigma_{Y}$,
    $\underline{P}_{\mathcal{M}_{k}}(E)=\inf_{P\in\mathcal{M}_{k}}P(E)=\min\{P_{0}(E),P_{c}(E)\},\qquad\underline{P}_{\mathcal{M}_{\infty}}(E)=P_{0}(E).$
    Fix any $\alpha^{\star}\in(0,1)\setminus S_{n+1}$, so $1-\alpha^{\star}\in(0,1)$. Then
    $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{\infty})=\bigcap\{E\in\Sigma_{Y}:P_{0}(E)\geq 1-\alpha^{\star}\}=\{0\},$
    whereas
    $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{k})=\bigcap\{E\in\Sigma_{Y}:\min(1_{\{0\}\subset E},1_{\{c\}\subset E})\geq 1-\alpha^{\star}\}=\{0,c\}\text{, }k\geq 1.$
    Let $V$ be any open neighborhood of $\{0\}$ with $c\notin V$ (e.g. $V=[0,\varepsilon)$ with $0<\varepsilon<c$). Then $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{\infty})\subset V$, but $\mathsf{IHDR}_{\alpha^{\star}}(\mathcal{M}_{k})=\{0,c\}\nsubseteq V$ for all $k$. Therefore $\mathsf{IHDR}_{\alpha^{\star}}$ is not upper hemicontinuous at $\mathcal{M}_{\infty}$ when $\mathscr{C}$ is endowed with the lower Vietoris topology. Hence the (full) Vietoris topology in (iv) is indispensable.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). In short, removing any one of (i)-(iv) (while keeping the rest, plus consonance) makes at least one of $\kappa$, $\mathsf{CRED}$, or $\mathsf{IHDR}$ non-u.h.c.; The hypotheses are collectively minimal. $\triangle$
An in-depth study of some of the properties of $\mathbf{UHCont}$ is carried out in Section 6. Now, instead, we briefly introduce structure $\mathbf{WMeas}_{\text{uc}}$, which will be explored in depth in Appendix A. There, we prove that it is indeed a category, and that the Full CP Diagram commutes in $\mathbf{WMeas}_{\text{uc}}$ as well, under essentially the same conditions of Theorem 5.
Definition 6 (Structure $\mathbf{WMeas}_{\text{uc}}$).
Define a structure $\mathbf{WMeas}_{\text{uc}}$ as follows,
  • Objects: Compact Polish spaces $(X,\Sigma_{X})$, with $\Sigma_{X}=\mathcal{B}(X)$, the Borel $\sigma$-algebra.
  • Morphisms: Weakly measurable (see Definition 26 in Appendix C), uniformly compact-valued (see Definition 27 in Appendix C) correspondences $\Phi:X\rightrightarrows Y$.
The Full Conformal Prediction Diagram commuting when we focus either on the continuity or on the measurability aspects of the conformal prediction correspondence $\kappa$ further testifies that, under our assumptions, uncertainty quantification is a built-in feature of Conformal Prediction methods.
4. Unifying Bayesian, Frequentist, and Imprecise Reasoning
In this section, we show how a diagram involving Bayesian, Conformal, and Imprecise Probabilistic prediction methods, commutes asymptotically in category $\mathbf{UHCont}$. Since the diagram commuting is an extension of the Full CP Diagram in Theorem 5, we retain assumptions (i)-(iv).
Let $(\Theta,\Sigma_{\Theta})$ be a measurable parameter space, and call $\lambda_{\Theta}$ a fixed $\sigma$-finite Borel measure on $(\Theta,\Sigma_{\Theta})$. Call $\Delta_{\Theta}^{\mathrm{dens}}$ the space of finitely additive probability measures on $\Theta$ admitting a density with respect to $\lambda_{\Theta}$ (Bell and Hagood,, 1988). [^10] This requirement essentially restricts us to focus on countably additive probabilities, see e.g. also here. Endow $\Delta_{\Theta}^{\mathrm{dens}}$ with the $L^{1}(\lambda_{\Theta})$ topology. Denote by $\Delta_{Y}^{\text{dens}}$ the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$. Consider now the following conditions,
  1. (v)
    Measurability. There exist de Finetti kernels $k_{\iota}:\Theta\times Y^{\iota}\to\mathbb{R}_{+}$, $\iota\in\{n,n+1\}$, such that, under parameter $\theta$, the joint density of $y^{\iota}$ is $k_{\iota}(\theta,y^{\iota})$ (a density on $Y^{\iota}$ with respect to $\lambda_{Y}^{\otimes\iota}$). For each $\iota\in\{n,n+1\}$ and each $y^{\iota}\in Y^{\iota}$, $\theta\mapsto k_{\iota}(\theta,y^{\iota})$ is $\Sigma_{\Theta}$-measurable, and for each $\theta\in\Theta$, $y^{\iota}\mapsto k_{\iota}(\theta,y^{\iota})$ is Borel. The posterior predictive is
    $p_{P}(y_{n+1}\mid y^{n})=\frac{\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))P(\mathrm{d}\theta)}{\int_{\Theta}k_{n}(\theta,y^{n})P(\mathrm{d}\theta)},\quad P\in\Delta_{\Theta}^{\mathrm{dens}}.$
  2. (vi)
    Pointwise $y$-continuity a.e. For $\lambda_{\Theta}$-almost all $\theta$, the maps $y^{\iota}\mapsto k_{\iota}(\theta,y^{\iota})$ are continuous on $Y^{\iota}$, for $\iota\in\{n,n+1\}$.
  3. (vii)
    Integrable envelope. There exists $M(\theta)\equiv M\in L^{\infty}(\lambda_{\Theta})$ with $k_{\iota}(\theta,y^{\iota})\leq M(\theta)$ for every $y^{\iota}\in Y^{\iota}$, $\iota\in\{n,n+1\}$, and $\lambda_{\Theta}$-almost all $\theta$.
  4. (viii)
    Positive marginal likelihood. For every $y^{n}\in Y^{n}$ and every admissible prior probability measure $P\in\Delta_{\Theta}^{\mathrm{dens}}$,
    $D_{P}(y^{n})\coloneqq\int_{\Theta}k_{n}(\theta,y^{n})P(\text{d}\theta)=\int_{\Theta}k_{n}(\theta,y^{n})p(\theta)\lambda_{\Theta}(\text{d}\theta)>0.$
For $P\in\Delta_{\Theta}^{\mathrm{dens}}$ having density $p$, and data $y^{n}\in Y^{n}$, set
$p_{P}\left(y_{n+1}\mid y^{n}\right)\coloneqq\frac{N_{P}(y_{n+1},y^{n})}{D_{P}(y^{n})},$
where $N_{P}(y_{n+1},y^{n})\coloneqq\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))p(\theta)\lambda_{\Theta}(\text{d}\theta)$. By (viii), it is well-defined.
Because $y\mapsto p_{P}(y\mid y^{n})$ is continuous (we prove this in Lemma 7) and $Y^{n+1}$ is compact by (i), the natural codomain for such a map is the Banach space $C(Y^{n}\times Y)$ with norm $\|f\|_{\infty}\coloneqq\sup_{(y^{n},y)\in Y^{n}\times Y}|f(y^{n},y)|$. Define
$\mathsf{BCP}:Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}\rightarrow Y^{n}\times C(Y^{n}\times Y),\quad\mathsf{BCP}\bigl{(}y^{n},P\bigr{)}\coloneqq\bigl{(}y^{n},\psi_{(y^{n},P)}\bigr{)},$ (10)
where
$\psi_{(y^{n},P)}(y_{n+1})\coloneqq-p_{P}(y_{n+1}\mid y^{n}).$
Here we abuse notation, and write $\psi_{(y^{n},P)}(y_{n+1})\equiv\psi_{P}(y^{n},y_{n+1})$ to highlight the dependence of the non-conformity score on the collected dataset $y^{n}$. We call $\mathsf{BCP}$ the function in (10), an acronym for Bayesian Conformal Prediction (Fong and Holmes,, 2021). Next, we have an important lemma, that we prove in Appendix B.
Lemma 7.
Under (i), (v)-(viii), the following hold for every fixed $n\in\mathbb{N}$,
  • For every $P\in\Delta_{\Theta}^{\mathrm{dens}}$, the predictive map $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ is jointly continuous on $Y^{n}\times Y=Y^{n+1}$;
  • The map $\mathsf{BCP}$ is continuous when $C(Y^{n}\times Y)\equiv C(Y^{n+1})$ is equipped with the sup-norm.
Conditions (v)-(viii) are indispensable for Lemma 7 to hold. This is examined in the following remark.
Remark 2 (Minimality of (v)-(viii)).
The four conditions (v)(viii) are minimal for Lemma 7. If any single hypothesis is dropped, we can construct a Bayesian model for which $\mathsf{BCP}$ fails to be continuous (or even well-defined). In particular,
  • Drop (v) (measurability of $k_{\iota}$). Without $\Sigma_{\Theta}$-measurability in $\theta$, the integrals defining $N_{P}$ and $D_{P}$ may not be Lebesgue integrals. Without Borel measurability in $y^{\iota}$, $(y^{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$ and $y^{n}\mapsto D_{P}(y^{n})$ need not even be measurable, so continuity cannot hold.
  • Drop (vi) (pointwise $y$-continuity a.e.). Steps 1-2 in the proof require continuity in the data arguments: For $\lambda_{\Theta}$-almost all $\theta$, $y^{\iota}\mapsto k_{\iota}(\theta,y^{\iota})$ must be continuous. If $k_{n+1}(\theta,\cdot)$ has a jump, then $N_{P}$ can jump in $(y_{n+1},y^{n})$; If $k_{n}(\theta,\cdot)$ has a jump, then $D_{P}$ can jump in $y^{n}$. Either breaks continuity of $p_{P}=N_{P}/D_{P}$.
  • Drop (vii) (integrable envelope). The dominated convergence steps (for $N_{P}$ and $D_{P}$ as functions of the data) and the uniform $L^{1}$ control in the prior argument rely on an integrable envelope $M\in L^{\infty}(\lambda_{\Theta})$ bounding $k_{n+1}$ and $k_{n}$, for almost all $\theta$. Without such $M$, we can have $P_{m}\to P$ in $L^{1}$ but $N_{P_{m}}\not\to N_{P}$ (and similarly for $D_{P}$), so $\mathsf{BCP}$ is discontinuous.
  • Drop (viii) (positive marginal likelihood). If $D_{P}(y^{n})=0$ for some $y^{n}$, then $p_{P}(\cdot\mid y^{n})$ is undefined (division by zero), so $\mathsf{BCP}$ is not even well-defined there, let alone continuous.
In practice, conditions (v)(viii) are almost built in to reasonable Bayesian models. First, in standard dominated parametric families (Gaussian, exponential families, finite mixtures with continuous component densities), the likelihood kernel $f_{\theta}(y)$ is continuous in $y$, and the joint density $k_{n}(\theta,y^{n})=\prod_{i=1}^{n}f_{\theta}(y_{i})$ is Borel in $y^{n}$ and $\Sigma_{\Theta}$-measurable in $\theta$, so (v) holds.
Condition (vi) ensures that, as the data $y^{n}$ vary, the integrands in $N_{P}$ and $D_{P}$ vary continuously (for $\lambda_{\Theta}$-almost all $\theta$), so dominated convergence applies. In practice, standard kernels (Normal, Laplace, Student-$t$, etc.) are smooth in $y$.
The uniform-in-$y$ bound in (vii) is what allows swapping limits and integrals in the proof of Lemma 7. Because $Y$ is compact by (i) and $k_{\iota}(\theta,\cdot)$ is continuous, one can take
$M_{n}(\theta)\coloneqq\sup_{y^{n}\in Y^{n}}k_{n}(\theta,y^{n}),\qquad M_{n+1}(\theta)\coloneqq\sup_{y^{n+1}\in Y^{n+1}}k_{n+1}(\theta,y^{n+1}),$
and use $M(\theta)\geq\max\{M_{n}(\theta),M_{n+1}(\theta)\}$. Under mild conditions on $\Theta$ (so that the kernels are uniformly bounded), $M\in L^{\infty}(\lambda_{\Theta})$, so (vii) holds. [^11] Alternatively, we can ask that $M$ is in $L^{1}$, and topologize $\Delta^{\text{dens}}_{\Theta}$ with the weighted topology $L^{1}(M\text{d}\lambda_{\Theta})$.
Since $N_{P}$ involves $k_{n+1}$, the prior integrability condition $\int M(\theta)P(\mathrm{d}\theta)<\infty$ guarantees finiteness and continuity of the predictive. This is usually automatic with common priors when the envelope $M$ has at most moderate growth in $\theta$ (e.g., exponential-family models with conjugate priors).
Finally, (viii) rules out division by zero in $p_{P}(y_{n+1}\mid y^{n})=N_{P}/D_{P}$. In well-posed models, the prior assigns positive mass to parameter values that could have generated the observed data (i.e., $k_{n}(\theta,y^{n})>0$ on a set of $\theta$ with positive prior mass), ensuring $D_{P}(y^{n})>0$.
In short, (v)(viii) are minimal technical machinery to ensure the posterior predictive is well defined and depends continuously on both the data $y^{n}$ and the prior $P$, and they are typically satisfied in practical modeling. $\triangle$
Before the main result of this section, we need to introduce two further conditions, one of which on the posterior predictive Cumulative Density Function (CDF). Recall that we denote by $\Delta_{Y}^{\text{dens}}$ the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$.
  1. (ix)
    Well-behaved CDF (uniform near the target). Fix $\alpha\in(0,1)$. For each $(y^{n},P)\in Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}$, define
    $F(c;y^{n},P)\coloneqq\int_{\{y:p_{P}(y\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\lambda_{Y}(\mathrm{d}y),\quad c(\alpha,y^{n},P)\coloneqq\inf\{c:F(c;y^{n},P)\geq\alpha\}.$
[…]
    $\big{[}c(\alpha,y^{n},P)-\delta,c(\alpha,y^{n},P)+\delta\big{]}.$
[…]
Call then $\mathsf{QUANT}:[0,1]\times Y^{n}\times\Delta_{\Theta}^{\text{dens}}\rightrightarrows Y$ the set-valued map that extracts the $\alpha$-level set of the posterior predictive distribution. That is,
$(\alpha,y^{n},P)\mapsto\mathsf{QUANT}(\alpha,y^{n},P)\coloneqq H\bigl{(}c(\alpha,y^{n},P),y^{n},P\bigr{)},$
where
$H:\mathbb{R}\times Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}\rightrightarrows Y,\qquad H\bigl{(}c,y^{n},P\bigr{)}\coloneqq\{y\in Y:p_{P}(y\mid y^{n})\geq c\}.$
Lemma 8.
Under assumptions (i)-(ix), the correspondence $\mathsf{QUANT}$ is upper hemicontinuous.
Lemma 8 is proved in Appendix B. Let us add a brief discussion on condition (ix). In many continuous Bayesian models, the posterior predictive density $y\mapsto p_{P}(y\mid y^{n})$ is continuous and strictly positive on compact $Y$. Then, for every $c$ with $\lambda_{Y}\left(\{y:p_{P}(y\mid y^{n})=c\}\right)=0$, the map
$F(c;y^{n},P)=\int_{\{p_{P}(\cdot\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\lambda_{Y}(\mathrm{d}y)$
is continuous (by dominated convergence). Moreover, $F$ is strictly increasing at a level $c_{0}$ whenever $c_{0}$ lies in the interior of the essential range of $p_{P}(\cdot\mid y^{n})$, i.e., there exists $\varepsilon>0$ such that
$\lambda_{Y}\left(\{c_{0}-\varepsilon<p_{P}(y\mid y^{n})\leq c_{0}\}\right)>0\quad\text{and}\quad\lambda_{Y}\left(\{c_{0}<p_{P}(y\mid y^{n})\leq c_{0}+\varepsilon\}\right)>0.$
In particular, if $c_{0}=c(\alpha,y^{n},P)$ satisfies this condition, then $F(\cdot;y^{n},P)$ is continuous and strictly increasing on $c_{0}-\varepsilon,c_{0}+\varepsilon]$, so (ix) holds. This is automatic, for example, if $Y$ is connected and $\lambda_{Y}$ assigns positive measure to every nonempty open set.
Condition (ix) may fail if: (i) $p_{P}$ has a plateau of positive $\lambda_{Y}$-measure at some level $c_{\star}$ (so $F$ is locally flat there); (ii) the predictive is not absolutely continuous with respect to $\lambda_{Y}$ (atoms create jumps of $F$); or (iii) $p_{P}$ never attains values in a neighborhood of $c(\alpha,y^{n},P)$ (a “gap” in its range, e.g., on disconnected $Y$), which precludes local strict increase.
The last condition that we need is the following.
  1. (x)
    Leave-One-Out (LOO) stability and conditional Glivenko-Cantelli (GC). Recall that $\mathfrak{P}$ denotes the sampling law of the exchangeable process $(\mathcal{Y}_{i})_{i\geq 1}$. For $(y^{n},P)$ and $y\in Y$, define
    $r_{n}(y)\coloneqq p_{P}(y\mid y^{n}),\qquad r_{n,-i}(y)\coloneqq p_{P}(y\mid y^{n+1}_{-i}),\quad 1\leq i\leq n.$
    1. (a)
      LOO stability.
      $\mathsf{err}_{n}(y^{n},P)\coloneqq\sup_{y\in Y}\max_{1\leq i\leq n}\bigl{|}r_{n,-i}(y)-r_{n}(y)\bigr{|}=o_{\mathfrak{P}}(1).$
    2. (b)
      Conditional GC. Let
[…]
      We have
      $\sup_{t\in\mathbb{R}}\bigl{|}F_{n}(t;y^{n},P)-F(t;y^{n},P)\bigr{|}=o_{\mathfrak{P}}(1).$
A natural question at this point is why we need to introduce condition (x). Full CP compares $r_{n,-i}(y_{i})$ with $r_{n}(y)$; The left-hand side depends on the candidate $y$ through $y^{n+1}_{-i}$. To recover the population $\alpha$-level set threshold defined by $F(\cdot;y^{n},P)$, we need two probabilistic approximations, $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\{r_{n,-i}(y_{i})\leq r_{n}(y)\}]\approx\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\{r_{n}(y_{i})\leq r_{n}(y)\}]\approx F\bigl{(}r_{n}(y);y^{n},P\bigr{)},$ uniformly in $y$. The first approximation is the LOO stability (x).(a); The second is the conditional GC property (x).(b). Neither follows from (i)(ix) (which are analytic/topological).
Conditions like (x) are mild and hold in common settings: (i) In regular parametric Bayes (e.g., smooth exponential families on compact $Y$ with conjugate priors), $r_{n,-i}$ and $r_{n}$ differ by $O_{\mathfrak{P}}(1/n)$ uniformly in $y$ (posterior contraction / Lipschitz dependence on empirical statistics), yielding (x).(a); (ii) In regular dominated models on compact $Y$ (parametric or suitably controlled nonparametric), the posterior predictives $r_{n}$ converge uniformly to a continuous limit $r$, and the posterior predictive laws $Q_{n}(\cdot)\coloneqq\int_{(\cdot)}p_{P}(y\mid y^{n})\lambda_{Y}(\text{d}y)$ converge in probability (as random probability measures) to a probability measure $Q$ with density $r$, i.e. $\|r_{n}-r\|_{\infty}\xrightarrow{\mathfrak{P}}0$, $Q_{n}\to Q\text{ in probability}$, $\frac{\text{d}Q}{\text{d}\lambda_{Y}}=r$. If, in addition, there are no plateaus at the target level (equivalently, $Q(\{r=t\})=0$ for the relevant thresholds $t$), then (x).(b) holds by a uniform Law of Large Numbers applied to the sublevel sets $\{r\leq t\}$. Strengthening both to almost sure versions upgrades all conclusions from “in probability” to “almost surely”.
Lemmas 7 and 8 are crucial in light of the following, which is the main result of this section.
Theorem 9 (Unifying Bayes, Conformal, and Imprecise Prediction in $\mathbf{UHCont}$).
Assume that (i)-(x) and consonance hold. Fix any $\alpha\in(0,1)$ and define $\beta_{n}(\alpha)\coloneqq\min\{k/(n+1):k/(n+1)>\alpha\}$. Then,
$d_{H}\left(\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)}),\mathsf{QUANT}_{\alpha}(y^{n},P)\right)\xrightarrow{\mathfrak{P}}0.$
Consequently, since $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$ under consonance,
$d_{H}\left(\mathsf{IHDR}_{\alpha}\bigl{(}\mathsf{CRED}(y^{n},\psi_{(y^{n},P)})\bigr{)},\mathsf{QUANT}_{\alpha}(y^{n},P)\right)\xrightarrow{\mathfrak{P}}0.$
Equivalently, the diagram
$Y^{n}\times\mathscr{F}$$\mathscr{C}$$Y^{n}\times\Delta^{\text{dens}}_{\Theta}$$\Sigma_{Y}$$\scriptstyle{\mathsf{CRED}}$$\scriptstyle{\kappa_{\alpha}}$$\scriptstyle{\mathsf{IHDR}_{\alpha}}$$\scriptstyle{\mathsf{BCP}}$$\scriptstyle{\mathsf{QUANT}_{\alpha}}$
commutes asymptotically in probability in $\mathbf{UHCont}$ (in the Hausdorff distance $d_{H}$).
If $Y$ is compact metric, [^12] We also tacitly assume that the level set $\{y:r_{n}(y)=c(\alpha,y^{n},P)\}$ has empty interior (equivalently, that $\{r_{n}>c(\alpha,y^{n},P)\}$ is dense in $\{r_{n}\geq c(\alpha,y^{n},P)\}$). the set convergence above is in the Hausdorff distance $d_{H}$; Otherwise a similar result can be proved with convergence taken in the Vietoris topology on closed subsets of $Y$.
Proof of Theorem 9.
First, we point out that Fong and Holmes, (2021) show that the negative posterior predictive density is a valid non-conformity score (i.e. it is an element of $\mathscr{F}$ in (4)). Fix then $\alpha\in(0,1)$ and write $r_{n}(y)\coloneqq p_{P}(y\mid y^{n})$. For $y\in Y$ define the full CP transducer
$\pi_{n}(y)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}[\{r_{n,-i}(y_{i})\leq r_{n}(y)\}],$
with the convention $y_{n+1}=y$ and $r_{n,-(n+1)}(y_{n+1})\equiv r_{n}(y)$. Then $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\}$.
Step 1 (LOO reduction via localized bracketing). Set
[…]
Fix $\eta>0$ and define the $\alpha$-band
$\mathcal{B}_{\eta}(y^{n},P)\coloneqq\bigl{\{}y\in Y:|F(r_{n}(y);y^{n},P)-\alpha|\leq\eta\bigr{\}}.$
For every $y\in Y$,
$F_{n}\bigl{(}r_{n}(y)-\delta_{n};y^{n},P\bigr{)}\leq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\bigl{[}r_{n,-i}(y_{i})\leq r_{n}(y)\bigr{]}\leq F_{n}\bigl{(}r_{n}(y)+\delta_{n};y^{n},P\bigr{)}.$
Hence, for $y\in\mathcal{B}_{\eta}(y^{n},P)$,
$\Bigl{|}\tfrac{1}{n}$ $\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}-F_{n}\bigl{(}r_{n}(y);y^{n},P\bigr{)}\Bigr{|}$
$\leq\sup_{\begin{subarray}{c}|t-r_{n}(y)|\leq\delta_{n}\end{subarray}}\bigl{|}F_{n}(t;y^{n},P)-F_{n}(r_{n}(y);y^{n},P)\bigr{|}$
$\leq\sup_{\begin{subarray}{c}|t-r_{n}(y)|\leq\delta_{n}\end{subarray}}\bigl{|}F(t;y^{n},P)-F(r_{n}(y);y^{n},P)\bigr{|}+2\sup_{u\in\mathbb{R}}\bigl{|}F_{n}(u;y^{n},P)-F(u;y^{n},P)\bigr{|}.$
By (ix), $c\mapsto F(c;y^{n},P)$ is continuous in a neighborhood of $c(\alpha,y^{n},P)$; Since $y\mapsto r_{n}(y)$ is continuous (Lemma 7) and $Y$ is compact, there is a deterministic modulus $\omega(\cdot)$ with
$\sup_{y\in\mathcal{B}_{\eta}(y^{n},P)}\sup_{|t-r_{n}(y)|\leq\delta}\bigl{|}F(t;y^{n},P)-F(r_{n}(y);y^{n},P)\bigr{|}\leq\omega(\delta),\qquad\omega(\delta)\downarrow 0\text{ as }\delta\downarrow 0.$
Together with (x).(b) we obtain the localized reduction
$\sup_{y\in\mathcal{B}_{\eta}(y^{n},P)}\Bigl{|}\tfrac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}-F\bigl{(}r_{n}(y);y^{n},P\bigr{)}\Bigr{|}=o_{\mathfrak{P}}(1).$
Finally, recall that
$\pi_{n}(y)=\frac{1}{n+1}\sum_{i=1}^{n+1}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}=\frac{n}{n+1}\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}+\frac{1}{n+1},$
so, since $n/(n+1)\to 1$,
$\sup_{y\in\mathcal{B}_{\eta}(y^{n},P)}\bigl{|}\pi_{n}(y)-F\bigl{(}r_{n}(y);y^{n},P\bigr{)}\bigr{|}=o_{\mathfrak{P}}(1).$
Step 2 (Decisions agree up to a shrinking $r$-tube). Let $c^{\star}\coloneqq c(\alpha,y^{n},P)$. By (ix), $c\mapsto F(c;y^{n},P)$ is strictly increasing near $c^{\star}$, so for any small $\delta>0$ define
$\gamma_{+}(\delta)\coloneqq F(c^{\star}+\delta;y^{n},P)-\alpha>0,\qquad\gamma_{-}(\delta)\coloneqq\alpha-F(c^{\star}-\delta;y^{n},P)>0,$
with $\gamma_{\pm}(\delta)\downarrow 0$ as $\delta\downarrow 0$. Recall that $\delta_{n}\coloneqq\sup_{y\in Y}\max_{i\leq n}|r_{n,-i}(y)-r_{n}(y)|\xrightarrow{\mathfrak{P}}0$. By LOO stability, for every $y\in Y$,
$F_{n}\bigl{(}r_{n}(y)-\delta_{n};y^{n},P\bigr{)}\leq G_{n}(y)\leq F_{n}\bigl{(}r_{n}(y)+\delta_{n};y^{n},P\bigr{)},$
where $G_{n}(y)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{r_{n,-i}(y_{i})\leq r_{n}(y)\}$. Fix $\delta>0$ and write $c^{\star}=c(\alpha,y^{n},P)$. By (ix), $c\mapsto F(c;y^{n},P)$ is strictly increasing and continuous near $c^{\star}$. Thus, there is a modulus $\omega(\cdot)\downarrow 0$ such that for small $h$,
$\bigl{|}F(c^{\star}\pm\delta\pm h;y^{n},P)-F(c^{\star}\pm\delta;y^{n},P)\bigr{|}\leq\omega(|h|).$
Let $\mathscr{D}_{n}\coloneqq\sup_{t\in\mathbb{R}}\bigl{|}F_{n}(t;y^{n},P)-F(t;y^{n},P)\bigr{|}\xrightarrow{\mathfrak{P}}0$ by (x).(b). Then:
(Left inclusion). If $r_{n}(y)\geq c^{\star}+\delta$, then
$G_{n}(y)$ $\geq F_{n}\bigl{(}r_{n}(y)-\delta_{n}\bigr{)}\geq F\bigl{(}r_{n}(y)-\delta_{n}\bigr{)}-\mathscr{D}_{n}$
$\geq F\bigl{(}c^{\star}+\delta-\delta_{n}\bigr{)}-\mathscr{D}_{n}\geq\alpha+\gamma_{+}(\delta)-\omega(\delta_{n})-\mathscr{D}_{n},$
where $\gamma_{+}(\delta)\coloneqq F(c^{\star}+\delta;y^{n},P)-\alpha>0$. Hence,
$\pi_{n}(y)=\tfrac{n}{n+1}G_{n}(y)+\tfrac{1}{n+1}\geq\alpha+\gamma_{+}(\delta)-\omega(\delta_{n})-\mathscr{D}_{n}+\tfrac{1}{n+1}\geq\beta_{n}(\alpha),$
for all large $n$.
(Right inclusion). If $r_{n}(y)\leq c^{\star}-\delta$, then
$G_{n}(y)$ $\leq F_{n}\bigl{(}r_{n}(y)+\delta_{n}\bigr{)}\leq F\bigl{(}r_{n}(y)+\delta_{n}\bigr{)}+\mathscr{D}_{n}$
$\leq F\bigl{(}c^{\star}-\delta+\delta_{n}\bigr{)}+\mathscr{D}_{n}\leq\alpha-\gamma_{-}(\delta)+\omega(\delta_{n})+\mathscr{D}_{n},$
with $\gamma_{-}(\delta)\coloneqq\alpha-F(c^{\star}-\delta;y^{n},P)>0$, thus $\pi_{n}(y)<\beta_{n}(\alpha)$ for all large $n$.
Consequently, for any fixed $\delta>0$, with probability $\to 1$,
$\{y:r_{n}(y)\geq c^{\star}+\delta\}\subseteq\ \{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\}\subseteq\ \{y:r_{n}(y)\geq c^{\star}-\delta\}.$ (11)
Thus any disagreement between $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})=\{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\}$ and the target set $\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}=\{y:r_{n}(y)\geq c^{\star}\}$ is confined to the “shrinking tube”
$\mathcal{T}_{\delta}(y^{n},P)\coloneqq\{y\in Y:|r_{n}(y)-c^{\star}|\leq\delta\}.$
Since $r_{n}$ is continuous on compact $Y$ (Lemma 7), the level sets $\{r_{n}\geq c^{\star}\pm\delta\}$ converge to $\{r_{n}\geq c^{\star}\}$ in Hausdorff distance as $\delta\downarrow 0$. Combined with (11), this yields the desired set convergence.
Step 3 (from inclusions to the target set). By (ix) and Lemma 7, the map $y\mapsto F(r_{n}(y);y^{n},P)$ is continuous on $Y$ at the decision level $\alpha$, and (see Lemma 8) the superlevel sets $\{y:F(r_{n}(y);y^{n},P)\geq t\}$ vary continuously in $t$ near $t=\alpha$. Therefore, as $\delta\downarrow 0$, the Hausdorff distance between $\{y:F(r_{n}(y);y^{n},P)\geq\alpha+\delta\}$ and $\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}$ (and similarly for $\alpha-\delta$) tends to 0. Combining with the inclusions above yields
$d_{H}\left(\{y:\pi_{n}(y)\geq\beta_{n}(\alpha)\},\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}\right)\xrightarrow{\mathfrak{P}}0,$
i.e. $d_{H}\bigl{(}\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)}),\mathsf{QUANT}_{\alpha}(y^{n},P)\bigr{)}\xrightarrow{\mathfrak{P}}0.$
Step 4 (Hausdorff convergence in probability). By Lemma 7 and (ix), the target $\alpha$-level set $\mathcal{Q}_{\alpha}(y^{n},P)\coloneqq\{y:F(r_{n}(y);y^{n},P)\geq\alpha\}=\mathsf{QUANT}_{\alpha}(y^{n},P)$ is a closed subset of the compact space $Y$. The sandwich above implies that, on events with probabilities $\to 1$, the inner and outer deviations of $\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)})$ from $\mathcal{Q}_{\alpha}(y^{n},P)$ shrink to 0. Therefore,
$d_{H}\left(\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)}),\mathcal{Q}_{\alpha}(y^{n},P)\right)\xrightarrow{\mathfrak{P}}0.$
Step 5 (Top route). Under consonance, Theorem 5 gives $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$ for any score. Composing with $\mathsf{BCP}$ yields the second convergence in probability. ∎
Theorem 9 is the central payoff of our categorical treatment, and perhaps the greatest joy (so far!) of looking at Conformal Prediction from a categorical angle. It confirms the intuition put forth in Martin, (2022) and Caprio et al., 2025a (, Section 5.2), that CP is a bridge between the Bayesian, frequentist, and imprecise approaches to predictive statistical reasoning. When seen as u.h.c. correspondences, the model-based Bayesian and imprecise methods, and the model-free conformal construction, yield (asymptotically) the same $\alpha$-level prediction region, thus showing a very profound connections between these three only apparently far apart prediction mechanisms.
4.1. Conformal Prediction and e-Posteriors
Theorem 9 allows us to tie CP to the concept of e-posterior (Grünwald,, 2023) as well, thus further testifying the central role of CP in modern Machine Learning.
To see this, recall that $\Delta_{Y}^{\text{dens}}$ denotes the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$. Since we observe the realizations $y^{n}=(y_{1},\ldots,y_{n})$ of the first $n$ terms $\mathcal{Y}_{1},\ldots,\mathcal{Y}_{n}$ of an exchangeable process (as pointed out in Section 2.1), the likelihood density $p(y^{n}\mid\theta)\equiv\ell(y^{n}\mid\theta)$ can be written as (de Finetti,, 1974, 1975)
$\ell(y^{n}\mid\theta)=\int\left[\prod_{i=1}^{n}h(y_{i}\mid\varsigma)\right]G_{\theta}(\text{d}\varsigma),$
where $h(\cdot\mid\varsigma)$ is the density of an element $L_{\varsigma}\in\Delta_{Y}^{\text{dens}}$ (given a latent parameter $\varsigma$), and $G_{\theta}$ is a mixing distribution indexed by the parameter $\theta\in\Theta$.
Consider then a credal prior set $\mathcal{P}_{\text{prior}}$, that is, a nonempty convex and weak-closed (and hence also closed in the $L^{1}(\lambda_{\Theta})$ topology) subset of $\Delta_{\Theta}^{\text{dens}}$. Let $\overline{P}$ be its upper probability, and $\overline{p}\coloneqq\sup_{P\in\mathcal{P}_{\text{prior}}}\frac{\text{d}P}{\text{d}\lambda_{\Theta}}$ be the upper prior density; The lower prior density $\underline{p}$ is defined similarly, with $\sup$ replaced by $\inf$.
Proposition 10 (The Upper Posterior is an e-Posterior).
Suppose that $\underline{p}\in L^{1}(\lambda_{\Theta})$, that $\underline{\ell}(y^{n})\coloneqq\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)>0$, for all $y^{n}\in Y^{n}$, and that the upper prior density $\overline{p}(\theta)\in(0,\infty)$, for all $\theta\in\Theta$. Then the upper posterior
$\overline{p}(\theta\mid y^{n})=\frac{\ell(y^{n}\mid\theta)\overline{p}(\theta)}{\underline{\ell}(y^{n})}$ (12)
is an e-posterior in the sense of Grünwald, (2023) if and only if
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),\quad\text{for all }\theta\in\Theta.$ (13)
We can interpret the necessary and sufficient condition in Proposition 10 as follows: The probability mass on which all priors agree is no larger than the minimum “peak plausibility” that any state enjoys under at least one prior. Since Proposition 10 is of general significance (the IP and e-values communities may be interested in it), we leave its proof in the main portion of our paper.
Proof of Proposition 10.
First, notice that the upper posterior density in (12) is well defined. In particular, it is an upper bound for the generalized Bayes’ upper posterior density (Walley,, 1991, Theorem 6.4.6).
Then, recall that an e-posterior is a function $p^{\text{e}}(\theta\mid y^{n})=1/Z_{\theta}(y^{n})$, where $Z_{\theta}(y^{n})$ is an e-variable, that is, a statistic for which $\mathbb{E}_{Y^{n}\sim L_{\theta}}[Z_{\theta}(Y^{n})]\leq 1$. Call $\lambda_{Y}^{\otimes n}$ the $n$-fold product of $\lambda_{Y}$; We have
$\mathbb{E}_{Y^{n}\sim L_{\theta}}\left[\frac{1}{\overline{p}(\theta\mid Y^{n})}\right]$ $=\mathbb{E}_{Y^{n}\sim L_{\theta}}\left[\frac{\underline{\ell}(Y^{n})}{\ell(Y^{n}\mid\theta)\overline{p}(\theta)}\right]$
$=\int_{Y^{n}}\frac{\underline{\ell}(y^{n})}{\ell(y^{n}\mid\theta)\overline{p}(\theta)}\ell(y^{n}\mid\theta)\lambda_{Y}^{\otimes n}(\text{d}y^{n})$
$=\frac{1}{\overline{p}(\theta)}\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})$
$\leq 1\iff\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})\leq\overline{p}(\theta),$
for all $\theta\in\Theta$. Now, notice that
$\int_{Y^{n}}\underline{\ell}(y^{n})\lambda_{Y}^{\otimes n}(\text{d}y^{n})$ $=\int_{Y^{n}}\int_{\Theta}\ell(y^{n}\mid\theta)\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\lambda_{Y}^{\otimes n}(\text{d}y^{n})$
$=\int_{\Theta}\underline{p}(\theta)\int_{Y^{n}}\ell(y^{n}\mid\theta)\lambda_{Y}^{\otimes n}(\text{d}y^{n})\lambda_{\Theta}(\mathrm{d}\theta)$
$=\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta),$
where the last-but-one equality comes from Fubini-Tonelli, and the last one comes from $\ell(y^{n}\mid\theta)$ being a proper density. Hence, the upper posterior density $\overline{p}(\theta\mid y^{n})$ is an e-posterior if and only if
$\int_{\Theta}\underline{p}(\theta)\lambda_{\Theta}(\mathrm{d}\theta)\leq\overline{p}(\theta),$ (14)
for all $\theta\in\Theta$. This concludes the proof. ∎
A consequence of Proposition 10 is that, if (14) holds and
$\overline{\Pi}(A)\equiv\overline{\Pi}_{(y^{n},\psi)}(A)=\int_{A}\underbrace{\int_{\Theta}\ell(y_{n+1}\mid\theta)\overline{p}(\theta\mid y^{n})\lambda_{\Theta}(\text{d}\theta)}_{\eqqcolon\overline{p}(y_{n+1}\mid y^{n})}\lambda_{Y}(\text{d}y_{n+1}),\quad\forall A\in\Sigma_{Y},$ (15)
then the $\alpha$-level Imprecise Highest Density Region $\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi}))$ in Theorem 9 is a prediction region that is associated with an e-posterior. The equality in (15) only holds if the upper posterior predictive density $\overline{p}(y_{n+1}\mid y^{n})$ collapses to a point mass. We will denote
$\overline{\Pi}_{\mathrm{B}}(A)\coloneqq\int_{A}\overline{p}(y\mid y^{n})\lambda_{Y}(dy),\qquad A\in\Sigma_{Y},$
so that (15) reads $\overline{\Pi}=\overline{\Pi}_{\mathrm{B}}$. The following result is proved in Appendix B.
Proposition 11 (Coincidence of Upper Probabilities implies Degeneracy).
Assume (15), and relax the assumption that $\lambda_{Y}$ is atomless. Then, there exists $y^{\ast}\in Y$ such that
$\overline{\Pi}_{\mathrm{B}}(\cdot)=\delta_{y^{\ast}}(\cdot)\quad\text{and}\quad\overline{P}\left(\{\theta:L_{\theta}=\delta_{y^{\ast}}\}\middle|y^{n}\right)=1.$
In particular, if $\overline{P}(\cdot\mid y^{n})\ll\lambda_{\Theta}$, then necessarily $\lambda_{Y}(\{y^{\ast}\})>0$ and there exists a density $\overline{p}(\cdot\mid y^{n})\in L^{1}(\lambda_{\Theta})$ supported on $\Theta^{\ast}\coloneqq\{\theta:L_{\theta}=\delta_{y^{\ast}}\}$, i.e.
$\overline{p}(\theta\mid y^{n})=0\quad\text{for }\lambda_{\Theta}\text{-almost all }\theta\notin\Theta^{\ast},\qquad\int_{\Theta^{\ast}}\overline{p}(\theta\mid y^{n})\lambda_{\Theta}(d\theta)=1.$
On $Y$ we have
$\overline{p}(y\mid y^{n})=\frac{\mathbbm{1}(y=y^{\ast})}{\lambda_{Y}(\{y^{\ast}\})}.$
If, in addition, the model is identifiable (the map $\theta\mapsto L_{\theta}$ is injective), then $\overline{P}(\cdot\mid y^{n})=\delta_{\theta^{\ast}}$.
If $\lambda_{Y}$ is atomless and $\overline{p}_{P}(\cdot\mid y^{n})\in L^{1}(\lambda_{Y})$, then (15) cannot hold for all $A\in\Sigma_{Y}$. Indeed, (15) would force $\overline{\Pi}_{\mathrm{B}}=\delta_{y^{\ast}}$, which is not absolutely continuous with respect to an atomless $\lambda_{Y}$. Although Proposition 11 shows that the equality in (15) holds only under very restrictive conditions, it nevertheless hints at the possibility of an arrow in the diagram of Theorem 9 connecting $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ and $\mathscr{C}$ via an e-posterior-based method. [^13] To be precise, for such an arrow to be a morphism of $\mathbf{UHCont}$, we need to substitute $Y^{n}\times\Delta_{\Theta}^{\text{dens}}$ with $Y^{n}\times\mathscr{C}_{\Theta}^{\text{dens}}$, where $\mathscr{C}_{\Theta}^{\text{dens}}$ denotes the set of convex and weak-closed subsets of $\Delta_{\Theta}^{\text{dens}}$. In that case, $\mathsf{BCP}$ and $\mathsf{QUANT}_{\alpha}$ both take as input the pair $(y^{n},\{P\})$, where $\{P\}\subset\mathcal{P}_{\text{prior}}$. A deeper study of the broader interest of the e-posterior approach to Conformal Prediction within a category theory framework is left for future research.
5. Conformal Prediction Regions as Images of the Imprecise Highest Density Region Functor
In this section, we see how the correspondence $\mathsf{IHDR}_{\alpha}$ can be cast as a functor between two well-known categories. As a consequence, the Conformal Prediction Region can be seen as a functor image.
Theorem 12 (Functoriality of the $\alpha$-IHDR).
Let $(Y,\Sigma_{Y}=2^{Y})$ be a measurable space, and let $\Delta_{Y}$ be the set of all finitely additive probability measures on $\Sigma_{Y}$ endowed with the weak topology. Let
$\mathscr{C}\coloneqq\bigl{\{}\mathcal{M}\subseteq\Delta_{Y}:\mathcal{M}\text{ is convex, nonempty, and weak${}^{\star}$-closed}\bigr{\}},$
and regard both $\mathscr{C}$ and $\Sigma_{Y}$ as (pre-)categories whose only morphisms are inclusions $\subseteq$”. Fix $\alpha\in[0,1]$. For $\mathcal{M}\in\mathscr{C}$, define
$B_{\mathcal{M}}^{\alpha}\coloneqq\bigcap\Bigl{\{}A\in\Sigma_{Y}:\underline{P}_{\mathcal{M}}(A)\geq 1-\alpha\Bigr{\}}.$
Set
$\mathsf{IHDR}_{\alpha}:\mathscr{C}\rightarrow\Sigma_{Y},\quad\mathcal{M}\mapsto B_{\mathcal{M}}^{\alpha},\qquad\mathsf{IHDR}_{\alpha}(\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2})\coloneqq B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha}.$
Then, $\mathsf{IHDR}_{\alpha}$ is a covariant functor.
Proof.
(i) Objects. Every $A$ in the defining intersection for $B_{\mathcal{M}}^{\alpha}$ is measurable; Arbitrary intersections of sets in $\Sigma_{Y}$ remain in $\Sigma_{Y}$, hence $B_{\mathcal{M}}^{\alpha}\in\Sigma_{Y}$. Thus $\mathsf{IHDR}_{\alpha}$ maps objects of $\mathscr{C}$ to objects of $\Sigma_{Y}$.
(ii) Morphisms (monotonicity). Suppose $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$. Let $y\in B_{\mathcal{M}_{1}}^{\alpha}$. If $y\notin B_{\mathcal{M}_{2}}^{\alpha}$, then by definition of $B_{\mathcal{M}_{2}}^{\alpha}$, there exists $A\in\Sigma_{Y}$ such that
$\underline{P}_{\mathcal{M}_{2}}(A)\geq 1-\alpha,\qquad y\notin A.$
Because $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$, the same inequality holds with $\mathcal{M}_{2}$ replaced by $\mathcal{M}_{1}$, so $A$ belongs to the family of sets whose intersection defines $B_{\mathcal{M}_{1}}^{\alpha}$. Consequently $y\in A$, contradicting $y\notin A$. Therefore $y\in B_{\mathcal{M}_{2}}^{\alpha}$, and $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}.$
Since $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}$ implies $\overline{P}_{\mathcal{M}_{1}}(A)\leq\overline{P}_{\mathcal{M}_{2}}(A)$ for all $A\in\Sigma_{Y}$, we have $\underline{P}_{\mathcal{M}_{1}}(A)=1-\overline{P}_{\mathcal{M}_{1}}(A^{c})\geq 1-\overline{P}_{\mathcal{M}_{2}}(A^{c})=\underline{P}_{\mathcal{M}_{2}}(A)$. Hence
$\{A\in\Sigma_{Y}:\underline{P}_{\mathcal{M}_{2}}(A)\geq 1-\alpha\}\subseteq\{A\in\Sigma_{Y}:\underline{P}_{\mathcal{M}_{1}}(A)\geq 1-\alpha\},$
and therefore $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}$.
Intuitively, enlarging a credal set can only decrease the lower probability (by the monotonicity of lower probabilities), so fewer events clear the $1-\alpha$ bar; Intersecting fewer sets yields a larger IHDR. Hence $\mathsf{IHDR}_{\alpha}$ maps the arrow $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ to the arrow $B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha}$.
(iii) Identities. For each $\mathcal{M}\in\mathscr{C}$, the identity morphism is $\mathcal{M}\hookrightarrow\mathcal{M}$. Since inclusion is reflexive, $B_{\mathcal{M}}^{\alpha}\hookrightarrow B_{\mathcal{M}}^{\alpha}$ is the identity in $\Sigma_{Y}$, so $\mathsf{IHDR}_{\alpha}(\mathrm{id}_{\mathcal{M}})=\mathrm{id}_{B_{\mathcal{M}}^{\alpha}}$.
(iv) Composition. Given $\mathcal{M}_{1}\subseteq\mathcal{M}_{2}\subseteq\mathcal{M}_{3}$, the composite morphism in $\mathscr{C}$ is $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{3}$. By monotonicity, $B_{\mathcal{M}_{1}}^{\alpha}\subseteq B_{\mathcal{M}_{2}}^{\alpha}\subseteq B_{\mathcal{M}_{3}}^{\alpha},$ so
$\mathsf{IHDR}_{\alpha}(\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{3})=B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{3}}^{\alpha}$ $=(B_{\mathcal{M}_{2}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{3}}^{\alpha})\circ(B_{\mathcal{M}_{1}}^{\alpha}\hookrightarrow B_{\mathcal{M}_{2}}^{\alpha})$
$=\mathsf{IHDR}_{\alpha}(\mathcal{M}_{2}\hookrightarrow\mathcal{M}_{3})\circ\mathsf{IHDR}_{\alpha}(\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}).$
Since $\mathsf{IHDR}_{\alpha}$ respects objects, morphisms, identities and composition, it is a covariant functor. ∎
Remark 3 (Contravariant Variants).
Replacing $B_{\mathcal{M}}^{\alpha}$ by its complement $Y\setminus B_{\mathcal{M}}^{\alpha}$ yields a contravariant functor. Allowing $\alpha$ itself to vary introduces a functor $\mathsf{IHDR}\colon[0,1]\times\mathscr{C}\rightarrow\Sigma_{Y}$ that is antitone in the first argument and covariant in the second. $\triangle$
To the best of our knowledge, Theorem 12 is the first time that an imprecise probabilistic notion such as the IHDR is framed as a functor between two categories, thus adding to the field of category-theoretic imprecise probability theory (Liell-Cock and Staton,, 2024).
We also note in passing that we will not delve into the examination of categories $(\mathscr{C},\subseteq)$ and $(\Sigma_{Y},\subseteq)$ because they are two special instances of the well-studied category $\mathbf{Sub}(X)$ of subsets of a generic set $X$, endowed with inclusion morphisms. Here we abuse notation: We write $(\mathscr{C},\subseteq)$ to denote the category whose objects are the elements of $\mathscr{C}$, and whose morphisms are the inclusions, and similarly for $(\Sigma_{Y},\subseteq)$.
We have the following corollary, which allows us to see the Conformal Prediction Region outputted by a Full Conformal Prediction procedure as an object image of functor $\mathsf{IHDR}_{\alpha}$.
Corollary 12.1 (Conformal Prediction as a Functor Image).
Suppose that the conformal transducer $\pi(\cdot,y^{n})$ is consonant, i.e.
$\sup_{y_{n+1}\in Y}\pi(y_{n+1},y^{n})=1.$
Then, for all $\alpha\in[0,1]$, all $y^{n}\in Y^{n}$, and all $\psi\in\mathscr{F}$, we have that
$\mathscr{R}_{\alpha}^{\psi}(y^{n})=\mathsf{IHDR}_{\alpha}(\mathcal{M}(\overline{\Pi})).$
That is, the $\alpha$-level Conformal Prediction Region $\mathscr{R}_{\alpha}^{\psi}(y^{n})$ is the image under functor $\mathsf{IHDR}_{\alpha}$ of a well-defined credal set
$\mathcal{M}(\overline{\Pi})\coloneqq\{P\in\Delta_{Y}:P(A)\leq\overline{\Pi}(A)\coloneqq\sup_{y_{n+1}\in A}\pi(y_{n+1},y^{n})\text{, }\forall A\in\Sigma_{Y}\},$
which is an object of category $(\mathscr{C},\subseteq)$.
Proof.
Immediate from Theorem 12 and Caprio et al., 2025a (, Proposition 5). ∎
Viewing the conformal prediction map $\mathsf{IHDR}_{\alpha}$ as a covariant functor immediately yields practical AI privacy benefits. Suppose that we have many sources producing credal sets $\mathcal{M}_{j}$. First, differentially-private or otherwise noise-adding transformation applied to each site’s credal set $\mathcal{M}_{j}$ propagates through the functor without violating the coverage guarantee, so statistical validity is preserved “for free”.
Second, collaboration can proceed in a strictly federated manner: Every agent needs to transmit only its (possibly privatized) summary object $\mathcal{M}_{j}$—never the underlying data—thereby minimizing exposure while still enabling the consortium to construct a globally valid prediction region.
Caveat. The functorial guarantee is monotone in the credal set: To preserve (and possibly enlarge) coverage after privatization, each site should output a superset $\widetilde{\mathcal{M}}_{j}\supseteq\mathcal{M}_{j}$ (e.g., add a weak buffer and take closed convex hull). Inner approximations $\widetilde{\mathcal{M}}_{j}\subset\mathcal{M}_{j}$ can shrink $B^{\alpha}_{\mathcal{M}}$ and reduce coverage. For federation, aggregation should be performed via the closed convex hull $\overline{\operatorname{co}}(\cdot)$,
$\mathcal{M}_{\mathrm{agg}}=\overline{\operatorname{co}}\Bigl{(}\bigcup_{j}\widetilde{\mathcal{M}}_{j}\Bigr{)},$
which ensures $B^{\alpha}_{\widetilde{\mathcal{M}}_{j}}\subseteq B^{\alpha}_{\mathcal{M}_{\mathrm{agg}}}$, and thus global validity whenever the true law lies in some $\widetilde{\mathcal{M}}_{j}$.
6. Further Properties of our Categories
In this section, we ask ourselves whether we can further characterize the categories that we introduced earlier in Sections 3. Throughout the section, if restrictions have to be imposed on the objects or morphisms of a category, we will report them as superscript and subscripts of the ensuing subcategory, respectively.
6.1. Monoidal Categories
The first, natural question is whether they are monoidal.
Alas, $\mathbf{UHCont}$ is not, as upper hemicontinuity is not preserved under product in general unless the correspondences are compact-valued. We can define a slight strengthening of $\mathbf{UHCont}$, though, which is monoidal.
[…]
Define a structure $\mathbf{UHCont}_{\text{c}}$ as follows,
  • Objects: Topological spaces $(X,\tau_{X})$.
  • Morphisms: Compact-valued, upper hemicontinuous correspondences $\Phi:X\rightrightarrows Y$.
Theorem 14 ($\mathbf{UHCont}_{\text{c}}$ is a Monoidal Category).
The structure $\mathbf{UHCont}_{\text{c}}$ is a monoidal category under the Cartesian product of topological spaces.
Proof.
We verify the monoidal category structure.
(1) Tensor Product of Objects and Morphisms. For objects $X,Y$, define $X\otimes Y\coloneqq X\times Y$, the Cartesian product with the product topology.
Given morphisms $\Phi_{1}:X\rightrightarrows X^{\prime}$ and $\Phi_{2}:Y\rightrightarrows Y^{\prime}$, define their tensor product as
$(\Phi_{1}\otimes\Phi_{2})(x,y)\coloneqq\Phi_{1}(x)\times\Phi_{2}(y).$
Since compact sets are preserved under finite products, $\Phi_{1}(x)\times\Phi_{2}(y)$ is compact for all $(x,y)\in X\times Y$. Also, the product of upper hemicontinuous correspondences with compact values is upper hemicontinuous with compact values (Aliprantis and Border,, 2006, Theorem 17.28). Thus, $\Phi_{1}\otimes\Phi_{2}$ is a morphism in $\mathbf{UHCont}_{\text{c}}$.
(2) Unit Object. Let the unit object be the singleton space $I\coloneqq\{*\}$, with the trivial topology. For any object $X$, the canonical identifications
$\lambda_{X}:I\times X\rightarrow X,\quad\rho_{X}:X\times I\rightarrow X$
are homeomorphisms. They define morphisms in $\mathbf{UHCont}_{\text{c}}$ via single-valued, continuous correspondences (e.g., $(*,x)\mapsto x$) with compact singleton values.
(3) Associativity. Define the associator as the canonical homeomorphism,
$\alpha_{X,Y,Z}:(X\times Y)\times Z\rightarrow X\times(Y\times Z).$
Its graph defines a single-valued, continuous correspondence with compact-valued singleton fibers. Hence, it is a morphism in the category.
(4) Bifunctoriality of $\otimes$. For morphisms $\Phi_{1}:X\rightrightarrows X^{\prime}$, $\Psi_{1}:X^{\prime}\rightrightarrows X^{\prime\prime}$ and $\Phi_{2}:Y\rightrightarrows Y^{\prime}$, $\Psi_{2}:Y^{\prime}\rightrightarrows Y^{\prime\prime}$, for each $(x,y)\in X\times Y$,
$\bigl{(}(\Psi_{1}\circ\Phi_{1})\otimes(\Psi_{2}\circ\Phi_{2})\bigr{)}(x,y)=\Bigl{(}\bigcup_{u\in\Phi_{1}(x)}\Psi_{1}(u)\Bigr{)}\times\Bigl{(}\bigcup_{v\in\Phi_{2}(y)}\Psi_{2}(v)\Bigr{)}=\bigcup_{(u,v)\in\Phi_{1}(x)\times\Phi_{2}(y)}\Psi_{1}(u)\times\Psi_{2}(v),$
which equals $\bigl{(}(\Psi_{1}\otimes\Psi_{2})\circ(\Phi_{1}\otimes\Phi_{2})\bigr{)}(x,y)$ by the definition of composition for correspondences. Identities are preserved since $(\mathrm{id}_{X}\otimes\mathrm{id}_{Y})(x,y)=\{x\}\times\{y\}$.
(5) Naturality of the associator and unitors. For morphisms $\Phi_{1}:X\rightrightarrows X^{\prime}$, $\Phi_{2}:Y\rightrightarrows Y^{\prime}$, $\Phi_{3}:Z\rightrightarrows Z^{\prime}$, we have
$\alpha_{X^{\prime},Y^{\prime},Z^{\prime}}\circ\bigl{(}(\Phi_{1}\otimes\Phi_{2})\otimes\Phi_{3}\bigr{)}=\bigl{(}\Phi_{1}\otimes(\Phi_{2}\otimes\Phi_{3})\bigr{)}\circ\alpha_{X,Y,Z},$
since both sides send $(x,y,z)$ to $\Phi_{1}(x)\times\bigl{(}\Phi_{2}(y)\times\Phi_{3}(z)\bigr{)}$ (after re-bracketing). For unitors, naturality reads
$\lambda_{X^{\prime}}\circ(\mathrm{id}_{I}\otimes\Phi_{1})=\Phi_{1}\circ\lambda_{X},\qquad\rho_{Y^{\prime}}\circ(\Phi_{2}\otimes\mathrm{id}_{I})=\Phi_{2}\circ\rho_{Y},$
because $\{*\}\times\Phi_{1}(x)=\Phi_{1}(x)$ and $\Phi_{2}(y)\times\{*\}=\Phi_{2}(y)$.
(6) Coherence Conditions. The pentagon and triangle commute, as in Top with the Cartesian product; Here $\alpha,\lambda,\rho$ are homeomorphisms (single-valued morphisms), so the standard coherence diagrams commute strictly.
(7) Symmetry. The flip map $\sigma_{X,Y}:X\times Y\to Y\times X$, $\sigma(x,y)=(y,x)$, is a homeomorphism, hence a morphism; The usual hexagon coherence holds. Thus $\mathbf{UHCont}_{\mathrm{c}}$ is symmetric monoidal. Moreover, the braiding is natural: $(\Phi_{2}\otimes\Phi_{1})\circ\sigma_{X,Y}=\sigma_{X^{\prime},Y^{\prime}}\circ(\Phi_{1}\otimes\Phi_{2})$. ∎
Given our previous proofs, it is easy to see that, under conditions (i)-(iv), the Full Conformal Prediction correspondence $\kappa$ is a morphism of $\mathbf{UHCont}_{\text{c}}$. This is because $Y$ is compact Hausdorff, and the graph of $\kappa$ is closed by the Proof of Theorem 5. By the Closed Graph Theorem (Lemma 3), this means that $\kappa$ is upper hemicontinuous and closed-valued; But $Y$ is compact Hausdorff, which implies that $\kappa$ is compact-valued.
Recall from Definition 6 that structure $\mathbf{WMeas}_{\text{uc}}$ has compact Polish spaces with the Borel $\sigma$-algebra for objects, and weakly measurable, uniformly compact-valued correspondences for morphisms. $\mathbf{WMeas}_{\text{uc}}$ is a monoidal category.
Theorem 15 ($\mathbf{WMeas}_{\text{uc}}$ is a Monoidal Category).
The structure $\mathbf{WMeas}_{\text{uc}}$ forms a monoidal category under the Cartesian product.
Proof.
Let the unit object be the singleton space $I\coloneqq\{*\}$, with the trivial topology, and
$(X,\Sigma_{X})\otimes(X^{\prime},\Sigma_{X^{\prime}})\coloneqq(X\times X^{\prime},\Sigma_{X}\otimes\Sigma_{X^{\prime}}),\qquad(F\otimes G)(x,x^{\prime})\coloneqq F(x)\times G(x^{\prime}).$
Since $X,X^{\prime}$ are Polish, $\Sigma_{X}\otimes\Sigma_{X^{\prime}}=\mathcal{B}(X\times X^{\prime})$.
1. Tensor is a bifunctor. For morphisms $F:X\Rightarrow Y$ and $G:X^{\prime}\Rightarrow Y^{\prime}$ the map $(F\otimes G)(x,x^{\prime})=F(x)\times G(x^{\prime})$ is
Uniformly compact. If $K_{F}\subset Y$ and $K_{G}\subset Y^{\prime}$ witness compactness, $K_{F}\times K_{G}$ is compact by Tychonoff, and $(F\otimes G)(x,x^{\prime})\subset K_{F}\times K_{G}$ for all $(x,x^{\prime})$.
Weakly measurable. Let $O=\bigcup_{k\in\mathbb{N}}(B_{k}\times C_{k})$ be an open subset of $Y\times Y^{\prime}$ written as a countable union of basic rectangles (recall that $Y,Y^{\prime}$ are second-countable, so any open set in $Y\times Y^{\prime}$ is a countable union of basic rectangles). Then,
$(F\otimes G)^{-1}(O)=\bigcup_{k\in\mathbb{N}}\bigl{(}\{x:F(x)\cap B_{k}\neq\emptyset\}\times\{x^{\prime}:G(x^{\prime})\cap C_{k}\neq\emptyset\}\bigr{)},$
a countable union of measurable rectangles, hence measurable in $\Sigma_{X}\otimes\Sigma_{X^{\prime}}$.
Functoriality is immediate: $(G_{1}\circ F_{1})\otimes(G_{2}\circ F_{2})=(G_{1}\otimes G_{2})\circ(F_{1}\otimes F_{2})$ because product distributes over union, and $1_{X}\otimes 1_{X^{\prime}}=1_{X\times X^{\prime}}$.
2. Associativity, unit, symmetry. Because $\otimes$ is the cartesian product on objects and the point-wise cartesian product on fibers, we have the canonical homeomorphisms
$\alpha_{X,Y,Z}:(X\times Y)\times Z\longrightarrow X\times(Y\times Z),\qquad\lambda_{X}:\{\ast\}\times X\longrightarrow X,\qquad\rho_{X}:X\times\{\ast\}\longrightarrow X,$
and the symmetry (flip) map
$\sigma_{X,Y}:X\times Y\longrightarrow Y\times X,(x,y)\mapsto(y,x).$
Each of these maps is single-valued and continuous, hence a morphism in $\mathbf{WMeas}_{\text{uc}}$ (singleton fibers are compact). They play the role of associator, left- and right-unitors, and symmetry, respectively.
3. Coherence. The associator, unitors, and symmetry just introduced are the usual homeomorphisms that witness the cartesian product as a symmetric monoidal structure in Set. Consequently the pentagon, triangle, and hexagon coherence diagrams commute exactly as they do in Top (hence also in Meas), and the same proofs apply verbatim in $\mathbf{WMeas}_{\text{uc}}$.
Hence $(\mathbf{WMeas}_{\mathrm{uc}},\otimes,\{\ast\})$ is a symmetric monoidal category. ∎
6.2. Functors Between Categories
Another question that we ask ourselves is whether there is a canonical functor between $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$ (or vice versa). The answer, as the attentive reader may expect, is no. The reason is that there exists a non-functorial gap between measurability and continuity, two different concepts that also motivated us to consider categories $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$ to begin with.
Functors do exist if we focus on restrictions of $\mathbf{WMeas}_{\text{uc}}$ and $\mathbf{UHCont}$. [^14] We defer the study of monoidal functors to future research. In particular, call $\mathbf{WMeas}_{\text{uc,uhc}}\subset\mathbf{WMeas}_{\text{uc}}$ the category whose objects are compact Polish spaces with the Borel $\sigma$-algebra, and whose morphisms are uniformly compact-valued, weakly measurable, and upper hemicontinuous correspondences. Then, the following holds.
Theorem 16 (Faithful Functor between $\mathbf{WMeas}_{\text{uc,uhc}}$ and $\mathbf{UHCont}$).
There exists a faithful functor
$\mathcal{F}:\mathbf{WMeas}_{\text{uc,uhc}}\rightarrow\mathbf{UHCont}$
defined by
  • $\mathcal{F}(X,\tau_{X},\mathcal{B}(\tau_{X}))\coloneqq(X,\tau_{X})$,
  • $\mathcal{F}(\Phi)\coloneqq\Phi$ (viewed as a topological correspondence).
Proof.
We verify that $\mathcal{F}$ is a functor.
(1) Identity. The identity correspondence $\mathrm{id}_{X}(x)=\{x\}$ is trivially compact-valued, upper hemicontinuous, and weakly measurable. Thus, $\mathcal{F}(\mathrm{id}_{X})=\mathrm{id}_{X}$ is a valid morphism in $\mathbf{UHCont}$.
(2) Composition. Let $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ be morphisms in $\mathbf{WMeas}_{\text{uc,uhc}}$.
Uniform compact bound. If $K_{\Psi}\subset Z$ is a uniform compact bound for $\Psi$, then
$(\Psi\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\Psi(y)\subset K_{\Psi}\quad\text{for all }x\in X,$
so $\Psi\circ\Phi$ is uniformly bounded.
Compact fibers. Since $\Phi(x)$ is compact, $\Psi$ has compact values, and $Z$ is compact Hausdorff, upper hemicontinuity of $\Psi$ implies $\operatorname{Gr}(\Psi)\subset Y\times Z$ is closed (by Lemma 3). Hence
$G_{x}\coloneqq\operatorname{Gr}(\Psi)\cap\bigl{(}\Phi(x)\times K_{\Psi}\bigr{)}$
is compact; Its projection onto $Z$ equals $(\Psi\circ\Phi)(x)$, so each $(\Psi\circ\Phi)(x)$ is compact.
Upper hemicontinuity. By (Aliprantis and Border,, 2006, Theorem 17.23), with $\Phi$ and $\Psi$ both u.h.c., the composition $\Psi\circ\Phi$ is u.h.c.
Weak measurability. Preserved under composition as in the proof of Theorem 23 (via Castaing selectors).
Thus $\Psi\circ\Phi$ is again a morphism in $\mathbf{WMeas}_{\text{uc,uhc}}$, and $\mathcal{F}(\Psi\circ\Phi)=\mathcal{F}(\Psi)\circ\mathcal{F}(\Phi)$.
(3) Faithfulness. If $\Phi\neq\Psi$ as correspondences in $\mathbf{WMeas}_{\text{uc,uhc}}$, then $\mathcal{F}(\Phi)\neq\mathcal{F}(\Psi)$ in $\mathbf{UHCont}$. So $\mathcal{F}$ is injective on morphisms. Therefore, $\mathcal{F}$ is a well-defined faithful functor. ∎
Functor $\mathcal{F}$ in Theorem 16 “forgets” the measurable structure but preserves the topological one. The condition of weak measurability ensures that the image $\Phi$ remains a valid morphism in both categories.
It is also worth mentioning that, in general, $\mathcal{F}$ does not have a limit in $\mathbf{UHCont}$ (Perrone,, 2024, Definition 3.1.5). This is because the representability conditions fail: Universal lifting of upper hemicontinuous cones is obstructed by the loss of measurability information via $\mathcal{F}$. This is a consequence of the incompatibility between continuity-based and measurability-based structures, exactly the same structural gap that made defining functor $\mathcal{F}$ nontrivial in the first place.
Let now $\mathbf{UHCont}_{\text{c}}^{\text{P}}\subset\mathbf{UHCont}$ be the category whose objects are compact Polish topological spaces, and whose morphisms are upper hemicontinuous, uniformly compact-valued correspondences. Then, the following holds.
Theorem 17 (Faithful Functor between $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{P}}$ and $\mathbf{WMeas}_{\text{uc}}$).
There exists a faithful functor
$\mathcal{G}:\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{P}}\rightarrow\mathbf{WMeas}_{\text{uc}}$
defined by
  • $\mathcal{G}(X,\tau_{X})\coloneqq(X,\tau_{X},\mathcal{B}(\tau_{X}))$,
  • $\mathcal{G}(\Phi)\coloneqq\Phi$, viewing the correspondence as a measurable one.
Proof.
If $X$ is compact Polish, then $(X,\tau_{X},\mathcal{B}(\tau_{X}))$ is a compact Polish measurable space, so $\mathcal{G}$ is well-defined on objects.
Let $\Phi:X\rightrightarrows Y$ be a morphism in $\mathbf{UHCont}_{\mathrm{c}}^{\mathrm{P}}$, i.e., upper hemicontinuous with uniformly compact values (hence closed values). Since $Y$ is compact Polish, $\Phi$ is automatically uniformly compact-valued (take $K_{\Phi}=Y$). We now show $\Phi$ is weakly measurable.
Weak measurability. Fix open $O\subseteq Y$ and define $\varphi_{O}(y)\coloneqq d(y,O^{c})$, where $d$ is the metric associated with the topology on $Y$. Set $h_{O}(x)\coloneqq\sup_{y\in\Phi(x)}\varphi_{O}(y)$. By Berge’s maximum theorem $h_{O}$ is upper semicontinuous. Hence
$\{x:\Phi(x)\cap O\neq\emptyset\}=\{x:h_{O}(x)>0\}=\bigcup_{m=1}^{\infty}\{x:h_{O}(x)\geq 1/m\},$
a countable union of closed sets; therefore it is measurable. Hence $\Phi$ is weakly measurable. Therefore $\mathcal{G}(\Phi)$ is a morphism in $\mathbf{WMeas}_{\mathrm{uc}}$.
Functoriality. Identities are preserved trivially. For composition, if $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$ are upper hemicontinuous, then $\Psi\circ\Phi$ is upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Because $\Psi$ is upper hemicontinuous with compact values and $Z$ is compact Hausdorff, Lemma 3 implies that $\operatorname{Gr}(\Psi)\subset Y\times Z$ is closed. For each $x\in X$, set $G_{x}\coloneqq\operatorname{Gr}(\Psi)\cap\bigl{(}\Phi(x)\times Z\bigr{)}.$ Since $\Phi(x)$ is compact and $Z$ is compact Hausdorff, the product $\Phi(x)\times Z$ is compact; Hence $G_{x}$ is compact as a closed subset of a compact set. The projection onto $Z$ satisfies
$\pi_{Z}(G_{x})=\bigcup_{y\in\Phi(x)}\Psi(y)=(\Psi\circ\Phi)(x),$
and since $\pi_{Z}:Y\times Z\to Z$ is continuous, $\pi_{Z}(G_{x})$ is compact. Therefore $(\Psi\circ\Phi)(x)$ is compact, and the same uniform bound $K_{\Psi}=Z$ works. Weak measurability of $\Psi\circ\Phi$ follows from the previous argument applied to the composite. Thus $\mathcal{G}(\Psi\circ\Phi)=\mathcal{G}(\Psi)\circ\mathcal{G}(\Phi)$.
Finally, $\mathcal{G}$ is the identity on morphisms, so it is injective on each $\mathrm{Hom}$-set and hence faithful. ∎
We also mention that, just as continuity cannot capture all measurable behavior, measurability cannot reconstruct topological properties like upper hemicontinuity. Therefore, $\mathcal{G}$ typically fails to admit a limit object in $\mathbf{WMeas}_{\text{uc}}$ (Perrone,, 2024, Definition 3.1.5) representing $\text{Cone}(-,\mathcal{G})$, for precisely the same reason $\mathcal{F}$ does: The two categories encode fundamentally distinct structures.
6.3. Monads
We now ask ourselves whether we can find monads on the categories $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$. The general answer is no, but we can find interesting subcategories for which a monad can be found.
Definition 18 (Structure $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$).
Define the structure $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ as follows,
  • Objects: Compact Hausdorff spaces $X$.
  • Morphisms: Compact-valued, upper hemicontinuous $\Phi:X\rightrightarrows Y$.
It is immediate to see that structure $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ is a subcategory of $\mathbf{UHCont}$. It admits a monad.
Theorem 19 (Vietoris lax monad on $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$).
Define an endofunctor $T$ and morphisms $\eta,\nu$ as follows:
  • On objects: $T(X)\coloneqq\mathcal{K}(X)$, the space of nonempty compact subsets of $X$ with the Vietoris topology.
  • On morphisms: for $\Phi:X\rightrightarrows Y$ and $K\in\mathcal{K}(X)$, set
    $T(\Phi)(K)\coloneqq\bigl{\{}L\in\mathcal{K}(Y):L\subset\Phi[K]\bigr{\}},\qquad\Phi[K]\coloneqq\bigcup_{x\in K}\Phi(x).$
  • Unit: $\eta_{X}:X\to\mathcal{K}(X)$, $\eta_{X}(x)=\{x\}$ (single-valued morphism).
  • Multiplication: $\nu_{X}:\mathcal{K}(\mathcal{K}(X))\rightrightarrows\mathcal{K}(X)$,
    $\nu_{X}(\mathcal{A})=\Bigl{\{}\bigcup_{A\in\mathcal{A}^{\prime}}A:\emptyset\neq\mathcal{A}^{\prime}\subseteq\mathcal{A}\text{ compact in }\mathcal{K}(X)\Bigr{\}},$
where $\mathcal{K}(\mathcal{K}(X))$ denotes the hyperspace whose elements are nonempty compact subsets of $\mathcal{K}(X)$. Then, $T$ is an endofunctor on $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$, $\eta$ and $\nu$ are natural (single-valued, resp. set-valued) morphisms, and $(T,\eta,\nu)$ satisfies the lax monad laws (with respect to inclusion of graphs),
$\mathrm{id}\subseteq\nu\circ T\eta,\qquad\nu\circ\eta T=\mathrm{id},\qquad\nu\circ T\nu\subseteq\nu\circ\nu T.$
Moreover, on the subcategory of single-valued continuous maps (CHaus), these inclusions become equalities, recovering the classical Vietoris monad.
Proof.
1) $T$ is a well-defined endofunctor. For a morphism $\Phi:X\rightrightarrows Y$, values of $T(\Phi)$ are nonempty compact subsets of $\mathcal{K}(Y)$. Indeed, $K\mapsto\Phi[K]$ is compact for each compact $K\subset X$ (since $\Phi$ has a closed graph and compact values, and $K$ is compact), so $T(\Phi)(K)=\mathcal{K}(\Phi[K])$ is compact in $\mathcal{K}(Y)$. To see upper hemicontinuity, it suffices to show the graph of $T(\Phi)$ is closed in $\mathcal{K}(X)\times\mathcal{K}(Y)$. Let $K_{m}\to K$ in $\mathcal{K}(X)$ and $L_{m}\to L$ in $\mathcal{K}(Y)$ with $L_{m}\subset\Phi[K_{m}]$. Pick any $y\in L$; Choose $y_{m}\in L_{m}$ with $y_{m}\to y$. For each $m$, pick $x_{m}\in K_{m}$ with $y_{m}\in\Phi(x_{m})$. By compactness of $X$ and the Vietoris convergence $K_{m}\to K$, the net $(x_{m})$ has a subnet $(x_{m_{\ell}})$ converging to some $x\in K$. Upper hemicontinuity of $\Phi$ then gives $y\in\Phi(x)$, hence $y\in\Phi[K]$. As $y\in L$ was arbitrary, $L\subset\Phi[K]$; Thus $(K,L)\in\operatorname{Gr}(T(\Phi))$. Therefore $T(\Phi)$ is compact-valued, upper hemicontinuous. Functoriality is immediate,
$T(\Psi\circ\Phi)(K)=\{L:L\subset(\Psi\circ\Phi)[K]\}$ $=\{L:L\subset\Psi[\Phi[K]]\}$
$=\bigcup_{M\in\mathcal{K}(\Phi[K])}\{L\in\mathcal{K}(Y):L\subset\Psi[M]\}=(T\Psi\circ T\Phi)(K),$
and $T(\mathrm{id})=\mathrm{id}$.
2) $\eta$ and $\nu$ are morphisms (natural in $X$). For $\eta_{X}$, singletons are compact and the map $x\mapsto\{x\}$ is continuous into $\mathcal{K}(X)$, hence a morphism. For $\nu_{X}$, values are compact: If $\mathcal{A}^{\prime}\subset\mathcal{A}$ is compact in $\mathcal{K}(X)$, then $\bigcup_{A\in\mathcal{A}^{\prime}}A$ is a compact subset of $X$. Upper hemicontinuity of $\nu_{X}$ follows from the Vietoris calculus: For a basic open set $\langle U_{1},\dots,U_{m}\rangle\subset\mathcal{K}(X)$,
$\nu_{X}^{-1}\bigl{(}\langle U_{1},\dots,U_{m}\rangle\bigr{)}\supseteq\Bigl{\langle}[U_{1}],\dots,[U_{m}],\langle U_{1}\cup\cdots\cup U_{m}\rangle\Bigr{\rangle},$
and the right hand side is open in $\mathcal{K}(\mathcal{K}(X))$ (see Definition 28); Hence $\nu_{X}$ is upper hemicontinuous.
Naturality (in the lax sense) holds by direct computation: For any $\Phi:X\rightrightarrows Y$,
$(T\Phi)\circ\eta_{X}\supseteq\eta_{Y}\circ\Phi\qquad\text{and}\qquad\nu_{Y}\circ T(T\Phi)\supseteq T(\Phi)\circ\nu_{X},$
since $\{y\}\subset\Phi(x)$ iff $y\in\Phi(x)$, and unions commute with direct images.
3) Lax monad laws. All inclusions are understood as inclusions of correspondences (i.e., of their graphs).
Left unit (lax): For $K\in\mathcal{K}(X)$,
$(T\eta_{X})(K)=\bigl{\{}\mathcal{C}\in\mathcal{K}(\mathcal{K}(X)):\mathcal{C}\subset\eta_{X}[K]\bigr{\}}.$
Thus
$(\nu_{X}\circ T\eta_{X})(K)=\bigl{\{}\textstyle\bigcup_{C\in\mathcal{C}^{\prime}}C:\emptyset\neq\mathcal{C}^{\prime}\subset\eta_{X}[K],\mathcal{C}^{\prime}\in\mathcal{K}(\mathcal{K}(X))\bigr{\}}=\{C\in\mathcal{K}(X):C\subset K\},$
so $\mathrm{id}_{\mathcal{K}(X)}\subseteq\nu_{X}\circ T\eta_{X}$ (take $\mathcal{C}^{\prime}=\eta_{X}[K]$ to recover $K$), but equality need not hold, hence laxity.
Right unit (strict): $\eta_{\mathcal{K}(X)}(K)=\{K\}$ and therefore $(\nu_{X}\circ\eta_{\mathcal{K}(X)})(K)=K$, so $\nu\circ\eta T=\mathrm{id}$ holds as equality.
Associativity (lax): For $\mathscr{B}\in\mathcal{K}(\mathcal{K}(\mathcal{K}(X)))$, elements of $(T\nu_{X})(\mathscr{B})$ are compact families $\mathcal{A}\subset\{\bigcup_{A\in\mathcal{C}}A:\mathcal{C}\in\mathscr{B}\}$, and
$(\nu_{X}\circ T\nu_{X})(\mathscr{B})=\Bigl{\{}\bigcup_{A\in\mathcal{A}^{\prime}}A:\mathcal{A}^{\prime}\subset(T\nu_{X})(\mathscr{B}),\mathcal{A}^{\prime}\in\mathcal{K}(\mathcal{K}(X))\Bigr{\}}.$
Each such union is a union of unions of sets from members of $\mathscr{B}$; Regrouping gives a union of sets from $\bigcup_{\mathcal{C}\in\mathscr{B}}\mathcal{C}$, so
$\nu_{X}\circ T\nu_{X}\subseteq\nu_{X}\circ\nu_{\mathcal{K}(X)},$
establishing the lax associativity.
4) Restriction to CHaus. If we restrict morphisms to single-valued continuous maps $f:X\to Y$, then $T(f)(K)=f[K]$, $\eta$ and $\nu$ are the classical unit and multiplication, and the three lax inclusions above become equalities, recovering the usual Vietoris monad. ∎
The construction in Theorem 19 adapts the Vietoris hyperspace mechanism to the correspondence category $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$. Because morphisms are upper hemicontinuous, compact-valued correspondences, the hyperspace assignment $T$ is a strict endofunctor: For morphisms $\Phi,\Psi$ we have
$T(\Psi\circ\Phi)=T(\Psi)\circ T(\Phi).$
The key point is that for compact $K\subset X$, $\Phi[K]$ is compact, so $T(\Psi\circ\Phi)(K)=\{L:L\subset\Psi[\Phi[K]]\}=(T\Psi\circ T\Phi)(K)$. The resulting structure is lax only at the level of monad laws, where
$\mathrm{id}\subseteq\nu\circ T\eta,\qquad\nu\circ\eta T=\mathrm{id},\qquad\nu\circ T\nu\subseteq\nu\circ\nu T.$
By contrast, on the classical category $\mathbf{CHaus}$ (compact Hausdorff spaces with continuous maps), the Vietoris hyperspace $X\mapsto\mathcal{K}(X)$ together with $T(f)(K)=f[K]$, $\eta(x)=\{x\}$, and $\mu(\mathcal{A})=\bigcup_{A\in\mathcal{A}}A$ defines a strict monad (Garner,, 2020). In that classical setting, the Eilenberg-Moore (EM) category of algebras for the Vietoris monad is equivalent to the category $\mathbf{ContLat}$ of continuous lattices with Scott-continuous lattice homomorphisms Day, (1975); Wyler, (1981).
Hence, the EM-$\mathbf{ContLat}$ equivalence applies to the classical Vietoris monad on $\mathbf{CHaus}$. For our correspondence-based (lax) variant on $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$, we would need a theory of lax algebras; We do not pursue this here.
We now show that, for any $n\in\mathbb{N}$, the (minimal) sufficient conditions for the Full Conformal Prediction diagram to commute in category $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ are the same that make it commute in category $\mathbf{UHCont}$.
Theorem 20 (The Conformal Prediction Diagram Commutes, $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$).
Assume that (i)-(iv) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$.
Proof.
Fix a nonempty subset $\mathscr{F}_{0}\subset\mathscr{F}$ that is closed, uniformly bounded, and equicontinuous on the compact domain $Y^{n}\times Y$. By Arzelà-Ascoli, $\mathscr{F}_{0}$ is compact in the sup-norm, hence $Y^{n}\times\mathscr{F}_{0}$ is compact Hausdorff. In what follows we restrict $\kappa_{\alpha}$ and $\mathsf{CRED}$ to $Y^{n}\times\mathscr{F}_{0}$.
We show that $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}$ are compact-valued and have closed graphs; By Lemma 3 they are upper hemicontinuous and hence morphisms in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. Throughout, $Y$ is compact Hausdorff by (i); Thus, closed subsets of $Y$ are compact.
(1) $\kappa_{\alpha}$ is compact-valued. Fix $(\alpha,y^{n},\psi)$ with $\alpha\notin S_{n+1}$ and set $\beta(\alpha)=\min\{g\in S_{n+1}:g>\alpha\}$ (as in (8)). For each $i$, the set
$\{y\in Y:\psi(y^{n+1}_{-i},y_{i})\geq\psi(y^{n},y)\}$
is closed in $y$ (difference of continuous functions $\geq 0$). Hence $y\mapsto\pi(y,y^{n},\psi)$ is upper semicontinuous as a finite average of indicators of closed sets, and
$\kappa_{\alpha}(y^{n},\psi)=\{y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\}$
is a superlevel set of a u.s.c. map, thus closed in $Y$, hence compact.
(2) $\mathsf{CRED}$ is compact-valued. We can write that
$\mathsf{CRED}(y^{n},\psi)=\bigcap_{A\in\Sigma_{Y}\text{, }A\text{ closed}}\Bigl{\{}P\in\Delta_{Y}:P(A)\leq\overline{\Pi}_{(y^{n},\psi)}(A)\Bigr{\}}.$
For each closed $A$, the map $P\mapsto P(A)$ is weak-continuous on the compact Hausdorff space $\Delta_{Y}$, and $(y^{n},\psi)\mapsto\overline{\Pi}_{(y^{n},\psi)}(A)$ is upper semicontinuous by Lemma 4. Thus each half-space is closed in $\Delta_{Y}$, and arbitrary intersections of closed sets are closed; Hence $\mathsf{CRED}(y^{n},\psi)$ is closed in $\Delta_{Y}$, therefore compact.
(3) $\mathsf{IHDR}_{\alpha}$ is compact-valued. In this compact Hausdorff setting we adopt the closed-sets presentation
$\mathsf{IHDR}_{\alpha}(\mathcal{M})\coloneqq\bigcap\bigl{\{}F\subseteq Y:F\text{ closed and }\underline{P}_{\mathcal{M}}(F)\geq 1-\alpha\bigr{\}}.$
Monotonicity of lower probabilities ensures this presentation agrees with the original one whenever $\mathcal{M}=\mathsf{CRED}(y^{n},\psi)$ and $\pi$ is consonant (by Caprio et al., 2025a (, Proposition 5), $\mathsf{IHDR}_{\alpha}(\mathsf{CRED}(y^{n},\psi))=\{y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\}$, which is closed). By construction the right-hand side is an intersection of closed subsets of $Y$, hence closed; Since $Y$ is compact Hausdorff, it is compact.
Having established compact-valuedness, the same closed-graph arguments used in the proof of Theorem 5 (together with Lemma 3) give upper hemicontinuity of $\kappa_{\alpha}$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}$; Thus they are morphisms of $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. Finally, Caprio et al., 2025a (, Proposition 5) yields the equality $\mathsf{IHDR}_{\alpha}(\mathsf{CRED}(y^{n},\psi))=\kappa_{\alpha}(y^{n},\psi)$ for all admissible $(\alpha,y^{n},\psi)$, i.e. the diagram commutes. ∎
Theorem 20 gives us minimal sufficient conditions so that (i) Full Conformal Prediction is a morphism of a category, $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$, that has a monad, and (ii) the Full Conformal Prediction Diagram commutes in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$. We now show that a subcategory of $\mathbf{WMeas}_{\text{uc}}$ admits a monad.
Theorem 21 (Vietoris monad on $\mathbf{WMeas}_{\mathrm{uc}}$ (with hemicontinuity)).
For this theorem only, strengthen the morphism condition in $\mathbf{WMeas}_{\mathrm{uc}}$ by additionally requiring both upper and lower hemicontinuity (still uniformly compact-valued and weakly measurable). [^15] It is easy to see that this class is still closed under composition. Define the endofunctor $\mathcal{K}$, unit $\eta$, and multiplication $\nu$ by
$\mathcal{K}(X)$ $\coloneqq\{C\subseteq X:C\neq\emptyset,C\text{ compact}\}\text{with the Vietoris topology (Borel $\sigma$-algebra),}$
$\mathcal{K}(\Phi)(C)$ $\coloneqq\{L\in\mathcal{K}(Y):L\subset\overline{\Phi[C]}\},\qquad C\in\mathcal{K}(X),$
$\eta_{X}(x)$ $\coloneqq\bigl{\{}\{x\}\bigr{\}},$
$\nu_{X}(\mathcal{C})$ $\coloneqq\bigl{\{}\textstyle\bigcup_{C\in\mathcal{C}}C\bigr{\}},\qquad\mathcal{C}\in\mathcal{K}(\mathcal{K}(X)).$
Then, $(\mathcal{K},\eta,\nu)$ is a lax monad on $\mathbf{WMeas}_{\mathrm{uc}}$ (under this added hemicontinuity standing assumption). Moreover, if $|X|>1$, the unit $\eta_{X}:X\to\mathcal{K}(X)$ is not an isomorphism (it is not surjective), so the monad is nontrivial.
Proof.
1. $\mathcal{K}$ is an endofunctor.
(Objects). If $X$ is compact Polish, then $\mathcal{K}(X)$ with the Vietoris topology is compact Polish; Its Borel $\sigma$-algebra makes $\mathcal{K}(X)$ an object of $\mathbf{WMeas}_{\mathrm{uc}}$.
(Morphisms). Let $\Phi:X\rightrightarrows Y$ be a morphism (uniformly compact-valued with common compact bound $K_{\Phi}\subset Y$, weakly measurable, and both upper and lower hemicontinuous).
Each fiber of $\mathcal{K}(\Phi)$ is the singleton $\{\overline{\Phi[C]}\}$, which lies in $\mathcal{K}(K_{\Phi})$. Hence, $\mathcal{K}(\Phi)$ is uniformly compact valued with uniform bound $\mathcal{K}(K_{\Phi})\subset\mathcal{K}(Y)$. By a standard hyperspace continuity result (Aubin and Frankowska,, 1990; Beer,, 1993), if $\Phi$ has compact values and is both upper and lower hemicontinuous, then the map
$H:\mathcal{K}(X)\longrightarrow\mathcal{K}(Y),\qquad H(C)\coloneqq\overline{\Phi[C]},$
is continuous for the Vietoris topology. Hence $\mathcal{K}(\Phi)$ is continuous (and therefore weakly measurable) with singleton compact fibers uniformly bounded by $K_{\Phi}$.
Functoriality. $\mathcal{K}(\mathrm{id}_{X})=\mathrm{id}_{\mathcal{K}(X)}$. For $\Psi:Y\rightrightarrows Z$ also hemicontinuous with compact values,
$\bigl{(}\mathcal{K}(\Psi)\circ\mathcal{K}(\Phi)\bigr{)}(C)=\bigl{\{}\overline{\Psi[\overline{\Phi[C]}]}\bigr{\}},\qquad\mathcal{K}(\Psi\circ\Phi)(C)=\bigl{\{}\overline{(\Psi\circ\Phi)[C]}\bigr{\}}.$
We claim $\overline{\Psi[\overline{\Phi[C]}]}=\overline{(\Psi\circ\Phi)[C]}$. One inclusion is monotonicity: $\Psi[\Phi[C]]\subset\Psi[\overline{\Phi[C]}]$, hence $\overline{(\Psi\circ\Phi)[C]}\subset\overline{\Psi[\overline{\Phi[C]}]}$. For the reverse, let $z\in\overline{\Psi[\overline{\Phi[C]}]}$. Then, there are sequences $y_{m}\in\overline{\Phi[C]}$ and $z_{m}\in\Psi(y_{m})$ with $z_{m}\to z$. By compactness, pass to subsequences with $y_{m}\to y\in\overline{\Phi[C]}$. Since $\Psi$ is upper hemicontinuous with compact values, its graph is closed, so $(y,z)\in\mathrm{Gr}(\Psi)$, i.e. $z\in\Psi(y)$. Because $y\in\overline{\Phi[C]}$, there are $u_{k}\in\Phi[C]$ with $u_{k}\to y$. By lower hemicontinuity of $\Psi$ at $y$, for every neighbourhood $W$ of $z$ there is a neighbourhood $N$ of $y$ with $\Psi(x)\cap W\neq\emptyset$ for all $x\in N$; Hence for large $k$ we can pick $v_{k}\in\Psi(u_{k})\cap W$. Thus $W$ meets $\Psi[\Phi[C]]$ for every neighbourhood $W$ of $z$, proving $z\in\overline{\Psi[\Phi[C]]}=\overline{(\Psi\circ\Phi)[C]}$. Therefore $\mathcal{K}(\Psi\circ\Phi)=\mathcal{K}(\Psi)\circ\mathcal{K}(\Phi)$.
2. The unit $\eta$ and multiplication $\nu$. Both $\eta_{X}$ and $\nu_{X}$ have singleton fibers, hence uniformly compact-valued. For weak measurability (indeed continuity), compute on subbasic opens
$\eta_{X}^{-1}\langle U\rangle=U,\qquad\nu_{X}^{-1}\langle U\rangle=\bigl{\{}\mathcal{C}\in\mathcal{K}(\mathcal{K}(X)):\exists C\in\mathcal{C},C\cap U\neq\emptyset\bigr{\}},$
and similarly for miss-type sets $[U]$; these are Vietoris-open. Naturality of $\eta$ and $\nu$ is immediate from $\mathcal{K}(\Phi)[\{x\}]=\{\overline{\Phi[x]}\}$ and $\mathcal{K}(\Phi)\bigl{[}\{\overline{\cup\mathcal{C}}\}\bigr{]}=\{\overline{\cup\mathcal{K}(\Phi)[\mathcal{C}]}\}$.
3. Monad laws. The unit and associativity hold in the lax sense:
$\mathrm{id}_{\mathcal{K}(X)}\subseteq\nu_{X}\circ\mathcal{K}(\eta_{X}),\qquad\nu_{X}\circ\eta_{\mathcal{K}(X)}=\mathrm{id}_{\mathcal{K}(X)},\qquad\nu_{X}\circ\mathcal{K}(\nu_{X})\subseteq\nu_{X}\circ\nu_{\mathcal{K}(X)}.$
Indeed, $\mathcal{K}(\eta_{X})(C)$ consists of all compact families of singletons drawn from $C$, so $\nu_{X}\circ\mathcal{K}(\eta_{X})(C)$ contains $C$ (taking the full family of singletons). For the right unit, $\eta_{\mathcal{K}(X)}(C)=\{\{C\}\}$ and $\nu_{X}$ returns $\{C\}$. For associativity, unions of (compact families of) unions yield no more than the union of all underlying members, giving the inclusion.
4. Non-triviality. If $|X|>1$, then $\eta_{X}:X\to\mathcal{K}(X)$ is not surjective (non-singleton compact sets are not in its image), hence no natural isomorphism $\mathcal{K}\cong\mathrm{Id}$ exists.
Thus $(\mathcal{K},\eta,\nu)$ is a lax monad on $\mathbf{WMeas}_{\mathrm{uc}}$ under the added hemicontinuity assumption. ∎
The construction in Theorem 21 is a version of classical Vietoris monad, adapted to our measurable setting: Objects are compact Polish spaces (with their Borel $\sigma$-algebras), and morphisms are uniformly compact-valued, weakly measurable correspondences. for the hyperspace construction and functoriality of $\mathcal{K}$, we strengthened morphisms to be both upper and lower hemicontinuous (a standard hypothesis ensuring continuity of the induced map on hyperspaces). This added assumption is only used to establish the (lax) monad; All results proved earlier for $\mathbf{WMeas}_{\mathrm{uc}}$—including Theorem 24—do not rely on the monad and remain valid verbatim.
Regarding Eilenberg-Moore algebras: On the category $\mathbf{CHaus}$ of compact Hausdorff spaces, the Vietoris monad (Garner,, 2020) has Eilenberg-Moore category equivalent to the category $\mathbf{ContLat}$ of continuous lattices with Scott-continuous homomorphisms (Day,, 1975; Wyler,, 1981). Since compact Polish spaces are precisely the compact metrizable Hausdorff spaces, restricting this equivalence to metrizable carriers yields the full subcategory of $\mathbf{ContLat}$ consisting of those algebras whose Lawson topology is metrizable.
We have a nice byproduct of focusing on the (sub)categories that we considered in this section.
Proposition 22 (Forgetful Embedding between $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ and $\mathbf{WMeas}_{\text{uc}}$).
There is a faithful functor
$F:\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}\rightarrow\mathbf{WMeas}_{\text{uc}},$
defined as follows.
  1. (Obj)
    For every metrizable compact Hausdorff space $X$, pick any compatible (complete separable) metric $d_{X}$; The resulting topological space is compact Polish. Set $F(X)\coloneqq(X,d_{X})$ equipped with its Borel $\sigma$–algebra.
  2. (Mor)
    For a morphism $\Phi:X\rightrightarrows Y$ in $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ (i.e. compact–valued and upper hemicontinuous), let
    $F(\Phi)\coloneqq\Phi\qquad(\text{same graph and fibers}).$
With these assignments, $F$ preserves identities and composition, and on every hom-set it is injective. Hence, $F$ is faithful.
The functor $F$ in Proposition 22 is defined on the full subcategory of $\mathbf{UHCont}^{\mathrm{cH}}_{\mathrm{c}}$ with metrizable objects, since $(X,\mathcal{B}(X))$ must be compact Polish to lie in $\mathbf{WMeas}_{\mathrm{uc}}$. When working within that metrizable subcategory (as we do here), the statement holds verbatim. The choice of compatible metric on $X$ does not affect $\mathcal{B}(X)$.
Proof of Proposition 22.
1. $F(\Phi)$ is a morphism in $\mathbf{WMeas}_{\text{uc}}$. Because $Y$ is compact, $\Phi(x)\subseteq Y$ for all $x$, so $K_{\Phi}\coloneqq Y$ is a common compact bound; Thus $F(\Phi)$ is uniformly compact-valued. To show weak measurability, fix an open $O\subseteq Y$ and define $\varphi_{O}(y)\coloneqq d_{Y}\bigl{(}y,(Y\setminus O)\bigr{)}$, where $d_{Y}$ is any compatible (complete) metric on $Y$. Set
$h_{O}(x)\coloneqq\sup_{y\in\Phi(x)}\varphi_{O}(y).$
Since $\Phi$ is upper hemicontinuous with compact values, Berge’s maximum theorem implies $h_{O}$ is upper semicontinuous; Hence
$\{x:\Phi(x)\cap O\neq\emptyset\}=\{x:h_{O}(x)>0\}$
is open (in particular, Borel). This is precisely the weak measurability condition for $F(\Phi)$.
2. Functoriality. $F$ leaves the underlying correspondence unchanged, so $F(\mathrm{id}_{X})=\mathrm{id}_{F(X)}$ and $F(\Psi\circ\Phi)=F(\Psi)\circ F(\Phi)$, for all composable $\Phi,\Psi$.
3. Faithfulness. On each hom-set, $F$ acts as the identity map $\Phi\mapsto\Phi,$ hence it is injective.
4. Domain of $F$. If every object of $\mathbf{UHCont}^{\mathrm{cH}}_{\text{c}}$ is assumed to be metrizable (equivalently, second-countable), then the object assignment is defined on the whole category. Otherwise, $F$ can still be applied to the metrizable subcategory, for which the conclusion holds verbatim. ∎
Alas, there is no faithful functor $E:\mathbf{WMeas}_{\mathrm{uc}}\to\mathbf{UHCont}$ that (i) is the identity on objects and (ii) fixes all Dirac morphisms $\delta_{y}:\{*\}\to Y$, $\delta_{y}(*)=\{y\}$. To see this, take $X=Y=[0,1]$ (usual topology and Borel $\sigma$-algebra) and define
[…]
Both $F$ and $G$ are morphisms in $\mathbf{WMeas}_{\mathrm{uc}}$ (singleton-valued and weakly measurable), with $F\neq G$. For each $x\in X$, let $\delta_{x}:\{*\}\to X$ be the Dirac morphism. Then $F\circ\delta_{x}=\delta_{F(x)}$ and $G\circ\delta_{x}=\delta_{G(x)}$. By the hypotheses and functoriality,
$E(F)\circ\delta_{x}=E(F\circ\delta_{x})=E(\delta_{F(x)})=\delta_{F(x)},$
so $E(F)(x)=\{F(x)\}$ for all $x$. Thus, $E(F)$ coincides pointwise with $F$, but $F$ is not upper hemicontinuous (its graph is not closed since $\mathbb{Q}$ and $\mathbb{R}\setminus\mathbb{Q}$ are both dense). This contradicts that $E(F)$ must be a morphism in $\mathbf{UHCont}$, so no such faithful functor $E$ exists.
7. Conclusion
In this paper, we took a Category Theory route to show that Full Conformal Prediction exhibits intrinsic UQ capabilities. This is a result of framing it as a morphism of two categories, UHCont and $\textbf{WMeas}_{\text{uc}}$, and of showing that it is part of a diagram that commutes. As byproduct, we were able to prove that CP methods bridge Bayesian, frequentist, and imprecise approaches to statistical prediction. We also studied the properties of such categories, and we pointed out how a Conformal Prediction Region can be seen as a functor image.
This work opens up the study of Categorical Conformal Prediction, and hints at many unsolved problems. For example, it is an open question how to leverage the structure of the functional-analysis-flavored categories UHCont and $\textbf{WMeas}_{\text{uc}}$ to learn further aspects of the CP methodology. In particular, because the morphisms of these categories are correspondences, we are interested in what the Kleisli categories (Kleisli,, 1962, 1965), (Perrone,, 2024, Section 5.1) arising from the subcategories of $\mathbf{UHCont}$ and $\mathbf{WMeas}_{\text{uc}}$ admitting a monad, can tell us about Conformal Prediction.
In addition, it is left to inspect whether studying CP as a morphism of different categories can allow us to remove the need for assumptions altogether, in particular compactness of the state space $Y$.
Finally, it is left for the future to show whether our results extend to Split Conformal Prediction, and to any other extension of classical Conformal Prediction that ML researchers introduced in recent years.
Acknowledgments
We wish to express our gratitude to Sam Staton and Paolo Perrone for insightful discussions on the possibility of formulating Conformal Prediction in the Category Theory framework. We are also extremely grateful to Sam Staton for his invitation to the University of Oxford, where the idea in the present paper germinated. We are indebted to Nicola Gambino for his assistance in improving the proof of Theorem 15, and for suggesting that we investigate the category of algebras for the Vietoris monad.
Finally, we thank Yusuf Sale and Eyke Hüllermeier for stimulating discussions on the Uncertainty Quantification capabilities of Conformal Prediction, Sangwoo Park for his idea that led to Proposition 10, and Alessandro Zito for his patience in helping us shape the abstract and introduction.
Appendix A Conformal Prediction as a Weakly measurable Morphism
This section shows that $\mathbf{WMeas}_{\text{uc}}$ is a category, and proves that, under minimal assumptions, the (Full) Conformal Prediction correspondence $\kappa$ is a morphism of $\mathbf{WMeas}_{\text{uc}}$ as well, embedded in a commuting diagram.
Recall that $\mathbf{WMeas}_{\text{uc}}$ is a structure with
  • Objects: Compact Polish spaces $(X,\Sigma_{X})$, with $\Sigma_{X}=\mathcal{B}(X)$, the Borel $\sigma$-algebra.
  • Morphisms: Weakly measurable (see Definition 26 in Appendix C), uniformly compact-valued (see Definition 27 in Appendix C) correspondences $\Phi:X\rightrightarrows Y$.
Theorem 23 ($\mathbf{WMeas}_{\text{uc}}$ is a category).
The structure $\mathbf{WMeas}_{\text{uc}}$ defines a category.
Proof.
We verify the four axioms.
(1) and (4) Identity morphisms and Identity laws. For any object $X$, $\operatorname{id}_{X}(x)=\{x\}$ is nonempty compact. Put $K_{\operatorname{id}_{X}}=X$ (compact by hypothesis); The uniformly compact-valued condition holds. For every open $O\subseteq X$ we have $\operatorname{id}_{X}^{-1}(O)=O\in\Sigma_{X}$, so $\operatorname{id}_{X}$ is a morphism. The usual set-theoretic equalities $\Phi\circ\operatorname{id}_{X}=\Phi$ and $\operatorname{id}_{Y}\circ\Phi=\Phi$ prove the identity laws.
(2) Associativity. For morphisms $\Phi\colon X\rightrightarrows Y$, $\Psi\colon Y\rightrightarrows Z$ and $\Theta\colon Z\rightrightarrows W$, the associativity of union gives $\bigl{(}(\Theta\circ\Psi)\circ\Phi\bigr{)}(x)=(\Theta\circ(\Psi\circ\Phi))(x)$, for every $x\in X$.
(3) Closure under composition. Let $\Phi\colon X\rightrightarrows Y$ and $\Psi\colon Y\rightrightarrows Z$ be morphisms. Write $K_{\Phi}\subseteq Y$ and $K_{\Psi}\subseteq Z$ for the uniform compact bounds, see Definition 27.(ii).
Compactness of fibers. For each $x\in X$,
$(\Psi\circ\Phi)(x)=\bigcup_{y\in\Phi(x)}\Psi(y)\subseteq\bigcup_{y\in\Phi(x)}K_{\Psi}\subseteq K_{\Psi}.$
Since $K_{\Psi}$ is compact and the union is taken inside it, $(\Psi\circ\Phi)(x)$ is a closed subset of a compact space, hence compact. [^16] For simplicity, here we implicitly assume that $|\Phi(x)|<\infty$, otherwise we would also need upper hemicontinuity of the morphisms, and we would end up working with $\mathbf{WMeas}_{\text{uc,uhc}}$. Thus compactness holds with the same compact bound $K_{\Psi}$, so the uniformly compact-valued condition also holds.
Weak measurability. Fix an open $O\subseteq Z$ and set $H\coloneqq\Psi\circ\Phi$. Since $X$ is standard Borel and $Y$ is Polish, $\Phi$ being weakly measurable with closed values implies that its graph $\operatorname{Gr}(\Phi)\subset X\times Y$ is Borel. By the Lusin-Novikov Uniformization Theorem, there exist Borel maps $f_{j}:X\to Y$, $j\in\mathbb{N}$, such that, for each $x$, the finite set $\Phi(x)$ is exactly $\{f_{j}(x):1\leq j\leq N(x)\}$ for some finite $N(x)$, and we may fix $f_{j}(x)=f_{1}(x)$ for $j>N(x)$.
Then, for every $x$,
$H(x)\cap O\neq\emptyset\iff\exists j\text{ such that }\Psi\bigl{(}f_{j}(x)\bigr{)}\cap O\neq\emptyset,$
and therefore
$H^{-1}(O)=\bigcup_{j=1}^{\infty}f_{j}^{-1}\bigl{(}\Psi^{-1}(O)\bigr{)}.$
Since each $f_{j}$ is Borel and $\Psi$ is weakly measurable, $\Psi^{-1}(O)\in\Sigma_{Y}$ and hence $f_{j}^{-1}\bigl{(}\Psi^{-1}(O)\bigr{)}\in\Sigma_{X}$. A countable union of Borel sets is Borel, so $H^{-1}(O)\in\Sigma_{X}$. Thus $H$ is weakly measurable, completing closure. Therefore, $\mathbf{WMeas}_{\text{uc}}$ satisfies the axioms of a category. ∎
Similarly to what we did in Section 3, we now ask ourselves what are the minimal conditions that make correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ uniformly compact-valued and weakly measurable. As we shall see in Theorem 24, they are the same as in Section 3, with (i), (ii), and (iv) slightly strengthened to
  1. (i’)
    $Y$ is compact metric and $\Sigma_{Y}$ is its Borel $\sigma$-field.
  2. (ii’)
    Each $\psi\in\mathscr{F}$ is jointly continuous on $Y^{n}\times Y$; $\mathscr{F}$ is compact Polish in the uniform ($\sup$-norm) topology.
  1. (iv’)
    When working with $\mathbf{WMeas}_{\text{uc}}$, $\Delta_{Y}$ denotes the set of countably additive probability measures on $Y$, endowed with the weak topology inherited from $C(Y)^{\star}$. [^17] This is needed to make $\Delta_{Y}$ metrizable: see proof of weak measurability of $\mathsf{CRED}$ in the next page.
Since $Y$ is compact metric, $C(Y)$ is separable; Hence $\Delta_{Y}$ is compact metrizable. As a consequence, $\mathscr{C}$ is compact metrizable (the Vietoris topology coincides with the Hausdorff metric topology on nonempty closed subsets of a compact metric space).
In Remark 4, we show that (i’) is needed to make $\mathsf{IHDR}$ weakly measurable, and (ii’) to make $Y^{n}\times\mathscr{F}$ Polish. We are now ready for the main result of this section.
Theorem 24 (The Conformal Prediction Diagram Commutes, $\mathbf{WMeas}_{\text{uc}}$).
Assume that (i’), (ii’), (iii), (iv’) hold, and that the conformal transducer $\pi$ is consonant. Consider the restrictions $\kappa_{\alpha}$ and $\mathsf{IHDR}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$. That is, the Full CP Diagram
$Y^{n}\times\mathscr{F}$$\mathscr{C}$$\Sigma_{Y}$$\scriptstyle{\mathsf{CRED}}$$\scriptstyle{\kappa_{\alpha}}$$\scriptstyle{\mathsf{IHDR}_{\alpha}}$
commutes in $\mathbf{WMeas}_{\text{uc}}$.
Notice that, in the Full CP Diagram, $\Sigma_{Y}=\mathcal{B}(Y)$ by Definition 6.
Proof of Theorem 24.
Fix any $n\in\mathbb{N}$. We first show that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$ are uniformly compact-valued and weakly measurable.
Uniformly compact-valued. All three images lie inside the fixed compact spaces $Y$ (by (i’)) or $\Delta_{Y}$ (by (iv’)); We are now left to show that they themselves are compact-valued.
Notice that, under consonance, for all $y^{n}\in Y^{n}$ and all $\psi\in\mathscr{F}$, $\mathsf{CRED}(y^{n},\psi)$ is a credal set, so a closed subset of the compact space $\Delta_{Y}$, and so compact in turn.
Also, for all $\alpha\in[0,1]\setminus S_{n+1}$ and all credal sets $\mathcal{M}$, $\mathsf{IHDR}(\alpha,\mathcal{M})$ is an intersection of closed sets, and so closed itself. In turn, it is a closed subset of the compact space $Y$, so it is itself compact.
An argument similar to point (i) in the Proof of Theorem 5 shows that, for all $y^{n}\in Y^{n}$, all $\psi\in\mathscr{F}$, and all $\alpha\in[0,1]\setminus S_{n+1}$, the set $\kappa(\alpha,y^{n},\psi)=\{y:\pi(y,y^{n},\psi)>\alpha\}=\{y:\pi(y,y^{n},\psi)\geq\beta\}$ is closed, where $\beta\equiv\beta(\alpha)$ was defined in (8). Indeed, it is a superlevel set of $\pi(\cdot,y^{n},\psi)$, and the latter is upper semicontinuous in $\psi$ as seen in the proof of Lemma 4. In turn, the map $y\mapsto\pi(y,y^{n},\psi)$ is u.s.c. on the compact $Y$. Hence, $\{y:\pi(y,y^{n},\psi)\geq\beta(\alpha)\}$; $\kappa(\alpha,y^{n},\psi)$ is then a closed subset of the compact space $Y$, so it is itself compact.
Weak measurability of $\kappa$. Fix an open $G\subseteq Y$ and let $\{\overline{B}_{k}\}_{k\in\mathbb{N}}$ be a countable family of closed balls with rational centers and radii that forms a closed base of $Y$. Because $G$ is open,
$\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\iff\exists k\text{ such that }\overline{B}_{k}\subseteq G\text{and}\sup_{y\in\overline{B}_{k}}\pi(y,y^{n},\psi)\geq\beta(\alpha).$
For each $k$, define
$u_{k}(y^{n},\psi)\coloneqq\sup_{y\in\overline{B}_{k}}\pi(y,y^{n},\psi).$
By Lemma 4 (applied to the closed set $\overline{B}_{k}$), $u_{k}$ is upper semicontinuous (hence Borel) on $Y^{n}\times\mathscr{F}$. Recall that $\Lambda\coloneqq[0,1]\setminus S_{n+1}$; Since $\beta:\Lambda\to S_{n+1}$ is piecewise constant (Borel), we have
$\bigl{\{}(\alpha,y^{n},\psi):\kappa(\alpha,y^{n},\psi)\cap G\neq\emptyset\bigr{\}}=\bigcup_{\overline{B}_{k}\subseteq G}\bigcup_{s\in S_{n+1}}\Bigl{(}\{\alpha:\beta(\alpha)=s\}\times\{(y^{n},\psi):u_{k}(y^{n},\psi)\geq s\}\Bigr{)},$
a countable union of Borel sets. Hence $\kappa$ is weakly measurable.
Weak measurability of $\mathsf{CRED}$. Let $\mathscr{K}$ be the countable family of all finite unions of closed balls in $Y$ with rational centers and radii. Enumerate $\mathscr{K}=(K^{j})_{j\in\mathbb{N}}$. For each $j$, set
$u_{j}(y^{n},\psi)\coloneqq\sup_{y\in K^{j}}\pi(y,y^{n},\psi).$
Each $u_{j}$ is u.s.c. on $Y^{n}\times\mathscr{F}$ (finite maxima of u.s.c.). Fix an open $G\subseteq\Delta_{Y}$. Since $\Delta_{Y}$ is compact metric, write $G=\bigcup_{m=1}^{\infty}K_{m}$ with $K_{m}\uparrow G$ compact. For each $j$ and $\ell\in\mathbb{N}$, set the closed thickening $K^{j,\ell}\coloneqq\{y\in Y:\mathrm{dist}(y,K^{j})\leq 1/\ell\}$, where $\mathrm{dist}(\cdot)$ is any metric on $Y$, and define
$v_{j,\ell}(y^{n},\psi)\coloneqq\sup_{y\in K^{j,\ell}}\pi(y,y^{n},\psi).$
By Lemma 4 (applied to the closed set $K^{j,\ell}$), each $v_{j,\ell}$ is upper semicontinuous on $Y^{n}\times\mathscr{F}$, and $v_{j,\ell}\downarrow u_{j}$ pointwise as $\ell\to\infty$ (nested closed sets).
For each $j$ and $\ell$, choose $\phi_{j,\ell}\in C(Y,[0,1])$ with $\phi_{j,\ell}\downarrow\mathbbm{1}_{K^{j}}$ pointwise and $\operatorname{supp}(\phi_{j,\ell})\subseteq K^{j,\ell}$; For instance, $\phi_{j,\ell}(y)=\max\{0,1-\ell\mathrm{dist}(y,K^{j})\}$, so that $\phi_{j,\ell}\leq\mathbbm{1}_{K^{j,\ell}}$. Replace $\widetilde{C}_{m,r,\ell}$ by
$\widehat{C}_{m,r,\ell}\coloneqq\Bigl{\{}((y^{n},\psi),P)\in(Y^{n}\times\mathscr{F})\times K_{m}:\textstyle\int\phi_{j,\ell}\text{d}P\leq v_{j,\ell}(y^{n},\psi)\text{, for }1\leq j\leq r\Bigr{\}}.$
Since $((y^{n},\psi),P)\mapsto\int\phi_{j,\ell}\text{d}P$ is continuous in $P$ and $(y^{n},\psi)\mapsto v_{j,\ell}(y^{n},\psi)$ is u.s.c., the map $((y^{n},\psi),P)\mapsto\int\phi_{j,\ell}\text{d}P-v_{j,\ell}(y^{n},\psi)$ is lower semicontinuous; Hence each $\widehat{C}_{m,r,\ell}$ is closed. With
$\widehat{C}_{m,r}\coloneqq\bigcap_{\ell=1}^{\infty}\widehat{C}_{m,r,\ell},\qquad\widehat{\Pi}_{m,r}\coloneqq\mathrm{pr}_{Y^{n}\times\mathscr{F}}(\widehat{C}_{m,r}),$
where $\mathrm{pr}_{Y^{n}\times\mathscr{F}}(\cdot)$ denotes the projection operator on $Y^{n}\times\mathscr{F}$, the same compactness/projection argument yields that $\widehat{\Pi}_{m,r}$ is compact.
For $(y^{n},\psi)$, set $\widehat{S}_{m,r}(y^{n},\psi)=\{P\in K_{m}:((y^{n},\psi),P)\in\widehat{C}_{m,r}\}$. Then, $\{\widehat{S}_{m,r}\}_{r\geq 1}$ is decreasing and
$\bigcap_{r\geq 1}\widehat{S}_{m,r}(y^{n},\psi)=\mathsf{CRED}(y^{n},\psi)\cap K_{m}.$
Indeed, if $P\in\mathsf{CRED}(y^{n},\psi)$, then for every $j,\ell$, $\phi_{j,\ell}\leq\mathbbm{1}_{K^{j,\ell}}$ gives $\int\phi_{j,\ell}\text{d}P\leq P(K^{j,\ell})\leq v_{j,\ell}(y^{n},\psi)$, so $P\in\widehat{S}_{m,r}$ for all $r$. Conversely, if $P\in\bigcap_{r,\ell}\widehat{S}_{m,r}(y^{n},\psi)$, then for all $j,\ell$, $\int\phi_{j,\ell}\text{d}P\leq v_{j,\ell}(y^{n},\psi)$. Letting $\ell\to\infty$ and using $\phi_{j,\ell}\downarrow\mathbbm{1}_{K^{j}}$ (bounded convergence) and $v_{j,\ell}\downarrow u_{j}$ yields $P(K^{j})\leq u_{j}(y^{n},\psi)$ for all $j$; By the finite-union approximation of closed sets used above, $P\in\mathsf{CRED}(y^{n},\psi)$. Therefore,
$\{(y^{n},\psi):\mathsf{CRED}(y^{n},\psi)\cap K_{m}\neq\emptyset\}=\bigcap_{r=1}^{\infty}\widehat{\Pi}_{m,r},$
which is Borel; Hence $\mathsf{CRED}$ is weakly measurable.
Weak measurability of $\mathsf{IHDR}_{\alpha}$. Let $\mathcal{U}$ be a fixed countable base of open sets of $Y$ (e.g., open balls with rational centers/radii). For $U\in\mathcal{U}$ define
$S_{U}(\mathcal{M})\coloneqq\sup_{P\in\mathcal{M}}P(U),\qquad\mathcal{M}\in\mathscr{C}.$
Because $Y$ is compact metric, the weak topology on $\Delta_{Y}$ coincides with the weak topology of measures; By the Portmanteau theorem, $P\mapsto P(U)$ is lower semicontinuous on $\Delta_{Y}$ for every open $U$. On the hyperspace $\mathscr{C}$, the map $\mathcal{M}\mapsto\sup_{P\in\mathcal{M}}f(P)$ is lower semicontinuous whenever $f$ is lower semicontinuous. [^18] One way to see this on a compact metric domain: Approximate $f$ from below by an increasing sequence of continuous functions $f_{m}\uparrow f$; Each $\mathcal{M}\mapsto\sup_{P\in\mathcal{M}}f_{m}(P)$ is continuous (uniform continuity of $f_{m}$ and Hausdorff convergence), and their pointwise supremum is lower semicontinuous. Hence each $S_{U}$ is lower semicontinuous on $\mathscr{C}$, and so the sets $\{\mathcal{M}:S_{U}(\mathcal{M})>\alpha\}$ are open in $\mathscr{C}$. We claim the following characterization,
$y\in\mathsf{IHDR}_{\alpha}(\mathcal{M})\quad\Longleftrightarrow\quad S_{U}(\mathcal{M})>\alpha\text{ for every }U\in\mathcal{U}\text{ with }y\in U.$ (16)
Indeed, $\underline{P}_{\mathcal{M}}(A)=1-\overline{P}_{\mathcal{M}}(A^{c})$ where $\overline{P}_{\mathcal{M}}(B)=\sup_{P\in\mathcal{M}}P(B)$. Thus $y\in\mathsf{IHDR}_{\alpha}(\mathcal{M})$ if and only if for every measurable $A$ with $y\notin A$ we have $\underline{P}_{\mathcal{M}}(A)<1-\alpha$, i.e., for every open neighborhood $U\ni y$, $\overline{P}_{\mathcal{M}}(U)>\alpha$. Since $\mathcal{U}$ is a base, it suffices to check $U\in\mathcal{U}$. Now, let $G\subseteq Y$ be open. Using (16),
$\{\mathcal{M}\in\mathscr{C}:\mathsf{IHDR}_{\alpha}(\mathcal{M})\cap G\neq\emptyset\}=\bigcup_{\begin{subarray}{c}U\in\mathcal{U}\text{, }U\subseteq G\end{subarray}}\{\mathcal{M}\in\mathscr{C}:S_{U}(\mathcal{M})>\alpha\},$
a countable union of open sets in $\mathscr{C}$. Hence $\mathsf{IHDR}_{\alpha}$ is weakly measurable.
We have now shown that $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}_{\alpha}$ are uniformly compact-valued and weakly measurable, and thus morphisms of $\mathbf{WMeas}_{\text{uc}}$. Fix now any $\alpha\in[0,1]\setminus S_{n+1}$. It is easy to see that the restriction $\kappa_{\alpha}$ remains uniformly compact-valued and weakly measurable.
Now, recall that the product of two Polish spaces is Polish. Since $Y$ is compact metric by (i’), it is also measurable Polish; $\mathscr{C}$ is compact Polish by (iv’); $Y^{n}\times\mathscr{F}$ is compact Polish in the product topology by (i’) and (ii’). Hence, these spaces are all objects of $\mathbf{WMeas}_{\text{uc}}$.
The fact that the Full Conformal Prediction Diagram commutes in $\mathbf{WMeas}_{\text{uc}}$, then, is an immediate consequence of Caprio et al., 2025a (, Proposition 5). ∎
The same considerations that we put forth in Section 3 on the importance of Theorem 5, still hold here for Theorem 24. The following remark shows that the conditions in Theorem 24 are truly minimal.
Remark 4 (Minimality of Hypotheses (i’), (ii’), (iii), (iv’)).
Similarly to Remark 1, we prove that each hypothesis is indispensable for the simultaneous uniformly compact-valuedness and weak measurability of the three correspondences $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$.
  • Drop (i’) (compactness of $Y$). Let $Y=\mathbb{R}$, $n=1$, $\psi(y^{1},y)=-y$, $\alpha=\frac{3}{4}$. Then $\kappa(\alpha,y^{1},\psi)=[y^{1},\infty)$ is not compact; Uniform compact-valuedness fails.
  • Drop (ii’) (Polishness of $\mathscr{F}$).
    (a) Completeness/Polishness. Let $\mathscr{F}$ be the set of polynomials on $[0,1]$ with the sup-norm subspace topology from $C([0,1])$. Write $\mathscr{F}=\bigcup_{n\geq 0}V_{n}$, where $V_{n}$ is the finite-dimensional space of degree-$\leq n$ polynomials. Each $V_{n}$ is closed and nowhere dense in $C([0,1])$, hence $\mathscr{F}$ is a countable union of nowhere dense sets in its own topology, i.e. $\mathscr{F}$ is meagre in itself. No nonempty Polish space is meagre in itself (Baire), so $\mathscr{F}$ is not Polish. Consequently $Y^{n}\times\mathscr{F}$ is not Polish and is not an object of $\mathbf{WMeas}_{\mathrm{uc}}$.
    (b) Continuity. Let $Y=[0,1]$, $n=1$, and fix $\alpha=\tfrac{3}{4}$ (so $\beta(\alpha)=1$). Choose an analytic non-Borel set $B\subset[0,1]$ and a nonempty open interval $O\subset[0,1]\setminus B$; Set $A\coloneqq B\cup O$ (analytic, non-Borel). Define the score
    $\psi(y^{1},y)\coloneqq\mathbbm{1}[y\notin A].$
    For $x=y^{1}$, check (using the Full CP transducer and $\beta(\alpha)=1$) that
    $\kappa\!\left(\tfrac{3}{4},x,\psi\right)=\begin{cases}[0,1],&x\in A,\\[2.0pt] [0,1]\setminus A,&x\notin A.\end{cases}$
    Taking $G\coloneqq O\subset A$ (open) we obtain
    $\{x:\kappa(\tfrac{3}{4},x,\psi)\cap G\neq\emptyset\}=A,$
    which is analytic non-Borel. Hence $\kappa$ fails to be weakly measurable.
  • Drop (iii) (no-tie rule). Same $\alpha=1$ argument in Remark 1.
  • Drop (iv’) (weak on $\Delta_{Y}$ and Vietoris on $\mathscr{C}$).
    (i) Put total variation on $\Delta_{Y}$. Equip $\Delta_{Y}$ with the total variation (TV) metric instead of the weak topology. Let $\psi\equiv 0$. Then, all nonconformity scores are equal, hence $\pi\equiv 1$ and
    $\mathsf{CRED}(y^{n},\psi)=\Delta_{Y}\quad\text{ for every }y^{n}.$
    But $(\Delta_{Y},\mathrm{TV})$ is not compact; Thus $\mathsf{CRED}(y^{n},\psi)$ itself is not compact, and $\mathsf{CRED}$ is not compact-valued (hence not uniformly compact-valued). Therefore the weak topology on $\Delta_{Y}$ in (iv’) is essential.
    (ii) Keep weak on $\Delta_{Y}$ but replace Vietoris on $\mathscr{C}$ by the lower Vietoris (hit) topology. Let $Y=[0,1]$, $A=\{0\}$, $P_{0}=\delta_{0}$, $P_{k}=\delta_{1/k}$, and $\mathcal{M}_{k}=\operatorname{co}(\{P_{0},P_{k}\})$, $\mathcal{M}_{\infty}=\{P_{0}\}$. Then $\mathcal{M}_{k}\to\mathcal{M}_{\infty}$ in the hit topology, but $\underline{P}_{\mathcal{M}_{k}}(A)=0$ while $\underline{P}_{\mathcal{M}_{\infty}}(A)=1$, so $\mathcal{M}\mapsto\underline{P}_{\mathcal{M}}(A)$ is not lower semicontinuous; The weak measurability argument for $\mathsf{IHDR}$ breaks.
Of course, consonance is needed to apply Caprio et al., 2025a (, Proposition 5). Hence (i’), (ii’), (iii), (iv’), together with consonance, are jointly minimal for the uniform compact-valuedness and weak measurability of $\kappa$, $\mathsf{CRED}$, and $\mathsf{IHDR}$. $\triangle$
Appendix B Further Proofs
Proof of Lemma 4.
First, we show that, if a generic set $K$ is compact and a generic map $f:K\to\mathbb{R}$ is upper semicontinuous (u.s.c.), then $f$ attains its maximum on $K$. Choose $x_{m}$ with $f(x_{m})\geq\sup_{K}f-\tfrac{1}{m}$. A compactness subsequence $x_{m_{k}}\to x_{\star}\in K$ exists. Upper semicontinuity gives $f(x_{\star})\geq\limsup f(x_{m_{k}})=\sup_{K}f.$
Now, we introduce auxiliary maps, and show their continuity. For $i\in\{1,\dots,n+1\}$, put
$g_{i}(y_{n+1},y^{n},\psi)\coloneqq\psi\bigl{(}y^{n+1}_{-i},y_{i}\bigr{)}-\psi\bigl{(}y^{n},y_{n+1}\bigr{)}.$
Because (a) the evaluation $\psi\mapsto\psi(y^{n},y)$ is continuous in the sup-norm, and (b) coordinate deletion $y^{n+1}\mapsto(y^{n+1}_{-i},y_{i})$ is continuous, each $g_{i}$ is continuous on $Y^{n+1}\times\mathscr{F}$.
Let us now show the upper semicontinuity of $\pi_{\psi}$. Fix $(y^{n+1},\psi)$. The set $\{g_{i}\geq 0\}$ is closed; For $k=0,\dots,n+1$ define
$R_{k}\coloneqq\Bigl{\{}(y^{n+1},\psi):\pi_{\psi}(y_{n+1},y^{n})\geq\tfrac{k}{n+1}\Bigr{\}}.$
At least $k$ of the inequalities $g_{i}\geq 0$ must hold, so $R_{k}=\bigcup_{\lvert I\rvert=k}\bigcap_{i\in I}\{g_{i}\geq 0\}$, a finite union of finite intersections of closed sets; Hence $R_{k}$ is closed and $\pi_{\psi}$ is u.s.c.
We are now ready to prove the main statement. Fix any closed set $A\in\Sigma_{Y}$. By (i), it is also compact. Let $\bigl{(}y^{n}_{m},\psi_{m}\bigr{)}_{m\geq 1}$ converge to $(y^{n},\psi)$ and set $L\coloneqq\limsup_{m\to\infty}\overline{\Pi}_{(y^{n}_{m},\psi_{m})}(A)$. [^19] If $Y$ is not first countable, we can switch from a sequence to a net argument. We must show $L\leq\overline{\Pi}_{(y^{n},\psi)}(A)$.
Step 1 (Discretization). Because $\pi_{\psi}$ takes values in the finite set $\{0,\tfrac{1}{n+1},\dots,1\}$, there is an integer $k\in\{0,\dots,n+1\}$ and a subsequence (still indexed by $m$) with $\overline{\Pi}_{(y^{n}_{m},\psi_{m})}(A)=k/(n+1)$, for all $m$.
Step 2 (Select Maximizers). By the attained maximum result that we presented earlier, pick $y_{n+1}^{(m)}\in A$ such that $\pi_{\psi_{m}}\bigl{(}y_{n+1}^{(m)},y^{n}_{m}\bigr{)}=k/(n+1)$. Compactness of $A$ yields a further subsequence and a limit $y_{n+1}\in A$ with $y_{n+1}^{(m)}\to y_{n+1}$.
Step 3 (Pass to the Limit). For every $i$, the continuity of $g_{i}$ gives $g_{i}\bigl{(}y_{n+1}^{(m)},y^{n}_{m},\psi_{m}\bigr{)}\to g_{i}\bigl{(}y_{n+1},y^{n},\psi\bigr{)}.$ Since at least $k$ of the $g_{i}$’s are non-negative for each $m$, suppose, for contradiction, that strictly fewer than $k$ are non-negative at the limit. Then there exists $\delta>0$ such that, for all sufficiently large $m$, fewer than $k$ indices satisfy $g_{i}\bigl{(}y_{n+1}^{(m)},y^{n}_{m},\psi_{m}\bigr{)}\geq-\delta$, contradicting the fact that every $m$ has $k$ indices with the same expression $\geq 0$. Hence at least $k$ of the limiting $g_{i}$’s are $\geq 0$, and
$\pi_{\psi}(y_{n+1},y^{n})\geq\frac{k}{n+1}.$
Step 4 (Take the Supremum). Because $y_{n+1}\in A$,
$\overline{\Pi}_{(y^{n},\psi)}(A)\geq\pi_{\psi}(y_{n+1},y^{n})\geq\frac{k}{n+1}=L.$
Thus $\limsup_{m\to\infty}\overline{\Pi}_{(y^{n}_{m},\psi_{m})}(A)\leq\overline{\Pi}_{(y^{n},\psi)}(A)$, establishing upper semicontinuity. ∎
Proof of Lemma 7.
Throughout, let $P\in\Delta_{\Theta}^{\mathrm{dens}}$ be fixed and write $\text{d}P/\text{d}\lambda_{\Theta}=p$.
Step 1. Continuity of the numerator $N_{P}$. Define
$g(\theta,y_{n+1},y^{n})\coloneqq k_{n+1}(\theta,(y^{n},y_{n+1}))p(\theta);$
We have $\lvert g\rvert\leq p(\theta)M(\theta)$ by (vii). Assumption (vi) gives continuity of $(y_{n+1},y^{n})\mapsto g(\theta,y_{n+1},y^{n})$ for $\lambda_{\Theta}$-almost all $\theta$. The envelope $pM\in L^{1}(\lambda_{\Theta})$ by (vii); Hence, dominated convergence yields continuity of $(y_{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$.
Step 2. Continuity and positivity of the denominator $D_{P}$. The integrand in $D_{P}(y^{n})$ is $g$, but “stopped” at observation $n$, so the same argument as Step 1 shows that $y^{n}\mapsto D_{P}(y^{n})$ is continuous. Positivity follows from (viii).
Step 3. Joint (and uniform) continuity of the predictive. Because $Y^{n+1}$ is compact by (i) and $D_{P}$ never vanishes,
$p_{P}(y_{n+1}\mid y^{n})=\frac{N_{P}(y_{n+1},y^{n})}{D_{P}(y^{n})}$
is continuous in $(y_{n+1},y^{n})$ and, by compactness, uniformly continuous. This proves joint continuity of the map $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$.
Step 4. Continuity in the prior. Let $P_{m}\to P$ in $L^{1}(\lambda_{\Theta})$, with densities $p_{m}\coloneqq\text{d}P_{m}/\text{d}\lambda_{\Theta}$. From (vii),
$\bigl{\|}N_{P_{m}}-N_{P}\bigr{\|}_{\infty}\leq\int_{\Theta}\lvert p_{m}-p\rvert M\text{d}\lambda_{\Theta}\xrightarrow[m\to\infty]{}0,$
and similarly $\|D_{P_{m}}-D_{P}\|_{\infty}\to 0.$ Because $D_{P}$ is bounded below on the compact set $Y^{n}$ (by Step 2 there is a constant $\underline{d}_{P}\coloneqq\min_{y^{n}}D_{P}(y^{n})>0$) and $D_{P_{m}}$ converges uniformly, we have
$\inf_{y^{n}}D_{P_{m}}(y^{n})\geq\frac{1}{2}\underline{d}_{P}\quad\text{for all sufficiently large }m.$
Consequently
$\bigl{\|}p_{P_{m}}(\cdot\mid y^{n})-p_{P}(\cdot\mid y^{n})\bigr{\|}_{\infty}\xrightarrow[m\to\infty]{}0\quad\text{uniformly in }y^{n}\in Y^{n},$
and division remains well behaved thanks to the common lower bound $\tfrac{1}{2}\underline{d}_{P}$.
Step 5. Continuity of $\mathsf{BCP}$. Fix a point $(y^{n},P)$ in the domain. If $y^{n}_{m}\to y^{n}\text{ and }P_{m}\to P$ (as in Step 4), then
$\|\psi_{(y^{n}_{m},P_{m})}-\psi_{(y^{n},P)}\|_{\infty}\leq\|p_{P_{m}}(\cdot\mid y^{n}_{m})-p_{P}(\cdot\mid y^{n}_{m})\|_{\infty}+\|p_{P}(\cdot\mid y^{n}_{m})-p_{P}(\cdot\mid y^{n})\|_{\infty}.$
The first term vanishes uniformly in $y^{n}$ by Step 4, and the second vanishes by Step 3. Hence, $\|\psi_{(y^{n}_{m},P_{m})}-\psi_{(y^{n},P)}\|_{\infty}\to 0$, and the first component $y^{n}_{m}\to y^{n}$ by assumption. Therefore, $\mathsf{BCP}$ is continuous. ∎
Proof of Lemma 8.
Fix $(c_{0},y^{n}_{0},P_{0})$ and let $V\supset H(c_{0},y^{n}_{0},P_{0})$ be open in $Y$. Put $C\coloneqq Y\setminus V$, which is compact. Define
$\Phi(c,y^{n},P)\coloneqq\sup_{y\in C}\bigl{(}p_{P}(y\mid y^{n})-c\bigr{)}.$
Since $(y,c,y^{n},P)\mapsto p_{P}(y\mid y^{n})-c$ is continuous and $C$ is compact, $\Phi$ is upper semicontinuous. Moreover, for every $y\in C$ we have $p_{P_{0}}(y\mid y^{n}_{0})<c_{0}$, hence $\Phi(c_{0},y^{n}_{0},P_{0})<0$. By upper semicontinuity, there exists a neighborhood $U$ of $(c_{0},y^{n}_{0},P_{0})$ with $\Phi(c,y^{n},P)<0$ for all $(c,y^{n},P)\in U$. Equivalently, $H(c,y^{n},P)\cap C=\emptyset$, i.e. $H(c,y^{n},P)\subset V$. Thus $H$ is upper hemicontinuous in $(c,y^{n},P)$ and closed-valued.
Next we show continuity of $c(\alpha,y^{n},P)$. By Lemma 7, $(y,y^{n},P)\mapsto p_{P}(y\mid y^{n})$ is continuous and $Y$ is compact, so $(c,y^{n},P)\mapsto F(c;y^{n},P)=\int_{\{p_{P}(\cdot\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\lambda_{Y}(\text{d}y)$ is jointly continuous at every point where the level set $\{y:p_{P}(y\mid y^{n})=c\}$ has zero $p_{P}(\cdot\mid y^{n})$-mass. Assumption (ix) (continuity and strict increase of $F$ in $c$ at the quantile) ensures this holds at $c=c(\alpha,y^{n},P)$. Therefore, $F$ is jointly continuous in $(c,y^{n},P)$ near the target, and the parametric inverse function theorem for monotone maps implies that $(\alpha,y^{n},P)\mapsto c(\alpha,y^{n},P)$ is continuous.
Finally, we put everything together to show the upper hemicontinuity of $\mathsf{QUANT}$. View $\mathsf{QUANT}$ as the composition
$(\alpha,y^{n},P)\xmapsto{c}\bigl{(}c(\alpha,y^{n},P),y^{n},P\bigr{)}\xmapsto{H}H\bigl{(}c(\alpha,y^{n},P),y^{n},P\bigr{)}.$
As we have shown before, $H$ is upper hemicontinuous in its three arguments, and $c(\alpha,y^{n},P)$ is continuous, hence the singleton-valued correspondence associated with $c$ (that we still denote by $c$ for ease of notation) is upper hemicontinuous. The composition of two upper hemicontinuous correspondences is again upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Therefore, $\mathsf{QUANT}$ is upper hemicontinuous. ∎
Proof of Proposition 11.
We prove our statement in four steps.
Step 1 (Maxitivity versus additivity). For disjoint $A,B\in\Sigma_{Y}$,
$\overline{\Pi}(A\cup B)=\max\{\overline{\Pi}(A),\overline{\Pi}(B)\},\qquad\overline{\Pi}_{\mathrm{B}}(A\cup B)=\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B).$
Equality of $\overline{\Pi}$ and $\overline{\Pi}_{\mathrm{B}}$ forces
$\max\bigl{\{}\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\bigr{\}}=\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(B).$ (17)
Step 2 (A 0-11 law). If $\overline{\Pi}_{\mathrm{B}}(A),\overline{\Pi}_{\mathrm{B}}(B)\in(0,1)$, the right-hand side of (17) is strictly larger than the left, a contradiction. Hence
$\overline{\Pi}_{\mathrm{B}}(A)\in\{0,1\},\qquad\forall A\in\Sigma_{Y}.$ (18)
Equivalently, taking $B=A^{c}$ in (17) and using $\overline{\Pi}_{\mathrm{B}}(A)+\overline{\Pi}_{\mathrm{B}}(A^{c})=1$ yields $\overline{\Pi}_{\mathrm{B}}(A)\in\{0,1\}$ for every $A\in\Sigma_{Y}$.
Step 3 (0-1 regular Borel probabilities on compact Hausdorff are Dirac). By (18), $\overline{\Pi}_{\mathrm{B}}$ is $\{0,1\}$-valued. Let $\mathcal{F}\coloneqq\{F\subseteq Y:F\text{ closed and }\overline{\Pi}_{\mathrm{B}}(F)=1\}$. For $F_{1},F_{2}\in\mathcal{F}$,
$\overline{\Pi}_{\mathrm{B}}(F_{1}\cap F_{2})\geq\overline{\Pi}_{\mathrm{B}}(F_{1})+\overline{\Pi}_{\mathrm{B}}(F_{2})-1=1,$
so $\mathcal{F}$ has the finite intersection property. Compactness yields $K\coloneqq\bigcap_{F\in\mathcal{F}}F\neq\emptyset$; fix $y^{\ast}\in K$. If $U$ is any open neighborhood of $y^{\ast}$, then $Y\setminus U$ is closed and cannot belong to $\mathcal{F}$ (else it would contain $y^{\ast}$), hence $\overline{\Pi}_{\mathrm{B}}(Y\setminus U)=0$ and so $\overline{\Pi}_{\mathrm{B}}(U)=1$. Since $\overline{\Pi}_{\mathrm{B}}$ is regular on the compact Hausdorff space YY, [^20] On a compact Hausdorff space, every finite Borel measure is regular. Hence $\overline{\Pi}_{\mathrm{B}}$ is regular regardless of any domination by $\lambda_{Y}$. outer regularity gives
$\overline{\Pi}_{\mathrm{B}}(\{y^{\ast}\})=\inf\{\overline{\Pi}_{\mathrm{B}}(U):U\text{ open},y^{\ast}\in U\}=1.$
Thus $\overline{\Pi}_{\mathrm{B}}=\delta_{y^{\ast}}$.
Step 4 (Collapse of the posterior and the likelihood). From (15) and Step 3,
$\overline{\Pi}_{\mathrm{B}}(\cdot)=\int_{\Theta}L_{\theta}(\cdot)\overline{P}(\mathrm{d}\theta\mid y^{n})=\delta_{y^{\ast}}(\cdot).$
Hence $\overline{P}\big{(}\Theta^{\ast}\mid y^{n}\big{)}=1$ for $\Theta^{\ast}\coloneqq\{\theta:L_{\theta}=\delta_{y^{\ast}}\}$. In particular, if the model is identifiable, then $\Theta^{\ast}=\{\theta^{\ast}\}$ and $\overline{P}(\cdot\mid y^{n})=\delta_{\theta^{\ast}}$. ∎
Appendix C Background in Functional Analysis
The concepts we present in this section are studied in detail in Aliprantis and Border, (2006, Chapters 17 and 18).
A correspondence $\Phi:X\rightrightarrows Y$ from a set $X$ to a set $Y$ assigns to each $x\in X$ a subset $\Phi(x)\in 2^{Y}$. The image of a set $A\subseteq X$ under $\Phi$ is the set $\Phi(A)\coloneqq\cup_{x\in A}\Phi(x)$. Given two correspondences $\Phi:X\rightrightarrows Y$ and $\Psi:Y\rightrightarrows Z$, their composition is defined naturally as
$\Psi\circ\Phi:X\rightrightarrows Z,\quad(\Psi\circ\Phi)(x)\coloneqq\bigcup_{y\in\Phi(x)}\Psi(y).$
The graph of $\Phi$ is given by $\text{Gr}\Phi\coloneqq\{(x,y)\in X\times Y:y\in\Phi(x)\}$.
We now turn our attention to continuity and measurability notions for correspondences.
Definition 25 (Upper Hemicontinuity).
Let $X$ and $Y$ be topological spaces. A correspondence $\Phi:X\rightrightarrows Y$ is upper hemicontinuous at $x_{0}\in X$ if for every open set $V\subseteq\Phi(x_{0})$, there exists an open neighborhood $U$ of $x_{0}$ such that for all $x\in U$, $\Phi(x)\subseteq V$.
We say that $\Phi$ is upper hemicontinuous (u.h.c.) if it is upper hemicontinuous at every point $x\in X$.
As the name suggests, an alternative notion of continuity for correspondences, called lower hemicontinuity, can be defined (Aliprantis and Border,, 2006, Definition 17.2). In this work, we focus primarily on u.h.c. for its historical importance, especially in fields such as economics and game theory (Börgers,, 1991).
The following is the definition of how a correspondence can be considered measurable in the weakest possible sense.
Definition 26 (Weakly Measurable Correspondence).
Let $(X,\Sigma_{X})$ be a measurable space and $(Y,\tau_{Y})$ a topological space. A correspondence $\Phi:X\rightrightarrows Y$ is said to be weakly measurable if for every open set $O\subseteq Y$, the set
$\Phi^{u}(O)\coloneqq\{x\in X:\Phi(x)\cap O\neq\emptyset\}\in\Sigma_{X}.$
Notice that weak measurability has nothing to do with weak topologies, and that other (stronger) measurability notions can be found in Aliprantis and Border, (2006, Definition 18.1).
Definition 27 (Uniformly Compact-Valued Correspondences).
Let $(X,\Sigma_{X})$ and $(Y,\Sigma_{Y})$ be measurable Polish spaces. A correspondence $\Phi\colon X\rightrightarrows Y$ is called uniformly compact-valued if
  1. (i)
    (compact fibers) $\Phi(x)\subseteq Y$ is nonempty compact, for every $x\in X$;
  2. (ii)
    (uniform compact bound) there exists a one and the same compact set $K_{\Phi}\subseteq Y$ such that $\Phi(x)\subseteq K_{\Phi}$, for all $x\in X$.
Finally, we also recall the definition of the Vietoris topology.
Definition 28 (Vietoris Topology).
Let $X$ be a compact Hausdorff space. The Vietoris topology on the hyperspace $\mathcal{K}(X)$ of nonempty compact subsets of $X$ is the smallest topology making, for every open $U\subseteq X$,
$\langle U\rangle\coloneqq\{K\in\mathcal{K}(X):K\cap U\neq\emptyset\}\quad\text{and}\quad[U]\coloneqq\{K\in\mathcal{K}(X):K\subseteq U\}$
open. Equivalently, it is generated by finite intersections $U_{0}]\cap\langle U_{1}\rangle\cap\cdots\cap\langle U_{m}\rangle$ and, in particular, by the basic “hit-and-cover” sets
$\bigl{\langle}U_{1},\dots,U_{n}\bigr{\rangle}\coloneqq\{K\in\mathcal{K}(X):K\subseteq\textstyle\bigcup_{i=1}^{n}U_{i}\text{and}K\cap U_{i}\neq\emptyset\text{, }\forall i\},$
since $\langle U_{1},\dots,U_{n}\rangle=[U_{1}\cup\cdots\cup U_{n}]\cap\bigcap_{i=1}^{n}\langle U_{i}\rangle$.
Appendix D Background in Category Theory
In this section, we brush up on some Category Theory basic notions that we use throughout the paper when deriving our main results. The definitions that we provide all come from Perrone, (2024).
Definition 29 (Category).
A category $\mathbf{C}$ consists of
  • a collection $\mathbf{C}_{0}$ of objects (written $X,Y,Z,\dots$);
  • a collection $\mathbf{C}_{1}$ of morphisms (written $f,g,h,\dots$);
together with
  1. (1)
    for every morphism $f$ there are two distinguished objects called its source and target, written $s(f)$ and $t(f)$ (we abbreviate $f\colon X\rightarrow Y$ when $s(f)=X$, $t(f)=Y$);
  2. (2)
    for every object $X$ an identity morphism $\mathrm{id}_{X}\colon X\rightarrow X$;
  3. (3)
    for every composable pair $X\xrightarrow{f}Y\xrightarrow{g}Z$ (that is, for which $t(f)=s(g)$) a composite morphism $g\circ f\colon X\rightarrow Z$;
satisfying the unitality laws $f\circ\mathrm{id}_{X}=f=\mathrm{id}_{Y}\circ f$ and the associativity law $(h\circ g)\circ f=h\circ(g\circ f)$, whenever the expressions are well defined.
Categories are the building blocks in CT. It is natural, then, to ask how to relate them.
Definition 30 (Functor).
Let $\mathbf{C},\mathbf{D}$ be categories. A functor $F\colon\mathbf{C}\rightarrow\mathbf{D}$ assigns
  • to every object $X$ of $\mathbf{C}$ an object $FX$ of $\mathbf{D}$;
  • to every morphism $f\colon X\rightarrow Y$ in $\mathbf{C}$ a morphism $Ff\colon FX\rightarrow FY$ in $\mathbf{D}$,
such that
$F(\mathrm{id}_{X})=\mathrm{id}_{FX},\qquad F(g\circ f)=Fg\circ Ff$
for all composable $f,g$ in $\mathbf{C}$.
A functor $F\colon\mathbf{C}\rightarrow\mathbf{D}$ is called faithful if, for every objects $X,X^{\prime}$ of $\mathbf{C}$ and every morphism $f,f^{\prime}:X\rightarrow X^{\prime}$ of $\mathbf{C}$, we have that $Ff=Ff^{\prime}$ in $\mathbf{D}$ implies $f=f^{\prime}$ in $\mathbf{C}$.
Following up on the question that we asked ourselves before, we can query whether there is a convenient choice of relating two functors.
Definition 31 (Natural transformation).
Let $\mathbf{C},\mathbf{D}$ be categories and let
$F,G:\mathbf{C}\rightarrow\mathbf{D}$
be functors. A natural transformation $\alpha:F\Rightarrow G$ assigns to every object $X$ of $\mathbf{C}$ a morphism
$\alpha_{X}:FX\rightarrow GX$
such that for every morphism $f:X\rightarrow Y$ in $\mathbf{C}$
$Gf\circ\alpha_{X}=\alpha_{Y}\circ Ff.$
Natural transformation $\alpha$ is a natural isomorphism if for each object $X$ of $\mathbb{C}$, the component $\alpha_{X}:FX\rightarrow GX$ is an isomorphism. In that case, functors $F$ and $G$ are said to be (naturally) isomorphic.
Then, we turn our attention to a special structure on a category, called a monad, that can be thought of as a consistent way of extending spaces to include generalized elements and generalized functions of a specific type. It is of special interest in computer science.
Definition 32 (Monad on a category).
For a category $\mathbf{C}$, a monad is a triple
$(T,\eta,\mu)$
consisting of
  • a functor $T:\mathbf{C}\rightarrow\mathbf{C}$;
  • a natural transformation (the unit)
    $\eta:\operatorname{id}_{\mathbf{C}}\Rightarrow T;$
  • a natural transformation (the multiplication)
    $\mu:T\circ T\Rightarrow T;$
subject to the associativity and unit laws
$\mu\circ T\mu=\mu\circ\mu T,\qquad\mu\circ T\eta=\mu\circ\eta T=\operatorname{id}_{T}.$
Finally, we are interested in finding a category whose objects and morphisms can be multiplied together in a way that is associative and unital, similarly to what we have for monoids (but only up to isomorphisms). Such categories, called monoidal categories, have wide applications, ranging from information theory to probability.
Definition 33 (Monoidal category).
A monoidal category is a tuple
$(\mathbf{C},\otimes,I,\alpha,\lambda,\rho)$
where
  • $\mathbf{C}$ is a category;
  • $\otimes:\mathbf{C}\times\mathbf{C}\rightarrow\mathbf{C}$ is a product, the tensor product (Perrone,, 2024, Definition 1.3.49); [^21] We do not introduce the tensor product formally in this background section because its definition follows the natural intuition of how a (cartesian) product behaves.
  • II is an object of $\mathbf{C}$ (the unit);
  • $\alpha_{A,B,C}:A\otimes(B\otimes C)\rightarrow(A\otimes B)\otimes C$ (associator), $\lambda_{A}:I\otimes A\rightarrow A$, $\rho_{A}:A\otimes I\rightarrow A$ are natural isomorphisms,
satisfying Mac Lane’s pentagon and triangle equalities, for all objects $A,B,C,D$ of $\mathbf{C}$
$\alpha_{A,B,C\otimes D}\circ\alpha_{A\otimes B,C,D}=(\alpha_{A,B,C}\otimes\operatorname{id}_{D})\circ\alpha_{A,B\otimes C,D}\circ(\operatorname{id}_{A}\otimes\alpha_{B,C,D}),$
$(\operatorname{id}_{A}\otimes\lambda_{B})\circ\alpha_{A,I,B}=\rho_{A}\otimes\operatorname{id}_{B}.$